### v2

# UniHR: Hierarchical Representation Learning for Unified Knowledge Graph Link Prediction

Abstract
Beyond-triple fact representations including hyper-relational facts with auxiliary key-value pairs, temporal facts with additional timestamps, and nested facts implying relationships between facts, are gaining significant attention. However, existing link prediction models are usually designed for one specific type of facts, making it difficult to generalize to other fact representations. To overcome this limitation, we propose a Unified Hierarchical Representation learning framework (UniHR) for unified knowledge graph link prediction. It consists of a unified Hierarchical Data Representation (HiDR) module and a unified Hierarchical Structure Learning (HiSL) module as graph encoder. The HiDR module unifies hyper-relational KGs, temporal KGs, and nested factual KGs into triple-based representations. Then HiSL incorporates intra-fact and inter-fact message passing, focusing on enhancing the semantic information within individual facts and enriching the structural information between facts. Experimental results across 7 datasets from 3 types of KGs demonstrate that our UniHR outperforms baselines designed for one specific kind of KG, indicating strong generalization capability of HiDR form and the effectiveness of HiSL module. Code and data are available at https://github.com/Lza12a/UniHR.
UniHR: Hierarchical Representation Learning for Unified Knowledge Graph Link Prediction
Zhiqiang Liu1,3, Mingyang Chen2,3, Yin Hua1,3, Zhuo Chen2,3, Ziqi Liu4, Lei Liang4, Huajun Chen2,3, Wen Zhang1,3thanks: Corresponding Author. 1School of Software Technology, Zhejiang University 2College of Computer Science and Technology, Zhejiang University 3ZJU-Ant Group Joint Lab of Knowledge Graph 4Ant Group {zhiqiangliu,zhang.wen}@zju.edu.cn
1 Introduction
Large-scale knowledge graphs (KGs) such as WordNet Miller (1995), Freebase Bollacker et al. (2008), and Wikidata Vrandečić and Krötzsch (2014) have been widely applied in many areas like question answering Kaiser et al. (2021), recommendation systems Guo et al. (2020), and natural language processing Annervaz et al. (2018). However, the presence of missing facts within these KGs inevitably limit their applications. Therefore, the link prediction task has been introduced to predict missing elements within factual data. Current link prediction methods mainly focus on facts in the form of triple $\left(head\,entity,relation,tail\,entity\right)$.
Figure 1: A special KG consists of triple-based fact, hyper-relational fact, nested fact and temporal fact.
Despite the simplicity and unity of triple-based representation, it is difficult to adequately express complex facts, such as “Oppenheimer is educated at Harvard University for a bachelor degree in chemistry" shown in Figure 1. Therefore, existing researches Wang et al. (2021); Xiong et al. (2024); Xu et al. (2019) contribute to focusing on semantically richer facts. Figure 1 illustrates three specific types of facts: hyper-relational fact ((Oppenheimer, educated at, Harvard University), degree: bachelor, major: chemistry), temporal fact (Oppenheimer, honored with, Fermi Prize, 1963), nested fact ((Oppenheimer, born in, New York), imply, (Oppenheimer, nationality, The United States)). These forms of facts allow for expression of complex semantics and revelation of relationships between facts, extending beyond the triple-based representation. Thus in recent years, Hyper-relational KGs (HKG) Chung et al. (2023), Temporal KGs (TKG) Xu et al. (2023a), and Nested factual KGs (NKG) Xiong et al. (2024) attract wide research interests.
Recent studies have demonstrated the effectiveness of various embedding strategies for these beyond-triple representations Xiong et al. (2023). However, these methods are usually designed for specific representation forms, e.g., StarE Galkin et al. (2020) customizes graph neural network to implement message passing on hyper-relational facts, For nested factual KGs, BiVE Chung and Whang (2023) connects two levels of facts throgh a simple linear layer. In addition, GeomE+ Xu et al. (2023a) et al. temporal KG embedding methods contain time-aware scoring functions to adapt timestamps. Although these methods perform well on specific type of facts, it is evident that such customized methods are difficult to generalize to other types of KGs. Therefore, establishing a unified representation learning method for multiple types of KGs is worth to investigate.
To overcome the challenges mentioned above, we propose a Unified Hierarchical Representation learning method (UniHR), which includes a Hierarchical Data Representation (HiDR) module and a Hierarchical Structure Learning (HiSL) module as the graph encoder. HiDR module standardizes hyper-relational facts, nested factual facts, and temporal facts into the form of triples without loss of information. Furthermore, HiSL module captures local semantic information during intra-fact message passing and then utilizes inter-fact message passing to enrich the global view of nodes to obtain better node embeddings based on HiDR form. Finally, the updated embeddings are fed into decoders for link prediction. Experimental results demonstrate that our UniHR achieves state-of-the-art performance on HKG and NKG datasets, and competitive performance on TKG datasets, revealing strong generalization capability of HiDR form and effectiveness of HiSL module. Our contributions can be summarized as follows:
  • 1.
    We emphasize the value of investigating unified KG representation method, including unified symbolic representation and unfied representation learning method for different KGs.
  • 2.
    We propose the first unified KG representation learning framework UniHR, across different types of KGs, including a hierarchical data representation module and a hierarchical structure learning module.
  • 3.
    We conduct link prediction experiments on 7 datasets across 3 types of KGs. Compared to methods designed for one kind of KG, UniHR achieves the best or competitive results, verifying strong generalization capability.
2 Preliminaries
In this section, we introduce the definition of four types of existing knowledge graphs (KGs): triple-based KG, hyper-relational KG, nested factual KG and temporal KG, along with link prediction tasks on these types of KGs.
Triple-based Knowledge Graph.
A common triple-based KG $\mathcal{G}_{KG}\,\text{=}\,\{\mathcal{V},\mathcal{R},\mathcal{F}\}$ represents facts as triples, denoted as =$\mathcal{F}\,\text{=}\{\left(h,r,t\right)|h,t\in\mathcal{V},$ $r\in\mathcal{R}\}$, where $\mathcal{V}$ is the set of entities and $\mathcal{R}$ is the set of relations.
Link Prediction on Triple-based KG. The link prediction on triple-based KGs involves answering a query $\left(h,r,?\right)$ or (?$,r,t\right)$, where the missing element ‘????’ is an entity in $\mathcal{V}$.
Hyper-relational Knowledge Graph.
A hyper-relational KG (HKG) $\mathcal{G}_{HKG}\,\text{=}\,\{\mathcal{V},\mathcal{R},\mathcal{F}\}$ consists of hyper-relational facts, abbreviated as H-Facts, denoted as =$\mathcal{F}\,\text{=}\{((h,r,t),\{(k_{i}\text{:}\,v_{i})\}_{i=1}^{m})$ $|\,h,t,v_{i}\in\mathcal{V},r,k_{i}\in\mathcal{R}\}$. Typically, we refer to $\left(h,r,t\right)$ as the main triple in the H-Fact and $\left\{\left(k_{i}\text{:}\,v_{i}\right)\right\}_{i=1}^{m}$ as $m$ auxiliary key-value pairs.
Link Prediction on Hyper-relational KG. Similar to link prediction on triple-based KGs, the link prediction on HKGs aims to predict entities in the main triple or the key-value pairs. Symbolically, the aim is to predict the missing element, denoted as ‘?’ for queries $((h,r,t),(k_{1}\text{:}\,v_{1}),\ldots(k_{i}\text{:}\,?)),$ ((?$,r,t),\{(k_{i}\text{:}v_{i})\}_{i=1}^{m})$ or $((h,r,?),\{(k_{i}\text{:}v_{i})\}_{i=1}^{m})$.
Nested Factual Knowledge Graph.
A nested factual KG (NKG) can be represented as $\mathcal{G}_{NKG}\,\text{=}\,\{\mathcal{V},\mathcal{R},\mathcal{F},\hat{\mathcal{R}},\hat{\mathcal{F}}\}$, which is composed of two levels of facts, called atomic facts and nested facts. =$\mathcal{F}\,\text{=}\left\{\left(h,r,t\right)|h,t\in\mathcal{V},r\in\mathcal{R}\right\}$ is the set of atomic facts, where $\mathcal{V}$ is a set of atomic entities and $\mathcal{R}$ is a set of atomic relations. ^=$\hat{\mathcal{F}}\,\text{=}\,\{\left(\mathcal{F}_{i},\hat{r},\mathcal{F}_{j}\right)|$ $\mathcal{F}_{i},\mathcal{F}_{j}\in\mathcal{F},\hat{r}\in\hat{\mathcal{R}}\}$ is the set of nested facts, where $\hat{\mathcal{R}}$ is the set of nested relations.
Link Prediction on Nested Factual KG. The link prediction on the NKGs is performed on the atomic facts or nested facts. We refer to the link prediction on atomic facts as Base Link Prediction, and the link prediction on nested facts as Triple Prediction. For base link prediction, given a query $\left(h,r,?\right)$ or (?$,r,t\right)$, the aim is to predict the missing atomic entity ‘????’ from $\mathcal{V}$. For triple prediction, given a query (?$,\hat{r},\mathcal{F}_{j}\right)$ or $\left(\mathcal{F}_{i},\hat{r},?\right)$, the aim is to predict the missing atomic fact ‘????’ from $\mathcal{F}$.
Temporal Knowledge Graph.
A temporal KG (TKG) $\mathcal{G}_{TKG}\,\text{=}\,\{\mathcal{V},\mathcal{R},\mathcal{F},\mathcal{T}\}$ is composed of quadruple-based facts, which can be represented as =$\mathcal{F}\,\text{=}\,\{(h,r,t,[\tau_{b},\tau_{e}])|h,t\in\mathcal{V},r\in\mathcal{R},\tau_{b},\tau_{e}\in\mathcal{T}\}$, where $\tau_{b}$ is the begin time, $\tau_{e}$ is the end time, $\mathcal{V}$ is the set of entities, $\mathcal{R}$ is the set of relations and $\mathcal{T}$ is the set of timestamps.
Link Prediction on Temporal KG. The link prediction on TKGs aims to predict missing entities ‘????’ in $\mathcal{V}$ for two types of queries (?$,r,t,[\tau_{b},\tau_{e}]\right)$ or $\left(h,r,?,[\tau_{b},\tau_{e}]\right)$.
3 Related Works
Link Prediction on Hyper-relational Knowledge Graph.
Earlier HKG representation learning methods e.g., m-TransH Wen et al. (2016), RAE Zhang et al. (2018) have generalized the triple-based approach to HKG and loosely represent the combinations of key-value pairs. Galkin et al. customize StarE Galkin et al. (2020) based on CompGCN Vashishth et al. (2019) for hyper-relational facts to capture the interaction information of key-value pairs with the main triple in the message passing stage, and achieves impressive results, demonstrating that the structural information of the graph in HKGs is also important. GRAN Guan et al. (2021) introduces edge-aware bias into the vanilla transformer attention Vaswani et al. (2017), while HyNT Chung et al. (2023) designs a qualifier encoder for HKG. They both focus on intra-fact dependencies but ignore the global structural information. Due to the existence of its particular key-value pairs on H-Facts, there are many limitations in capturing global structural information through the widely available triple-based GNNs.
Link Prediction on Nested Factual Knowledge Graph.
Chung et al. Chung and Whang (2023) first introduced the concept of nested facts. They also propose BiVE which projects atomic facts to fact nodes in the encoding phase via a simple linear layer and scores both atomic facts and nested facts using the quaternion-based KGE scoring functions like QuatE Zhang et al. (2019) or BiQUE Guo and Kok (2021). Based on BiVE, NestE Xiong et al. (2024) represents the fact nodes as a $1\times 3$ embedding matrix and the nested relations as a $3\times 3$ matrix to avoid information loss, embedding them into hyperplanes with different dimensions. These methods only capture the semantic information between atomic facts and nested facts while ignoring global structural information. Meanwhile, due to the complexity of this representation, common triple-based GNNs have difficulty in message passing between atomic fact and nested fact.
Link Prediction on Temporal Knowledge Graph.
Recent studies in temporal knowledge graph representation learning have focused on enhancing performance by designing special time-aware scoring functions. Models such as TTransE Leblay and Chekol (2018), HyTE Dasgupta et al. (2018), TeRo Xu et al. (2020), and TGeomE+ Xu et al. (2023a) incorporate temporal-aware module into the KGE score function in various ways. However, existing models for link prediction seldom directly utilize GNNs to perceive time information for enhancing entity and relation embeddings. Therefore, we believe it is a direction worth exploring.
4 Methodology
In this section, we introduce our method, a Unified Hierarchical Representation learning framework (UniHR), which includes a Hierarchical Data Representation (HiDR) module and a Hierarchical Structure Learning (HiSL) module. Our workflow can be divided into the following three steps: 1) Given a KG $\mathcal{G}$ of any type, we represent it into $\mathcal{G}^{HiDR}$ under the HiDR form. 2) The $\mathcal{G}^{HiDR}$ will be encoded by HiSL module with the enhancement of semantic information within individual facts and structural information between facts on the whole graph. 3) In the phase of decoding, the updated embeddings of nodes and edges are fed into transformer decoders to obtain the plausibility score of facts.
Figure 2: Diverse facts are translated into the HiDR form. Blue, light blue, and red circles denote atomic nodes, relation nodes, and fact nodes. Black, light blue, and red arrows denote atomic relations, connected relations, and nested relations. The triples they connect correspond to atomic facts, connected facts, and nested facts.
4.1 Hierarchical Data Representation
To overcome the differences in the representation of multiple types KGs, we introduce a Hierarchical Data Representation module, abbreviated as HiDR. Different from labelled RDF representation Ali et al. (2022), we constrain “triple" to be considered as the basic units of HiDR form, then HiDR could continuous benefit from the model developments of triple-based KGs, which is the most active area about link prediction over KGs.
As shown in Fig. 2, in order to ensure comprehensive representation of facts, we introduce three hierarchical types of nodes and three connected relations in HiDR. Inspired by the nested fact form Xiong et al. (2024), we denote original entities within three types of KGs as atomic nodes and complement fact nodes for HKGs and TKGs lacking a designated fact node. To facilitate the interaction between fact nodes and relations explicitly, we incorporate relation nodes into the graph, represented as $e_{r}$ for each $r$. These relation nodes are derived from transforming the relation edges in the original KG. It is important to facilitate direct access of fact nodes to the relevant atomic nodes during message passing process. To achieve this, we introduce three connected relations: has relation, has head entity and has tail entity, which establish directly connections between atomic nodes and fact nodes. Ultimately, we denote the (main) triple $\left(h,r,t\right)$ in original fact as three connected facts: $\left(f,has\,relation,e_{r}\right),\left(f,has\,head\,entity,h\right),$ $(f,has\,tail\,entity,t)$, and a atomic fact $(h,r,t)$, where $f$ is fact node. Formally, the definition of HiDR form is as follows:
Definition 1.
Hierarchical Data Representation:
A KG represented as the HiDR form is denoted as $\mathcal{G}^{HiDR}\,\text{=}\,\{\mathcal{V}^{HiDR},\mathcal{R}^{HiDR},\mathcal{F}^{HiDR}\},$ where $\mathcal{V}^{HiDR}\,\text{=}\,\mathcal{V}_{a}\cup\mathcal{V}_{r}\cup\mathcal{V}_{f}$ is a joint set of atomic node set ($\mathcal{V}_{a}$), relation node set ($\mathcal{V}_{r}$), fact node set ($\mathcal{V}_{f}$). $\mathcal{R}^{HiDR}\,\text{=}\,\mathcal{R}_{a}\cup\mathcal{R}_{n}\cup\mathcal{R}_{c}$ is a joint set of atomic relation set ($\mathcal{R}_{a}$), nested relation set ($\mathcal{R}_{n}$), connected relation set $\mathcal{R}_{c}\,\text{=}\{has\,relation,has\,head\,entity,$ $has\,tail\,entity\}$. The fact set $\mathcal{F}^{HiDR}\,\text{=}\,\mathcal{F}_{a}\cup\mathcal{F}_{c}\cup\mathcal{F}_{n}$ is jointly composed of three types of triple-based facts: atomic facts ($\mathcal{F}_{a}$), connected facts ($\mathcal{F}_{c}$) and nested facts ($\mathcal{F}_{n}$), where $\mathcal{F}_{a}\text{=}\,\{(v_{1},r,$ $v_{2})|\,v_{1},v_{2}\in\mathcal{V}_{a},r\in\mathcal{R}_{a}\}$, $\mathcal{F}_{c}\,\text{=}\,\{(v_{1},r,v_{2})|\,v_{1}\in\mathcal{V}_{f},r\in\mathcal{R}_{c},v_{2}\in\mathcal{V}_{a}\}$, $\mathcal{F}_{n}\,\text{=}\,\{(v_{1},r,v_{2})|\,v_{1},v_{2}\in\mathcal{V}_{f},r\in\mathcal{R}_{n}\}$.
Next, we introduce how to transform different types of KGs into HiDR form.
For hyper-relational knowledge graphs, we regard key-value pairs as complementary information for facts. Thus, we translate H-Facts $\mathcal{F}_{HKG}$ =$\text{=}\,\{((h,r,t),\{(k_{i}\text{:}\,v_{i})\}_{i=1}^{m}\}$ into the HiDR form that $\mathcal{G}^{HiDR}_{HKG}\,\text{=}\,\{\mathcal{V},\mathcal{R},\mathcal{F}^{HiDR}_{HKG}\}$ following the definition, where $\mathcal{F}_{c}\text{=}\{(f,has\,relation,e_{r}),(f,has$ $,\,head\,entity,h),(f,has\,tail\,entity,t),(f,k_{1},v_{1}),$$\ldots,(f,k_{m},v_{m})\}$, $\mathcal{F}_{a}\,\text{=}\,\{(h,r,t)\,|\,((h,r,t),\{(k_{i}\text{:}\,$ $v_{i})\}_{i=1}^{m})\in\mathcal{F}_{HKG})\}$ and $\mathcal{F}_{n}\text{=}\,\varnothing$.
For nested factual knowledge graphs, HiDR can naturally represent hierarchical facts, so we translate the atomic facts $\mathcal{F}_{NKG}\,\text{=}\,\{(h_{i},r_{i},t_{i})\}$ and the nested facts ^$\hat{\mathcal{F}}_{NKG}\,\text{=}\,\{((h_{1},r_{1},t_{1}),R,(h_{2},$ $r_{2},t_{2}))|(h_{i},r_{i},t_{i})\in\mathcal{F}_{NKG}\}$ into the form of HiDR that $\mathcal{G}^{HiDR}_{NKG}\,\text{=}\,\{\mathcal{V},\mathcal{R},\mathcal{F}^{HiDR}_{NKG}\}$ following the definition, where $\mathcal{F}_{a}\,\text{=}\,\{(h_{i},r_{i},t_{i})\,|\,(h_{i},r_{i},t_{i})\in$ $\mathcal{F}_{NKG}\}$, $\mathcal{F}_{c}\,\text{=}\{(f_{i},has\,head\,entity,h_{i}),(f_{i},has$ $tail\,entity,t_{i}),(f_{i},has\,relation,e_{r_{i}})|f_{i}\,\text{=}\,(h_{i},r_{i},$ $t_{i})\in\mathcal{F}_{NKG}\}$ and $\mathcal{F}_{n}\text{=}\{(f_{1},R,f_{2})|f_{i}\in\mathcal{F}_{NKG}\}$.
For temporal knowledge graphs, we regard the TKG as a special HKG, and convert timestamps to auxiliary key-value pairs in HKGs by adding two special atomic relations: begin and end, regarding timestamps as special numerical atomic nodes. Thus, we firstly translate the temporal facts in TKGs $\mathcal{F}_{TKG}\text{=}\,\{(h,r,t,[\tau_{b},\tau_{e}])\}$ into H-Facts form $\mathcal{F}_{TKG}^{HKG}\text{=}\,\{(h,r,t,begin\text{:}\tau_{b},end\text{:}\tau_{e})\}$. Then according to the previous transformation in HKG, it can be translated into the HiDR form that $\mathcal{G}_{TKG}^{HiDR}\,\text{=}\,\{\mathcal{V},\mathcal{R},\mathcal{F}_{TKG}^{HiDR}\}$ following the definition,where $\mathcal{F}_{a}\,\text{=}\{(h,r,t)\,|\,(h,r,t,begin\text{:}\,\tau_{b},end\text{:}$ $\tau_{e})\in\mathcal{F}_{TKG}^{HKG}\}$, $\mathcal{F}_{c}\,\text{=}\,\{(f,has\,relation,e_{r}),(f,$ $has\,head\,entity,h),\,(f,has\,tail\,entity,t),\,(f,$ $begin,\tau_{b}),(f,end,\tau_{e})\,|\,f\,\text{=}\,(h,r,t,begin\text{:}\,\tau_{b},end\text{:}$ $\,\tau_{e})\in\mathcal{F}_{TKG}^{HKG}\}$ and $\mathcal{F}_{n}\,\text{=}\,\varnothing$.
In summary, we could convert all above KGs into the HiDR form, and preserve the semantics in the original KGs without loss of information.
4.2 Hierarchical Structure Learning
In this section, we illustrate how various KGs in the form of HiDR can be encoded by the Hierarchical Structure Learning module, abbreviated as HiSL shown in Fig. 3.
Figure 3: HiSL module for intra-fact and inter-fact MP.
Representation Initialization.
We first initialize the embedding matrices $\mathbf{H}_{a}\in\mathbb{R}^{|\mathcal{V}_{a}|\times d}$ and $\mathbf{E}\in\mathbb{R}^{|\mathcal{R}^{HiDR}|\times d}$ for atomic nodes and all relation edges. Then we also initialize the embedding of relation node $\mathbf{H}_{r}\in\mathbb{R}^{|\mathcal{V}_{r}|\times d}$, which can be transformed from the relation edge $r$, define as:
$\mathbf{H}_{r}=\mathbf{E_{a}\cdot W_{r}},$ (1)
where $\mathbf{E_{a}}\subseteq\mathbf{E},\mathbf{W_{r}}\in\mathbb{R}^{d\times d}$ denote the atomic relation embeddings and a projection matrix. Then we initialize the fact node embeddings $\mathbf{H}_{f}$ to explicitly capture key information within facts by utilizing the embedding of (main) triple:
$\mathbf{h}_{f}=f_{m}([\mathbf{h}_{h};\mathbf{h}_{r};\mathbf{h}_{t}]),$ (2)
where $(h,r,t)\in\mathcal{F}_{a}$, the operation [;][$\cdot$;$\cdot$][ ⋅ ; ⋅ ] is the vector concatenation, $\mathbf{h}_{h},\mathbf{h}_{t}\subseteq\mathbf{H}_{a},\mathbf{h}_{r}\subseteq\mathbf{H}_{r}$ denote (main) triple embedding and $f_{m}\text{:}\;\mathbb{R}^{3d}\rightarrow\mathbb{R}^{d}$ is a 1-layer MLP.
For numerical atomic nodes, namely timestamps in temporal knowledge graphs, we utilize the Time2Vec Kazemi et al. (2019) to encode the timestamp $\tau$ into an embedding:
$\mathbf{h}_{\tau}=\omega_{p}\sin\left(f_{p}(\tau)\right)+f_{np}(\tau),$ (3)
where $f_{p}\text{:}\;\mathbb{R}^{1}\rightarrow\mathbb{R}^{d}$ is a 1-layer MLP as periodic function, $f_{np}\text{:}\;\mathbb{R}^{1}\rightarrow\mathbb{R}^{d}$ is a 1-layer MLP as non-periodic function, and $\omega_{p}\in\mathbb{R}^{1}$ is a learnable parameter for scaling the periodic features.
Intra-fact Message Passing.
In this stage, massage passing is conducted for fact nodes. Given a fact node $f_{k}\in\mathcal{V}_{f}$, we construct its constituent elements, i.e., one-hop neighbors, as the node set $\mathcal{V}_{k}\,\text{=}\,\{v\in\mathcal{N}_{f_{k}}\,|\,v\in\mathcal{V}_{a}\cup\mathcal{V}_{r}\}$, where $\mathcal{N}_{f_{k}}$ is the set of one-hop neighbors of fact node $f_{k}$. Then we retain the edges directly connected to fact node $f_{k}$, thereby constructing a subgraph $\mathcal{G}_{k}\,\text{=}\,\{\mathcal{V}_{k},\mathcal{R}_{k},\mathcal{F}_{k}\}\subseteq\mathcal{G}^{HiDR}$. For this subgraph, we employ the graph attention network (GAT) Brody et al. (2021) to aggregate local information, computing the attention score $\alpha_{i,j}$ between node $i\in\mathcal{V}_{k}$ and its neighbor $j$. The formula for calculating $\alpha_{i,j}$ in the $l$-th layer is as follows:
$\alpha_{i,j}^{l}=\frac{\exp\left(\mathbf{W}^{l}\left(\sigma\left(\mathbf{W}_{in}^{l}\mathbf{h}_{i}^{l}+\mathbf{W}_{out}^{l}\mathbf{h}_{j}^{l}\right)\right)\right)}{\sum\limits_{j^{\prime}\in\mathcal{N}_{i}}\exp\left(\mathbf{W}^{l}\left(\sigma\left(\mathbf{W}_{in}^{l}\mathbf{h}_{i}^{l}+\mathbf{W}_{out}^{l}\mathbf{h}_{j^{\prime}}^{l}\right)\right)\right)},$
where $\mathbf{h}_{i}^{l},\mathbf{h}_{j}^{l}\in\mathbb{R}^{d}$ represent the embeddings of node $i$ and its neighbor $j$ in $l$-th layer. And there are three learnable weight matrices $\mathbf{W}_{in}^{l},\mathbf{W}_{out}^{l}\in\mathbb{R}^{d\times d}$ and $\mathbf{W}^{l}\in\mathbb{R}^{d}$. We choose LeakyReLU as activation function $\sigma$. Then, the updated node embeddings are obtained by aggregating the information of neighbors according to the attention scores:
$\mathbf{h}_{i}^{l}=\mathbf{h}_{i}^{l}+\sum\limits_{j\in\mathcal{N}_{i}}\alpha_{i,j}^{l}\cdot\mathbf{W}_{out}^{l}\mathbf{h}_{j}^{l}.$
Inter-fact Message Passing.
At this stage, message passing is conducted on the whole graph $\mathcal{G}^{HiDR}$. Similar to previous work Vashishth et al. (2019), we use a non-parametric aggregation operator $\phi\left(\cdot\right)\text{:}\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ to obtain messages from neighbouring nodes and edges. Specifically, we employ the circular-correlation operator inspired from HolE Nickel et al. (2016), defined as:
$\phi\left(\mathbf{h}_{j},\mathbf{e}_{r}\right)=\mathbf{h}_{j}\star\mathbf{e}_{r}=\mathbf{F}^{-1}\left(\left(\mathbf{F}\mathbf{h}_{j}\right)\odot\overline{\left(\mathbf{F}\mathbf{e}_{r}\right)}\right)$
where $\mathbf{F}$ and $\mathbf{F}^{-1}$ denote the discrete fourier transform (DFT) matrix and its inverse matrix, the $\odot$ is the element-wise (Hadamard) product. Furthermore, in order to fully capture the heterogeneity of the graph, we classify edges along two dimensions: $\lambda(r)\in\left\{forward,reverse\right\}$ and $\tau(r)\in\{connected\,\,relation,atomic\,\,relation,nested\,$ $relation\}$ and adopt two relation-type specific learnable parameters $\mathbf{W}_{\lambda\left(r\right)}\in\mathbb{R}^{d\times d}$ and $\omega_{\tau\left(r\right)}\in$ $\mathbb{R}^{1}$ for more fine-grained aggregation:
$\mathbf{h}_{i}^{l+1}=\sum\limits_{(r,j)\in\mathcal{N}(i)}{\sigma\left(\omega_{\tau\left(r\right)}^{l}\right)\mathbf{W}_{{\lambda}\left(r\right)}^{{l}}\phi\left(\mathbf{h}_{j}^{l},\mathbf{e}_{r}^{l}\right)}+\mathbf{W}_{self}^{l}\mathbf{h}_{i}^{l}$
$\mathbf{e}_{r}^{l+1}=\mathbf{W}_{rel}^{l}\mathbf{e}_{r}^{l}$
where $\mathbf{W}^{l}_{self},\mathbf{W}^{l}_{rel}\in\mathbb{R}^{d\times d}$, $\sigma$ is a sigmoid activation function and $\mathcal{N}\left(i\right)$ is a set of immediate neighbors of $i$ for its outgoing edges $r$. We utilize $\phi\left(\cdot\right)$ to combine the information from both edge $r$ and node $j$, and then passes it to node $i$. Subsequently, node $i$ aggregates the information according to the types of edge $r$ separately to update its embedding, while edge $r$ is also projected into the same embedding space as the updated nodes.
4.3 Link Prediction Decoder
Since the query varies across different settings, we use the transformer Vaswani et al. (2017) as the decoder with the mask pattern. Specifically, we convert the updated node and edge embeddings into a sequence of fact embeddings, mask the elements to be predicted in facts with the $\bm{\left[M\right]}$ token as the input to the transformer. Finally, we obtain the embedding of output $\bm{\left[M\right]}$ in the last layer to measure the plausibility of the fact, denoted as $\mathbf{h}_{pre}$, and calculate the probability distribution of candidates, followed by training it using the cross-entropy loss function:
$P=\,\,{\rm Softmax}\left(f\left(\mathbf{h}_{pre}\right)[\mathbf{E};\mathbf{H}]^{\top}\right),$
$\mathcal{L}=\sum_{t=0}^{\left|\mathcal{R}\right|+\left|\mathcal{V}\right|}{y_{t}\log P_{t}}$
where $P\in\mathbb{R}^{\left|\mathcal{R}\right|+\left|\mathcal{V}\right|}$ represents the confidence scores of all candidates, $f\text{:}\;\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ is a 1-layer MLP, and $[\mathbf{E};\mathbf{H}]\in\mathbb{R}^{(\left|\mathcal{R}\right|+\left|\mathcal{V}\right|)\times d}$ is the embedding matrix of all candidate edges or nodes. The $P_{t}$ and $y_{t}$ are probability and ground truth of the $t$-th candidate. The final loss function $\mathcal{L}$ includes both node loss and edge loss during the predictions.
Table 1: Results of link prediction on HKG datasets. All baselines’ results are taken from Chung et al. (2023). The best results are written bold, while the second are underlined. # Tr. Params denotes the number of learnable parameters during training.
WikiPeople WD50K
Model # Tr. Params subject/object all entities # Tr. Params subject/object all entities
MRR Hits@1 Hits@10 MRR Hits@1 Hits@10 MRR Hits@1 Hits@10 MRR Hits@1 Hits@10
NaLP - 0.356 0.271 0.499 0.360 0.275 0.503 - 0.230 0.170 0.347 0.251 0.187 0.375
tNaLP - 0.358 0.288 0.486 0.361 0.290 0.490 - 0.221 0.163 0.331 0.243 0.182 0.360
RAM 27.34M 0.459 0.384 0.584 0.461 0.386 0.585 - 0.276 0.210 0.399 0.296 0.232 0.416
HINGE - 0.393 0.309 0.547 0.395 0.311 0.549 - 0.264 0.187 0.410 0.277 0.200 0.424
NeuInfer - 0.357 0.247 0.533 0.357 0.248 0.532 - 0.220 0.154 0.347 0.225 0.158 0.355
StarE 7.84M 0.458 0.364 0.611 - - - 10.39M 0.309 0.234 0.452 - - -
HyTransformer - 0.460 0.382 0.594 - - - - 0.304 0.231 0.443 - - -
GRAN 15.26M 0.462 0.366 0.610 0.465 0.371 0.613 18.51M 0.330 0.255 0.472 0.361 0.286 0.501
HyNT 23.02M 0.482 0.415 0.602 0.481 0.414 0.603 29.61M 0.333 0.259 0.474 0.360 0.287 0.500
UniHR 8.02M 0.491 0.417 0.618 0.493 0.420 0.621 10.55M 0.348 0.278 0.482 0.382 0.313 0.513
Table 2: Results of base link prediction (left) and triple prediction (right). All baselines’ results are taken from Xiong et al. (2024). For BiVE Chung and Whang (2023) and NestE Xiong et al. (2024), we pick their best variants.
Model FBHE/FBH DBHE FBH FBHE DBHE
MRR Hits@10 MRR Hits@10 #Tr. Params MR MRR Hits@10 #Tr. Params MR MRR Hits@10 #Tr. Params MR MRR Hits@10
Base link prediction Triple prediction
QuatE 0.354 0.581 0.264 0.440 - 145603.8 0.103 0.114 - 94684.4 0.101 0.209 - 26485.0 0.157 0.179
BiQUE 0.356 0.583 0.274 0.446 - 81687.5 0.104 0.115 - 61015.2 0.135 0.205 - 19079.4 0.163 0.185
Neural-LP 0.315 0.486 0.233 0.357 - 115016.6 0.070 0.073 - 90000.4 0.238 0.274 - 21130.5 0.170 0.209
DRUM 0.317 0.490 0.237 0.359 - 115016.6 0.069 0.073 - 90000.3 0.261 0.274 - 21130.5 0.166 0.209
AnyBURL 0.310 0.526 0.220 0.364 - 108079.6 0.096 0.108 - 83136.8 0.191 0.252 - 20530.8 0.177 0.214
BiVE 0.370 0.607 0.274 0.422 12.66M 6.20 0.855 0.941 12.67M 8.35 0.711 0.866 10.74M 3.63 0.687 0.958
NestE 0.371 0.608 0.289 0.443 12.11M 3.34 0.922 0.982 12.46M 3.05 0.851 0.962 10.22M 2.07 0.862 0.984
UniHR 0.401 0.619 0.300 0.455 4.12M 2.46 0.946 0.993 4.12M 5.20 0.793 0.890 3.67M 1.90 0.862 0.987
5 Experiment
5.1 Experiment Settings
Datasets.
For link prediction on HKGs, we select three benchmark datasets: WikiPeople Guan et al. (2021) , WD50K Galkin et al. (2020) and JF17K Wen et al. (2016). As for the NKGs, we choose FBH, FBHE and DBHE constructed by Chung and Whang (2023). Lastly, we use wikidata12k Dasgupta et al. (2018), a subset of wikidata Vrandečić and Krötzsch (2014) for link prediction on TKGs. The statics of datasets are given in Appendix D.
Evaluation metric.
We conduct link prediction across multiple settings, evaluating performance based on the rank of predicted facts. We use the MR (Mean Rank), MRR (Mean Reciprocal Rank) and Hits@K (K=1,3,10) as our evaluation metrics. And we choose to adopt the metrics used in prior works. Additionally, we employ filtering settings Bordes et al. (2013) during the evaluation process to eliminate existing facts in the dataset.
Baselines.
For link prediction on HKG, we compare our UniHR against NaLP Guan et al. (2021), tNaLP Guan et al. (2021), RAM Liu et al. (2021), HINGE Rosso et al. (2020), NeuInfer Guan et al. (2020), StarE Galkin et al. (2020), HyTransformer Yu and Yang (2021), GRAN Wang et al. (2021) and HyNT Chung et al. (2023). For link prediction on NKG, QuatE Zhang et al. (2019), BiQUE Guo and Kok (2021), Neural-LP Yang et al. (2017), DRUM Sadeghian et al. (2019), AnyBURL Meilicke et al. (2019), BiVE Chung and Whang (2023) and NestE Xiong et al. (2024) are chosen as baselines. BiVE and NestE are especially designed for NKG. We compare against following TKG link prediction methods: ComplEx-N3 Lacroix et al. (2018), TTransE Leblay and Chekol (2018), HyTE Dasgupta et al. (2018), TA-DistMult Garcia-Duran et al. (2018), ATiSE Xu et al. (2019), TeRo Xu et al. (2020), TASTER Wang et al. (2023), TGeomE+ Xu et al. (2023a).
Implementation details.
All experiments are conducted on a single Nvidia 80G A800 GPU and implemented with PyTorch. For base link prediction on NKGs, we also use augmented triples from Chung and Whang (2023) for training to ensure fairness. For triple prediction, due to the small size of training set, we conduct training based on fixed embeddings of entities obtained from the base link prediction and set $\omega_{nested\;relation}\text{=}0$ to prevent overfitting. We employ AdamW Kingma and Ba (2015) optimizer. Hyperparameters are chosen by using a grid search based on the MRR performance and details can be found in Appendix F.
5.2 Main Results
Table 3: Results of link prediction on wikidata12k.
Model wikidata12k
# Tr. Params MRR Hits@1 Hits@3 Hits@10
ComplEx-N3 - 0.248 0.143 - 0.489
TTransE - 0.172 0.096 0.184 0.329
HyTE 1.69M 0.253 0.147 - 0.483
TA-DistMult - 0.230 0.130 - 0.461
TeRo 12.91M 0.299 0.198 0.329 0.507
ATiSE 31.46M 0.252 0.148 0.288 0.462
TASTER 5.04M 0.327 0.239 - 0.514
TGeomE+ - 0.333 0.232 0.361 0.546
UniHR 3.68M 0.333 0.240 0.367 0.527
Table 4: Ablation studies on the HiSL module for triple prediction. Best results are boldfaced.
Variant FBH FBHE DBHE
MR MRR Hits@10 MR MRR Hits@10 MR MRR Hits@10
w/o initial $\textbf{h}_{f}$ 5.22 0.909 0.980 5.86 0.767 0.885 2.56 0.794 0.978
w/o $\textbf{W}_{r}$ 3.06 0.944 0.992 5.98 0.792 0.885 2.30 0.850 0.976
w/o intra-fact MP 3.97 0.897 0.972 5.26 0.754 0.883 2.02 0.842 0.983
w/o $\omega_{\tau(r)}$ 2.70 0.934 0.992 5.46 0.782 0.888 2.69 0.810 0.973
w/o $\textbf{W}_{\lambda(r)}$ 2.50 0.941 0.992 5.69 0.778 0.889 2.37 0.810 0.978
w/o inter-fact MP 2.61 0.913 0.991 4.56 0.776 0.887 2.11 0.827 0.986
UniHR 2.46 0.946 0.993 5.20 0.793 0.890 1.90 0.862 0.987
Table 5: Results of separate training and joint training on the HKG and TKG dataset, where identical entities and relations share the same embeddings. WikiPeople- and wikidata12k- represent the filtered test sets.
Model WikiPeople- wikidata12k-
subject/object all entities subject/object
MR MRR Hits@1 Hits@3 Hits@10 MR MRR Hits@1 Hits@3 Hits@10 MR MRR Hits@1 Hits@3 Hits@10
UniHR 835.8 0.486 0.412 0.528 0.617 829.0 0.488 0.414 0.531 0.620 818.7 0.314 0.220 0.345 0.509
UniHRJoint 692.7 0.488 0.409 0.533 0.629 686.5 0.490 0.414 0.536 0.632 489.5 0.315 0.222 0.346 0.498
Link prediction on HKG.
We compare our method with previous methods on the WD50K and WikiPeople datasets shown in Table 1. We can observe that the models based on transformers (i.e., StarE, GRAN, HyNT) demonstrate significantly better performance compared to other models. We attribute this to the transformer’s superior ability to capture long distance dependencies within H-Facts. Among these methods, it can be seen that our proposed UniHR achieves state-of-the-art results, which means our method effectively captures global structural information. Compared to the GNN-based method StarE, we achieve improvements of 3.9 points (12.6%) in MRR, 4.4 points (18.8%) in Hits@1 and 3.0 points (6.6%) in Hits@10 on WD50K. This indicates that the performance of StarE’s customized graph neural network is limited by its inability to flexibly capture key-value pair information. Moreover, since the embeddings for newly added fact nodes and relation nodes are computed from atomic facts, so our training parameters do not significantly increase.
Link prediction on NKG.
Our experiments on the NKGs consist of two tasks: base link prediction and triple prediction. Unlike previous quaternion-based methods Chung and Whang (2023); Xiong et al. (2024), UniHR significantly reduces the number of required training parameters. From the results in Table 2, we can see that our proposed UniHR obtains competitive results as the first method to capture global structural information of NKGs. For base link prediction task, UniHR achieves considerable improvements on all datasets. Of particular note, the MRR of FBHE/FBH and DBHE increases by 8.1% and 3.8%, respectively.
For triple prediction, we perform best on FBH and DBHE datasets, especially obtaining an improvement of 2.4 points in MRR on FBH, and achieve the second-best performance on FBHE, which suggests that structural information is also valuable for NKG and UniHR can effectively capture the heterogeneity of NKG to enhance node embeddings. In particular, as a unified method, we do not use the customized decoder for triples, while previous state-of-the-art methods do. We will further illustrate the effectiveness of UniHR with other decoders in Appendix C.
Link prediction on TKG.
As shown in Table 3, we achieve competitive results on the wikidata12k, even surpassing TGeomE+ by 3.4% on Hits@1 and 1.7% on Hits@3. However, existing temporal knowledge graph embedding methods, such as TGeomE+, often employ time-aware decoders, which are challenging to generalize to other types of KGs. In contrast, our approach efficiently encodes timestamps as atomic nodes only during initialization and learns temporal information through message passing on graph structure, demonstrating that graph structure information is also beneficial for temporal knowledge graphs, highlighting the effectiveness of our UniHR.
5.3 Further Analysis
Ablation study on HiSL.
We conduct ablation experiments on triple prediction, the most relevant task to fact nodes, using three NKG datasets. As shown in Table 4, all variant models with certain modules or parameters removed exhibit a decrease in performance. We can conclude that intra-fact and inter-fact message passing modules both play crucial roles, allowing UniHR to more fully represent the current fact node with enhanced structural information. We also change the initialization of fact node embeddings to learnable embeddings (w/o initial $\textbf{h}_{f}$) in HiSL. There is a performance decrease, indicating that initializing fact node representations is essential. It highlights key information in the facts, mitigating the noise introduced by excessive neighbors.
Potential of Joint Learning on Different Types of KGs.
We suppose that unified representation makes it possible to develop pre-trained models that integrate multiple types of KGs. To explore its potential, we conduct joint learning on different types of KGs. Therefore, we construct a hybrid dataset called wikimix which includes two subsets of Wikidata Vrandečić and Krötzsch (2014), namely HKG dataset WikiPeople and TKG dataset wikidata12k, which encompass 3547 identical entities and 18 identical relations. Due to the different types of facts, there are no identical facts in these two subsets. To further prevent data leakage, we filter out 537 entries from the wikidata12k test set whose main triples appear in the H-Facts of WikiPeople train set, and 384 entries from the WikiPeople test set whose main triples appear in wikidata12k train set. Statics of wikimix are given in Appendix D.
From the results in Table 5, it is evident that joint learning outperforms separate learning across most metrics. Notably, there are improvements of 17.1% and 40.2% in MR metric on wikipeople- and wikidata12k- datasets, respectively. This indicates that more complex structural interactions and diverse types of training data are beneficial. Moreover, our UniHR demonstrates good scalability and effectiveness in integrating multiple types of KGs.
6 Conclusion
In this paper, we propose UniHR, a unified hierarchical knowledge graph representation learning framework consisting of a Hierarchical Data Representation (HiDR) module and a Hierarchical Structure Learning (HiSL) module. The HiDR form unifies the hyper-relational facts, nested facts and temporal facts into the form of triples, overcoming the limitations of customized encoders for different forms of facts. Moreover, HiSL captures local semantic information within facts and global structural information between facts. Our approach achieves the best or competitive performance on link prediction tasks across three types of KGs.
Limitations
The limitations of our paper are summarized as follows:
Our UniHR
In this paper, our UniHR framework focuses on link prediction tasks under transductive settings with a single modality. In the future, we will investigate how to generalize our HiDR form to more complex tasks such as inductive reasoning Teru et al. (2020) and multi-modal scenarios Zhang et al. (2024), etc.
Joint Learning on Different Types of KGs
Constrained by computational resources, our analysis of the potential for joint training across multiple types of knowledge graphs focuse only on HKG and TKG. We believe the unification of knowledge graph representation learning methods is a developing trend that makes it possible to develop unified pre-trained models based on multiple types of KGs. In the future, we aim to explore joint training across more types of KG to demonstrate the advantages of integrating multi-type KG data.
Ethics Statement
In this paper, we explore the unified knowledge graph link prediction problem, aiming to complete various types of knowledge graphs using a unified model with deep learning techniques. Our training and evaluation are based on publicly available and widely used datasets of different types of knowledge graphs. Therefore, we believe this does not violate any ethics.
Appendix A Related Works
A.1 Link Prediction Methods for Triple-based KGs
Most existing techniques in KG representation learning are proposed for triple-based KGs. Among these techniques, knowledge graph embedding (KGE) models Bordes et al. (2013); Sun et al. (2018) have received extensive attention due to their effectiveness and simplicity. The idea is to project entities and relations in the KG to low-dimensional vector spaces, utilizing KGE scoring functions to measure the plausibility of triples in the embedding space. Typical methods include TransE Bordes et al. (2013), RotatE Sun et al. (2018), and ConvE Dettmers et al. (2018).
Depending on the KGE model alone has limitation of capturing complex graph structures, whereas augmenting global structural information with a graph neural network (GNN) Vashishth et al. (2019); Nathani et al. (2019); Xu et al. (2023b) proves to be an effective approach for enhancement. The paradigm of combining GNN as encoder with KGE scoring function as decoder helps to enhance the performance of KGE scoring function. These GNN methods design elaborate message passing mechanisms to capture the global structural features. Typically, CompGCN Vashishth et al. (2019) aggregates the joint embedding of entities and relations in the neighborhood via a parameter-efficient way and MA-GNN Xu et al. (2023b) learns global-local structural information based on multi-attention. These methods achieve impressive results on triple-based KGs but are hard to generalize to beyond-triple KGs.
Appendix B Results on JF17K
Table 6 shows the experimental results on JF17K. Due to the absence of a validation set in the JF17K dataset and the different ways of dividing the dataset across various baselines, we adopt the results reported in the original paper. Consistent with previous experiments on hyper-relational knowledge graphs, we also achieve state-of-the-art performance on JF17K among all baselines. In particular, we achieved 1.7 (2.9%) points in MRR and 1.5 (2.9%) points in Hits@1 compared to the method StarE which also utilises a graph neural network encoding and a simple transformer decoding, indicating that our hierarchical GNN HiSL could better capture the structure of hyper-relational facts.
Table 6: Link prediction on JF17K. All results of baselines are taken from the original paper. Best results are in bold.
Model JF17K
subject/object all entities
MRR Hits@1 Hits@10 MRR Hits@1 Hits@10
m-TransH 0.206 0.206 0.462 - - -
NaLP 0.221 0.165 0.331 0.366 0.290 0.516
HINGE 0.431 0.342 0.611 0.517 0.436 0.675
NeuInfer 0.449 0.361 0.624 0.473 0.397 0.618
StarE 0.574 0.496 0.725 - - -
HyTransformer 0.582 0.501 0.742 - - -
UniHR 0.591 0.511 0.745 0.621 0.545 0.768
Appendix C Decoder Analysis
To explore the effectiveness of our UniHR encoding further, we pair UniHR with different decoders and evaluated them on triple prediction task. In addition to the previously mentioned unified framework UniHR + Transformer, we also experiment on UniHR + ConvE with two scoring strategies. The ConvE Dettmers et al. (2018) is the decoder customized for triples and its scoring function is $vec\left(\sigma\left(\left[\mathbf{\tilde{h}}_{h};\mathbf{\tilde{e}}_{r}\right]*\psi\right)\right)$, where $\mathbf{\tilde{h}}_{h}$ and $\mathbf{\tilde{e}}_{r}$ represent reshaped 2D embeddings of head entity $h$ and relation $r$, and * is a convolution operator. The $vec\left(\cdot\right)$ and $\psi$ are denoted as the vectorization function and a set of convolution kernels.
Due to our special representation, there exists two scoring methods for atomic triples, thus we present the base link prediction results separately for each scoring method. The $s_{f}$ represents scoring triples $\left(f,has\;head\;entity,h\right)$and $\left(f,has\;tail\;entity,t\right)$, and $s_{t}$ represents scoring $\left(h,r,t\right)$. The performance of base link prediction is shown in Table 7. Notably, FBH and FBHE share identical atomic facts, resulting in the same performance. It can be observed that regardless of the scoring method employed, we both achieve competitive performance, especially with scoring $\left(h,r,t\right)$ on FBH and scoring $\left(f,has\;head\;entity,h\right)\,\left(f,has\;tail\;entity,t\right)$ on DBHE. We attribute the differences in performance under different scoring methods to dataset characteristics. The DBHE dataset is relatively smaller, and scoring method $s_{f}$ effectively alleviates overfitting problem. Conversely, for larger datasets FBH, scoring based on $\left(h,r,t\right)$ minimizes information loss.
Table 8 shows the results of triple prediction on three benchmark datasets. Among all baselines, Quate, Bique, Neural-LP, Drum, and AnyBURL struggle to model the mapping relationship between atomic facts and nested facts. Furthermore, prior works Chung and Whang (2023) do not guarantee that all atomic facts in the nested fact test set are present in the training set as entities, which shifts the problem from a transductive setting to an inductive setting, leading to significant performance gaps between these baselines. On most metrics, our method outperforms BiVE and NestE which are specifically modeled for nested facts. Notably, NestE fully preserves the semantics of atomic facts. However, on the FBHE dataset, UniHR + ConvE achieves an improvement of 0.58 (6.4%) points in MRR and 0.24 (2.4%) points in Hits@10 compared to the state-of-the-art model NestE and the second-best performance after UniHR + Transformer on the FBH and DBHE datasets, demonstrating UniHR’s powerful graph structure encoding capabilities. We also carry out ablation experiments on UniHR + ConvE as shown in Table 8. Performance declines are observed after removing any part of the HiSL module, showing the significance of HiSL for hierarchical encoding.
Table 7: Base link prediction on FBHE, FBH and DBHE. All baselines’ results are taken from Xiong et al. (2024). The best results among all models are written bold, while the second are underlined. The $s_{f}$ and $s_{h}$ denote $\left(f,has\;head\;entity,h\right)\ \left(f,has\;tail\;entity,t\right)$ and $\left(h,r,t\right)$ two types of scoring method respectively. For BiVE Chung and Whang (2023) and NestE Xiong et al. (2024), we pick their variants with best performance.
Model FBHE/FBH DBHE
MRR Hits@10 MRR Hits@10
QuatE 0.354 0.581 0.264 0.440
BiQUE 0.356 0.583 0.274 0.446
Neural-LP 0.315 0.486 0.233 0.357
DRUM 0.317 0.490 0.237 0.359
AnyBURL 0.310 0.526 0.220 0.364
BiVE 0.370 0.607 0.274 0.422
NestE 0.371 0.608 0.289 0.443
UniHR + ConvE $s_{h}$ 0.397 0.622 0.289 0.443
UniHR + ConvE $s_{f}$ 0.375 0.596 0.307 0.471
UniHR + Transformer 0.401 0.619 0.300 0.455
Table 8: Triple prediction on FBHE, FBH and DBHE. All baselines’ results are taken from Xiong et al. (2024). The best results among all models are written bold. For BiVE Chung and Whang (2023) and NestE Xiong et al. (2024), we pick their variants with best performance.
Model FBH FBHE DBHE
MR MRR Hits@10 MR MRR Hits@10 MR MRR Hits@10
QuatE Zhang et al. (2019) 145603.8 0.103 0.114 94684.4 0.101 0.209 26485.0 0.157 0.179
BiQUE Guo and Kok (2021) 81687.5 0.104 0.115 61015.2 0.135 0.205 19079.4 0.163 0.185
Neural-LP Yang et al. (2017) 115016.6 0.070 0.073 90000.4 0.238 0.274 21130.5 0.170 0.209
DRUM Sadeghian et al. (2019) 115016.6 0.069 0.073 90000.3 0.261 0.274 21130.5 0.166 0.209
AnyBURL Meilicke et al. (2019) 108079.6 0.096 0.108 83136.8 0.191 0.252 20530.8 0.177 0.214
BiVE Chung and Whang (2023) 6.20 0.855 0.941 8.35 0.711 0.866 3.63 0.687 0.958
NestE Xiong et al. (2024) 3.34 0.922 0.982 3.05 0.851 0.962 2.07 0.862 0.984
UniHR + Transformer 2.46 0.946 0.993 5.20 0.793 0.890 1.90 0.862 0.987
UniHR + ConvE 3.00 0.900 0.983 6.27 0.909 0.986 2.06 0.876 0.978
UniHR + ConvE w/o $\textbf{h}_{f}$ 4.39 0.887 0.979 10.10 0.865 0.970 2.76 0.798 0.961
UniHR + ConvE w/o intra-fact 6.54 0.859 0.959 18.10 0.871 0.968 5.82 0.665 0.900
UniHR + ConvE w/o inter-fact 12.56 0.864 0.961 20.56 0.864 0.966 10.75 0.764 0.951
Table 9: The statistics of diverse knowledge graphs dataset, where “with Q(%)" and “Arity" column respectively denote the number of facts with auxiliary key-value pairs and the range of arity of hyper-relational facts, the “with AF(%)" column denotes the number of atomic facts in nested facts.
Dataset Atomic Fact Entities Relations Train Valid Test with Q(%) Arity Nested Fact Nested Relation with AF(%) Period
Hyper-relational Knowledge Graph
WikiPeople 369866 34825 178 294439 37715 37712 9482(2.6%) 2-7 - - - -
WD50K 236507 47155 531 166435 23913 46159 32167(13.6%) 2-67 - - - -
JF17K 100947 28645 501 76379 - 24568 46320(45.9%) 2-6 - - - -
Nested Factual Knowledge Graph
FBH 310116 14541 237 248094 31011 31011 - - 27062 6 33157 -
FBHE 310116 14541 237 248094 31011 31011 - - 34941 10 33719 -
DBHE 68296 12440 87 54636 6830 6830 - - 6717 8 8206 -
Temporal Knowledge Graph
wikidata12k 40621 12554 24 32497 4062 4062 - - - - - 19-2020
Multiple types of Knowledge Graph
wikimix 409566 43832 184 326936 41777 3525(TKG)/37328(HKG) 9098(2.2%) 2-7 - - - 19-2020
Table 10: The major hyperparameters of our approach for all link prediction tasks.
Hyperparameter WikiPeople WD50K wikidata12k FBHEbase FBHbase DBHEbase FBHEtriple FBHtriple DBHEtriple
batch_size 2048 2048 2048 2048 2048 2048 2048 2048 2048
embedding dim 200 200 200 200 200 200 200 200 200
hidden dim 200 200 200 200 200 200 200 200 200
GNN_layer 2 2 2 2 2 2 2 2 2
GNN_intra-fact heads 4 4 4 4 4 4 4 4 4
GNN_intra-fact dropout 0.1 0.1 0.1 0.3 0.1 0.3 0.2 0.1 0.1
GNN_inter-fact activation tanh tanh tanh tanh tanh tanh tanh tanh tanh
GNN_dropout 0.1 0.1 0.1 0.3 0.1 0.3 0.2 0.1 0.1
transformer layers 2 2 2 2 2 2 2 2 2
transformer heads 4 4 4 4 4 4 4 4 4
transfomer activation gelu gelu gelu gelu gelu gelu gelu gelu gelu
decoder dropout 0.1 0.1 0.1 0.1 0.1 0.1 0.1 0.1 0.1
soft label for entity 0.2 0.2 0.4 0.2 0.2 0.3 0.2 0.2 0.2
soft label for relation 0.1 0.1 0.3 0.2 0.2 0.3 0.2 0.2 0.2
weight_decay 0.01 0.01 0.01 0.01 0.01 0.01 0.01 0.01 0.01
learning rate 5e-4 5e-4 5e-4 5e-4 5e-4 5e-4 5e-4 5e-4 5e-4
Algorithm 1 Message passing process of HiSL
Input: $\mathcal{G}^{HiDR}=\{\mathcal{V}_{a}\cup\mathcal{V}_{r}\cup\mathcal{V}_{f},\mathcal{R}_{a}\cup\mathcal{R}_{n}\cup\mathcal{R}_{c},\mathcal{F}_{a}\cup\mathcal{F}_{c}\cup\mathcal{F}_{n}\}$; the number of encoder layers $L$.
Output: The node embedding matrix $\mathbf{H}^{L}$; the edge embedding matrix $\mathbf{E}^{L}$.
1:  Initialize the embedding matrix of atomic nodes $\mathbf{H}_{a}^{0}\in\mathbb{R}^{\left|\mathcal{V}_{a}\right|\times d}$.
2:  Initialize the embedding matrix of three types of relations $\mathbf{E}^{0}=\{\mathbf{E}_{a}^{0},\mathbf{E}_{c}^{0},\mathbf{E}_{n}^{0}\}\in\mathbb{R}^{\left|\mathcal{R}_{a}\cup\mathcal{R}_{c}\cup\mathcal{R}_{n}\right|\times d}$.
3:  Initialize the embedding matrix of relation nodes $\mathbf{H}_{r}^{0}\leftarrow\mathbf{E}_{a}^{0}\cdot\mathbf{W}_{r}\in\mathbb{R}^{\left|\mathcal{V}_{r}\right|\times d}$.
4:  for  $h,r,t\in\mathcal{F}_{a}$ do
5:     Initialize the embedding of fact nodes $\mathbf{h}_{f}^{0}\leftarrow f_{m}([\mathbf{h}_{h};\mathbf{h}_{r};\mathbf{h}_{t}])$. $\triangleright$ Representation Initialization
6:  end for
7:  for  $l\leftarrow 1$ to $L$ do
8:     for $i\in\mathcal{V}$ do
9:        $\mathcal{M}^{l}_{i}\leftarrow\mathbf{h}_{i}^{l-1}$. $\triangleright$ Intra-fact Message Passing
10:        for  $i,r,j\in\mathcal{F}_{c}$ do
11:           $\alpha_{i,j}^{l-1}\leftarrow\frac{\exp\left(\mathbf{W}^{l-1}\left(\sigma\left(\mathbf{W}_{in}^{l-1}\mathbf{h}_{i}^{l-1}+\mathbf{W}_{out}^{l-1}\mathbf{h}_{j}^{l-1}\right)\right)\right)}{\sum_{j^{\prime}\in\mathcal{N}_{i}}{\exp\left(\mathbf{W}^{l-1}\left(\sigma\left(\mathbf{W}_{in}^{l-1}\mathbf{h}_{i}^{l-1}+\mathbf{W}_{out}^{l-1}\mathbf{h}_{j^{\prime}}^{l-1}\right)\right)\right)}}$.
12:           $\mathcal{M}^{l}_{i}\leftarrow\{\alpha_{i,j}^{l-1}\cdot\mathbf{W}_{out}^{l-1}\mathbf{h}_{j}^{l-1}\}\cup\mathcal{M}^{l}_{i}$.
13:        end for
14:     end for
15:     for $i\in\mathcal{V}$ do
[…]
17:     end $\triangleright$ Inter-fact Message Passing
18:     for $i\in\mathcal{V}$ do
19:        $\mathcal{M}^{l}_{i}\leftarrow\{\mathbf{W}_{self}^{{l-1}}\mathbf{h}_{i}^{{l-1}}\}$.
20:        for  $r,j\in\mathcal{N}_{i}$ do
21:           $\mathbf{m}^{l-1}_{(i,r,j)}\leftarrow\sigma\left(\omega_{\tau\left(r\right)}^{l-1}\right)\mathbf{W}_{{\lambda}\left(r\right)}^{{l-1}}\phi\left(\mathbf{h}_{{j}}^{{l-1}},\mathbf{e}_{{r}}^{{l-1}}\right)$.
22:           $\mathcal{M}^{l}_{i}\leftarrow\{\mathbf{m}^{l-1}_{(i,r,j)}\}\cup\mathcal{M}^{l}_{i}$.
23:        end for
24:     end for
25:     for $i\in\mathcal{V}$ do
[…]
27:     end for
28:     $\mathbf{E}^{l}\leftarrow\mathbf{W}_{rel}^{l-1}\mathbf{E}^{l-1}$.
29:  end for
30:  return  $\mathbf{H}^{L},\mathbf{E}^{L}$
Appendix D Datasets Statistics
Table 9 shows the details of the three hyper-relational knowledge graph benchmark datasets: WikiPeople, WD50K, JF17K, three nested factual knowledge graph benchmark datasets: FBH, FBHE, DBHE,and the temporal knowledge graph benchmark dataset wikidata12k. Among them, WikiPeople is a dataset derived from Wikidata Vrandečić and Krötzsch (2014) concerning entities type “human”. WikiPeople filter out the elements which have at least 30 mentions as key-value pairs. WD50K is a high-quality dataset extracting from Wikidata statements and avoiding the potential data leakage which allows triple-based models to memorize main fact in the H-Facts of test set. The “with Q(%)" column in Table 9 denote the number of facts with auxiliary key-value pairs and the “Arity" column denote range of the number of entities in hyper-relational facts. The nested factual knowledge graph datasets FBH and FBHE Chung and Whang (2023) are constructed based on FB15k237 Toutanova and Chen (2015) from Freebase Bollacker et al. (2008) while DBHE is based on DB15K Liu et al. (2019) from DBpedia Lehmann et al. (2015). FBH contains nested facts that can be only inferred inside the atomic facts, while FBHE and DBHE contain externally-sourced nested relation crawling from Wikipedia articles, e.g., NextAlmaMater and SucceededBy. Temporal knowledge graph dataset wikidata12K is also a subset of Wikidata Vrandečić and Krötzsch (2014), which represents the time information $\tau\in\mathcal{T}$ as time intervals.
Appendix E Pseudo-code of HiSL
The pseudo code of HiSL is outlined in Algorithm 1.
Appendix F Hyperparameter Settings
Here, we show the hyperparameter details for each link prediction task. To be specific, we tune the learning rate using the range $\left\{0.0001,0.0005,0.001\right\}$, the embedding dim using the range $\left\{50,100,200,400\right\}$, the GNN layer using the range $\left\{1,2,3\right\}$ and dropout using the range $\left\{0.1,0.2,0.3,0.4\right\}$. Additionally, we use smoothing label in the training phase from range $\left\{0.1,0.2,0.3\right\}$. The best hyperparameters obtained from the experiments are presented in Table 10.

### v3

@@
 Beyond-triple fact representations including hyper-relational facts with auxiliary key-value pairs, temporal facts with additional timestamps, and nested facts implying relationships between facts, are gaining significant attention.
-However, existing link prediction models are usually designed for one specific type of facts, making it difficult to generalize to other fact representations.
+However, constrained by complex fact representation forms, existing link prediction models for beyond-triple facts have difficulty achieving hierarchical fact modeling and generalizing the modules for one specific facts to other fact types.
 To overcome this limitation, we propose a Unified Hierarchical Representation learning framework (UniHR) for unified knowledge graph link prediction.
@@
 Then HiSL incorporates intra-fact and inter-fact message passing, focusing on enhancing the semantic information within individual facts and enriching the structural information between facts.
-Experimental results across 7 datasets from 3 types of KGs demonstrate that our UniHR outperforms baselines designed for one specific kind of KG, indicating strong generalization capability of HiDR form and the effectiveness of HiSL module.
+Empirical results demonstrate the effectiveness of UniHR and highlight the strong potential of unified representations.
 Code and data are available at https://github.com/Lza12a/UniHR.
-UniHR: Hierarchical Representation Learning for Unified Knowledge Graph Link Prediction
-Zhiqiang Liu1,3, Mingyang Chen2,3, Yin Hua1,3, Zhuo Chen2,3, Ziqi Liu4, Lei Liang4, Huajun Chen2,3, Wen Zhang1,3thanks: Corresponding Author.
-1School of Software Technology, Zhejiang University 2College of Computer Science and Technology, Zhejiang University 3ZJU-Ant Group Joint Lab of Knowledge Graph 4Ant Group {zhiqiangliu,zhang.wen}@zju.edu.cn
+Keywords:
+Knowledge graph representation learning Beyond-triple knowledge graph Link prediction.
 1 Introduction
-Large-scale knowledge graphs (KGs) such as WordNet Miller (1995), Freebase Bollacker et al.
-(2008), and Wikidata Vrandečić and Krötzsch (2014) have been widely applied in many areas like question answering Kaiser et al.
-(2021), recommendation systems Guo et al.
-(2020), and natural language processing Annervaz et al.
-(2018).
+Large-scale knowledge graphs (KGs) such as Freebase [3] and Wikidata [33] have been widely applied in many areas like question answering [17] and natural language processing [2].
 However, the presence of missing facts within these KGs inevitably limit their applications.
 Therefore, the link prediction task has been introduced to predict missing elements within factual data.
-Current link prediction methods mainly focus on facts in the form of triple $\left(head\,entity,relation,tail\,entity\right)$.
-Figure 1: A special KG consists of triple-based fact, hyper-relational fact, nested fact and temporal fact.
+Current link prediction methods mainly focus on the fact in the form of triple e.g., $(head\,entity,relation,$
+$tail\,entity)$.
+Figure 1: Comparison between the existing link prediction methods for specific beyond-triple knowledge graphs and our unified methods.
 Despite the simplicity and unity of triple-based representation, it is difficult to adequately express complex facts, such as “Oppenheimer is educated at Harvard University for a bachelor degree in chemistry" shown in Figure 1.
-Therefore, existing researches Wang et al.
-(2021); Xiong et al.
-(2024); Xu et al.
-(2019) contribute to focusing on semantically richer facts.
+Therefore, existing researches [34, 37, 38] contribute to focusing on semantically richer facts.
 Figure 1 illustrates three specific types of facts: hyper-relational fact ((Oppenheimer, educated at, Harvard University), degree: bachelor, major: chemistry), temporal fact (Oppenheimer, honored with, Fermi Prize, 1963), nested fact ((Oppenheimer, born in, New York), imply, (Oppenheimer, nationality, The United States)).
 These forms of facts allow for expression of complex semantics and revelation of relationships between facts, extending beyond the triple-based representation.
-Thus in recent years, Hyper-relational KGs (HKG) Chung et al.
-(2023), Temporal KGs (TKG) Xu et al.
-(2023a), and Nested factual KGs (NKG) Xiong et al.
-(2024) attract wide research interests.
-Recent studies have demonstrated the effectiveness of various embedding strategies for these beyond-triple representations Xiong et al.
-(2023).
-However, these methods are usually designed for specific representation forms, e.g., StarE Galkin et al.
-(2020) customizes graph neural network to implement message passing on hyper-relational facts, For nested factual KGs, BiVE Chung and Whang (2023) connects two levels of facts throgh a simple linear layer.
-In addition, GeomE+ Xu et al.
-(2023a) et al.
-temporal KG embedding methods contain time-aware scoring functions to adapt timestamps.
-Although these methods perform well on specific type of facts, it is evident that such customized methods are difficult to generalize to other types of KGs.
-Therefore, establishing a unified representation learning method for multiple types of KGs is worth to investigate.
+Thus in recent years, Hyper-relational KGs (HKG) [6], Temporal KGs (TKG) [40], and Nested factual KGs (NKG) [37] attract wide research interests.
+Recent studies have demonstrated the effectiveness of various embedding strategies for these beyond-triple representations [36].
+However, these methods are hard to achieve hierarchical fact modeling and usually designed for specific representation forms, e.g., StarE [12] customizes graph neural network to enhance inter-fact structure information between hyper-relational facts.
+For NKGs, BiVE [7] bridges intra-fact semantics between two levels of facts through a simple MLP.
+Moreover, GeomE+ [40] et al.
+TKG methods employ time-aware scoring functions to capture intra-fact collaborative information between factual semantics and temporal information in the timestamp.
+Although these methods perform well on specific types of facts, it is evident that such customized methods are difficult to generalize to other types of KGs for intra-fact structure and inter-fact semantic enhancements.
+Therefore, establishing a unified hierarchical representation learning method for multiple types of KGs is worth to investigate.
 To overcome the challenges mentioned above, we propose a Unified Hierarchical Representation learning method (UniHR), which includes a Hierarchical Data Representation (HiDR) module and a Hierarchical Structure Learning (HiSL) module as the graph encoder.
 HiDR module standardizes hyper-relational facts, nested factual facts, and temporal facts into the form of triples without loss of information.
-Furthermore, HiSL module captures local semantic information during intra-fact message passing and then utilizes inter-fact message passing to enrich the global view of nodes to obtain better node embeddings based on HiDR form.
+Furthermore, HiSL module captures local semantic information during intra-fact message passing and then utilizes inter-fact message passing to enrich the global structure information to obtain better node embeddings based on HiDR form.
 Finally, the updated embeddings are fed into decoders for link prediction.
-Experimental results demonstrate that our UniHR achieves state-of-the-art performance on HKG and NKG datasets, and competitive performance on TKG datasets, revealing strong generalization capability of HiDR form and effectiveness of HiSL module.
-Our contributions can be summarized as follows:
+Empirical results demonstrate the UniHR achieves state-of-the-art or competitive performance on most metrics and highlight the strong potential of unified representations.
+Our contributions can be summarized as follows.
   • 1.
@@
   • 2.
-    We propose the first unified KG representation learning framework UniHR, across different types of KGs, including a hierarchical data representation module and a hierarchical structure learning module.
+    To our knowledge, we propose the first unified KG representation learning framework UniHR, across different types of KGs, including a hierarchical data representation module and a hierarchical structure learning module.
   • 3.
-    We conduct link prediction experiments on 7 datasets across 3 types of KGs.
+    We conduct link prediction experiments on 9 datasets across 5 types of KGs.
 Compared to methods designed for one kind of KG, UniHR achieves the best or competitive results, verifying strong generalization capability.
@@
 In this section, we introduce the definition of four types of existing knowledge graphs (KGs): triple-based KG, hyper-relational KG, nested factual KG and temporal KG, along with link prediction tasks on these types of KGs.
-Triple-based Knowledge Graph.
-A common triple-based KG $\mathcal{G}_{KG}\,\text{=}\,\{\mathcal{V},\mathcal{R},\mathcal{F}\}$ represents facts as triples, denoted as =$\mathcal{F}\,\text{=}\{\left(h,r,t\right)|h,t\in\mathcal{V},$ $r\in\mathcal{R}\}$, where $\mathcal{V}$ is the set of entities and $\mathcal{R}$ is the set of relations.
 Link Prediction on Triple-based KG.
+A triple-based KG $\mathcal{G}_{KG}\,\text{=}\,\{\mathcal{V},\mathcal{R},\mathcal{F}\}$ represents facts as triples, denoted as =$\mathcal{F}\,\text{=}\{\left(h,r,t\right)|h,t\in\mathcal{V},$ $r\in\mathcal{R}\}$, where $\mathcal{V}$ is the set of entities and $\mathcal{R}$ is the set of relations.
 The link prediction on triple-based KGs involves answering a query $\left(h,r,?\right)$ or (?$,r,t\right)$, where the missing element ‘????’ is an entity in $\mathcal{V}$.
-Hyper-relational Knowledge Graph.
+Link Prediction on Hyper-relational KG.
 A hyper-relational KG (HKG) $\mathcal{G}_{HKG}\,\text{=}\,\{\mathcal{V},\mathcal{R},\mathcal{F}\}$ consists of hyper-relational facts, abbreviated as H-Facts, denoted as =$\mathcal{F}\,\text{=}\{((h,r,t),\{(k_{i}\text{:}\,v_{i})\}_{i=1}^{m})$ $|\,h,t,v_{i}\in\mathcal{V},r,k_{i}\in\mathcal{R}\}$.
 Typically, we refer to $\left(h,r,t\right)$ as the main triple in the H-Fact and $\left\{\left(k_{i}\text{:}\,v_{i}\right)\right\}_{i=1}^{m}$ as $m$ auxiliary key-value pairs.
-Link Prediction on Hyper-relational KG.
 Similar to link prediction on triple-based KGs, the link prediction on HKGs aims to predict entities in the main triple or the key-value pairs.
 Symbolically, the aim is to predict the missing element, denoted as ‘?’ for queries $((h,r,t),(k_{1}\text{:}\,v_{1}),\ldots(k_{i}\text{:}\,?)),$ ((?$,r,t),\{(k_{i}\text{:}v_{i})\}_{i=1}^{m})$ or $((h,r,?),\{(k_{i}\text{:}v_{i})\}_{i=1}^{m})$.
-Nested Factual Knowledge Graph.
+Link Prediction on Nested Factual KG.
 A nested factual KG (NKG) can be represented as $\mathcal{G}_{NKG}\,\text{=}\,\{\mathcal{V},\mathcal{R},\mathcal{F},\hat{\mathcal{R}},\hat{\mathcal{F}}\}$, which is composed of two levels of facts, called atomic facts and nested facts.
@@
 ^=$\hat{\mathcal{F}}\,\text{=}\,\{\left(\mathcal{F}_{i},\hat{r},\mathcal{F}_{j}\right)|$ $\mathcal{F}_{i},\mathcal{F}_{j}\in\mathcal{F},\hat{r}\in\hat{\mathcal{R}}\}$ is the set of nested facts, where $\hat{\mathcal{R}}$ is the set of nested relations.
-Link Prediction on Nested Factual KG.
 The link prediction on the NKGs is performed on the atomic facts or nested facts.
@@
 For triple prediction, given a query (?$,\hat{r},\mathcal{F}_{j}\right)$ or $\left(\mathcal{F}_{i},\hat{r},?\right)$, the aim is to predict the missing atomic fact ‘????’ from $\mathcal{F}$.
-Temporal Knowledge Graph.
-A temporal KG (TKG) $\mathcal{G}_{TKG}\,\text{=}\,\{\mathcal{V},\mathcal{R},\mathcal{F},\mathcal{T}\}$ is composed of quadruple-based facts, which can be represented as =$\mathcal{F}\,\text{=}\,\{(h,r,t,[\tau_{b},\tau_{e}])|h,t\in\mathcal{V},r\in\mathcal{R},\tau_{b},\tau_{e}\in\mathcal{T}\}$, where $\tau_{b}$ is the begin time, $\tau_{e}$ is the end time, $\mathcal{V}$ is the set of entities, $\mathcal{R}$ is the set of relations and $\mathcal{T}$ is the set of timestamps.
 Link Prediction on Temporal KG.
+A temporal KG (TKG) $\mathcal{G}_{TKG}\,\text{=}\,\{\mathcal{V},\mathcal{R},\mathcal{F},\mathcal{T}\}$ is composed of quadruple-based facts, which can be represented as =$\mathcal{F}\,\text{=}\,\{(h,r,t,$
+$[\tau_{b},\tau_{e}])|h,t\in\mathcal{V},r\in\mathcal{R},\tau_{b},\tau_{e}\in\mathcal{T}\}$, where $\tau_{b}$ is the begin time, $\tau_{e}$ is the end time, $\mathcal{V}$ is the set of entities, $\mathcal{R}$ is the set of relations and $\mathcal{T}$ is the set of timestamps.
 The link prediction on TKGs aims to predict missing entities ‘????’ in $\mathcal{V}$ for two types of queries (?$,r,t,[\tau_{b},\tau_{e}]\right)$ or $\left(h,r,?,[\tau_{b},\tau_{e}]\right)$.
@@
 Link Prediction on Hyper-relational Knowledge Graph.
-Earlier HKG representation learning methods e.g., m-TransH Wen et al.
-(2016), RAE Zhang et al.
-(2018) have generalized the triple-based approach to HKG and loosely represent the combinations of key-value pairs.
+Earlier HKG representation learning methods e.g., m-TransH [35], RAE [44] have generalized the triple-based approach to HKG and loosely represent the combinations of key-value pairs.
 Galkin et al.
-customize StarE Galkin et al.
-(2020) based on CompGCN Vashishth et al.
-(2019) for hyper-relational facts to capture the interaction information of key-value pairs with the main triple in the message passing stage, and achieves impressive results, demonstrating that the structural information of the graph in HKGs is also important.
-GRAN Guan et al.
-(2021) introduces edge-aware bias into the vanilla transformer attention Vaswani et al.
-(2017), while HyNT Chung et al.
-(2023) designs a qualifier encoder for HKG.
-They both focus on intra-fact dependencies but ignore the global structural information.
-Due to the existence of its particular key-value pairs on H-Facts, there are many limitations in capturing global structural information through the widely available triple-based GNNs.
+first customize StarE [12] based on CompGCN [31] for H-Facts to capture global structure information between H-Facts in the message passing stage, and achieves impressive results, demonstrating that the structure information of the graph in HKGs is also important.
+GRAN [15] introduces edge-aware bias into the vanilla transformer attention [32], while HyNT [6] designs a qualifier encoder for HKG.
+They both focus on intra-fact semantic dependencies but ignore the global structure information.
+In order to fully capture inter-fact and intra-fact information, HAHE [23] first adopts dual graph attention and edge-aware bias in the transformer attention layer for hierarchical modeling and achieves significant performance improvement.
+We consider the hierarchical attention mechanism of HAHE a good beginning, but it is customized for HKG and is difficult to generalize to other facts.
 Link Prediction on Nested Factual Knowledge Graph.
 Chung et al.
-Chung and Whang (2023) first introduced the concept of nested facts.
-They also propose BiVE which projects atomic facts to fact nodes in the encoding phase via a simple linear layer and scores both atomic facts and nested facts using the quaternion-based KGE scoring functions like QuatE Zhang et al.
-(2019) or BiQUE Guo and Kok (2021).
-Based on BiVE, NestE Xiong et al.
-(2024) represents the fact nodes as a $1\times 3$ embedding matrix and the nested relations as a $3\times 3$ matrix to avoid information loss, embedding them into hyperplanes with different dimensions.
-These methods only capture the semantic information between atomic facts and nested facts while ignoring global structural information.
-Meanwhile, due to the complexity of this representation, common triple-based GNNs have difficulty in message passing between atomic fact and nested fact.
+[7] first introduced nested facts to imply relationships between facts.
+They also propose BiVE which bridges semantics between atomic facts and fact nodes in the encoding phase via a simple MLP and scores both atomic facts and nested facts using the quaternion-based KGE scoring functions like QuatE [45] or BiQUE [16].
+Based on BiVE, NestE [37] represents the fact nodes as a $1\times 3$ embedding matrix and the nested relations as a $3\times 3$ matrix to avoid information loss, embedding them into hyperplanes with different dimensions.
+These methods only capture the intra-fact semantic information between atomic facts and nested facts.
+Meanwhile, due to the complexity of this representation, common triple-based GNNs have difficulty in message passing between atomic fact and nested fact for capturing inter-fact global information.
 Link Prediction on Temporal Knowledge Graph.
-Recent studies in temporal knowledge graph representation learning have focused on enhancing performance by designing special time-aware scoring functions.
-Models such as TTransE Leblay and Chekol (2018), HyTE Dasgupta et al.
-(2018), TeRo Xu et al.
-(2020), and TGeomE+ Xu et al.
-(2023a) incorporate temporal-aware module into the KGE score function in various ways.
-However, existing models for link prediction seldom directly utilize GNNs to perceive time information for enhancing entity and relation embeddings.
+Recent studies in temporal knowledge graph representation learning have focused on designing special time-aware scoring functions for inter-fact semantic modeling.
+Models such as TTransE [21], HyTE [8], TeRo [39], and TGeomE+ [40] incorporate temporal-aware module into the KGE score function in various ways.
+However, existing models seldom regard timestamps as node and directly utilize inter-fact message passing mechanism to perceive time information for enhancing entity and relation embeddings.
 Therefore, we believe it is a direction worth exploring.
@@
 Our workflow can be divided into the following three steps: 1) Given a KG $\mathcal{G}$ of any type, we represent it into $\mathcal{G}^{HiDR}$ under the HiDR form.
-2) The $\mathcal{G}^{HiDR}$ will be encoded by HiSL module with the enhancement of semantic information within individual facts and structural information between facts on the whole graph.
+2) The $\mathcal{G}^{HiDR}$ will be encoded by HiSL module to enhance the semantic information within individual facts and structural information between facts on the whole graph.
 3) In the phase of decoding, the updated embeddings of nodes and edges are fed into transformer decoders to obtain the plausibility score of facts.
-Figure 2: Diverse facts are translated into the HiDR form.
-Blue, light blue, and red circles denote atomic nodes, relation nodes, and fact nodes.
-Black, light blue, and red arrows denote atomic relations, connected relations, and nested relations.
-The triples they connect correspond to atomic facts, connected facts, and nested facts.
+Figure 2: Diverse facts are translated into HiDR form.
 4.1 Hierarchical Data Representation
 To overcome the differences in the representation of multiple types KGs, we introduce a Hierarchical Data Representation module, abbreviated as HiDR.
-Different from labelled RDF representation Ali et al.
-(2022), we constrain “triple" to be considered as the basic units of HiDR form, then HiDR could continuous benefit from the model developments of triple-based KGs, which is the most active area about link prediction over KGs.
+Different from labelled RDF representation [1], we constrain “triple" to be considered as the basic units of HiDR form, then HiDR could continuous benefit from the model developments of triple-based KGs, which is the most active area about link prediction over KGs.
 As shown in Fig.
 2, in order to ensure comprehensive representation of facts, we introduce three hierarchical types of nodes and three connected relations in HiDR.
-Inspired by the nested fact form Xiong et al.
-(2024), we denote original entities within three types of KGs as atomic nodes and complement fact nodes for HKGs and TKGs lacking a designated fact node.
+Firstly, we denote original entities within three types of KGs as atomic nodes and complement fact nodes for HKGs and TKGs lacking a designated fact node.
 To facilitate the interaction between fact nodes and relations explicitly, we incorporate relation nodes into the graph, represented as $e_{r}$ for each $r$.
@@
 To achieve this, we introduce three connected relations: has relation, has head entity and has tail entity, which establish directly connections between atomic nodes and fact nodes.
-Ultimately, we denote the (main) triple $\left(h,r,t\right)$ in original fact as three connected facts: $\left(f,has\,relation,e_{r}\right),\left(f,has\,head\,entity,h\right),$ $(f,has\,tail\,entity,t)$, and a atomic fact $(h,r,t)$, where $f$ is fact node.
+Ultimately, we denote the (main) triple $\left(h,r,t\right)$ in original fact as three connected facts: $\left(f,has\,relation,e_{r}\right),\left(f,has\,head\,entity,h\right),$ $(f,has\,tail\,entity,t)$, and an atomic fact $(h,r,t)$, where $f$ is fact node.
 Formally, the definition of HiDR form is as follows:
-Definition 1.
-Hierarchical Data Representation:
-A KG represented as the HiDR form is denoted as $\mathcal{G}^{HiDR}\,\text{=}\,\{\mathcal{V}^{HiDR},\mathcal{R}^{HiDR},\mathcal{F}^{HiDR}\},$ where $\mathcal{V}^{HiDR}\,\text{=}\,\mathcal{V}_{a}\cup\mathcal{V}_{r}\cup\mathcal{V}_{f}$ is a joint set of atomic node set ($\mathcal{V}_{a}$), relation node set ($\mathcal{V}_{r}$), fact node set ($\mathcal{V}_{f}$).
-$\mathcal{R}^{HiDR}\,\text{=}\,\mathcal{R}_{a}\cup\mathcal{R}_{n}\cup\mathcal{R}_{c}$ is a joint set of atomic relation set ($\mathcal{R}_{a}$), nested relation set ($\mathcal{R}_{n}$), connected relation set $\mathcal{R}_{c}\,\text{=}\{has\,relation,has\,head\,entity,$ $has\,tail\,entity\}$.
+Definition 1
+Hierarchical Data Representation (HiDR): A KG represented as the HiDR form is denoted as $\mathcal{G}^{HiDR}\,\text{=}\,\{\mathcal{V}^{HiDR},\mathcal{R}^{HiDR},\mathcal{F}^{HiDR}\},$ where $\mathcal{V}^{HiDR}\,\text{=}\,$
+$\mathcal{V}_{a}\cup\mathcal{V}_{r}\cup\mathcal{V}_{f}$ is a joint set of atomic node set ($\mathcal{V}_{a}$), relation node set ($\mathcal{V}_{r}$), fact node set ($\mathcal{V}_{f}$).
+$\mathcal{R}^{HiDR}\,\text{=}\,\mathcal{R}_{a}\cup\mathcal{R}_{n}\cup\mathcal{R}_{c}$ is a joint set of atomic relation set ($\mathcal{R}_{a}$), nested relation set ($\mathcal{R}_{n}$), connected relation set $\mathcal{R}_{c}\,\text{=}\{has\,relation,has\,head\,entity,has\,$
+$tail\,entity\}$.
 The fact set $\mathcal{F}^{HiDR}\,\text{=}\,\mathcal{F}_{a}\cup\mathcal{F}_{c}\cup\mathcal{F}_{n}$ is jointly composed of three types of triple-based facts: atomic facts ($\mathcal{F}_{a}$), connected facts ($\mathcal{F}_{c}$) and nested facts ($\mathcal{F}_{n}$), where $\mathcal{F}_{a}\text{=}\,\{(v_{1},r,$ $v_{2})|\,v_{1},v_{2}\in\mathcal{V}_{a},r\in\mathcal{R}_{a}\}$, $\mathcal{F}_{c}\,\text{=}\,\{(v_{1},r,v_{2})|\,v_{1}\in\mathcal{V}_{f},r\in\mathcal{R}_{c},v_{2}\in\mathcal{V}_{a}\}$, $\mathcal{F}_{n}\,\text{=}\,\{(v_{1},r,v_{2})|\,v_{1},v_{2}\in\mathcal{V}_{f},r\in\mathcal{R}_{n}\}$.
 Next, we introduce how to transform different types of KGs into HiDR form.
-For hyper-relational knowledge graphs, we regard key-value pairs as complementary information for facts.
-Thus, we translate H-Facts $\mathcal{F}_{HKG}$ =$\text{=}\,\{((h,r,t),\{(k_{i}\text{:}\,v_{i})\}_{i=1}^{m}\}$ into the HiDR form that $\mathcal{G}^{HiDR}_{HKG}\,\text{=}\,\{\mathcal{V},\mathcal{R},\mathcal{F}^{HiDR}_{HKG}\}$ following the definition, where $\mathcal{F}_{c}\text{=}\{(f,has\,relation,e_{r}),(f,has$ $,\,head\,entity,h),(f,has\,tail\,entity,t),(f,k_{1},v_{1}),$$\ldots,(f,k_{m},v_{m})\}$, $\mathcal{F}_{a}\,\text{=}\,\{(h,r,t)\,|\,((h,r,t),\{(k_{i}\text{:}\,$ $v_{i})\}_{i=1}^{m})\in\mathcal{F}_{HKG})\}$ and $\mathcal{F}_{n}\text{=}\,\varnothing$.
-For nested factual knowledge graphs, HiDR can naturally represent hierarchical facts, so we translate the atomic facts $\mathcal{F}_{NKG}\,\text{=}\,\{(h_{i},r_{i},t_{i})\}$ and the nested facts ^$\hat{\mathcal{F}}_{NKG}\,\text{=}\,\{((h_{1},r_{1},t_{1}),R,(h_{2},$ $r_{2},t_{2}))|(h_{i},r_{i},t_{i})\in\mathcal{F}_{NKG}\}$ into the form of HiDR that $\mathcal{G}^{HiDR}_{NKG}\,\text{=}\,\{\mathcal{V},\mathcal{R},\mathcal{F}^{HiDR}_{NKG}\}$ following the definition, where $\mathcal{F}_{a}\,\text{=}\,\{(h_{i},r_{i},t_{i})\,|\,(h_{i},r_{i},t_{i})\in$ $\mathcal{F}_{NKG}\}$, $\mathcal{F}_{c}\,\text{=}\{(f_{i},has\,head\,entity,h_{i}),(f_{i},has$ $tail\,entity,t_{i}),(f_{i},has\,relation,e_{r_{i}})|f_{i}\,\text{=}\,(h_{i},r_{i},$ $t_{i})\in\mathcal{F}_{NKG}\}$ and $\mathcal{F}_{n}\text{=}\{(f_{1},R,f_{2})|f_{i}\in\mathcal{F}_{NKG}\}$.
+For hyper-relational knowledge graphs, we regard key-value pairs as complementary information for facts. Thus, we translate the H-Facts $\mathcal{F}_{HKG}\,\text{=}$
+$\{((h,r,t),\{(k_{i}\text{:}\,v_{i})\}_{i=1}^{m}\}$ into the HiDR form that $\mathcal{G}^{HiDR}_{HKG}\,\text{=}\,\{\mathcal{V},\mathcal{R},\mathcal{F}^{HiDR}_{HKG}\}$ following the definition, where $\mathcal{F}_{c}\,\text{=}\,\{(f,has\,relation,e_{r}),(f,has$ $\,head\,entity,h),(f,has$
+$tail\,entity,t),(f,k_{1},v_{1}),\ldots,(f,k_{m},v_{m})\}$, $\mathcal{F}_{a}\,\text{=}\,\{(h,r,t)\,|\,((h,r,t),\{(k_{i}\text{:}\,v_{i})\}_{i=1}^{m})\in\mathcal{F}_{HKG})\}$ and $\mathcal{F}_{n}\text{=}\,\varnothing$.
+For nested factual knowledge graphs, HiDR can naturally represent hierarchical facts, so we translate the atomic facts $\mathcal{F}_{NKG}\,\text{=}\,\{(h_{i},r_{i},t_{i})\}$ and the nested facts ^$\hat{\mathcal{F}}_{NKG}\,\text{=}\,\{((h_{1},r_{1},t_{1}),R,(h_{2},$ $r_{2},t_{2}))|(h_{i},r_{i},t_{i})\in\mathcal{F}_{NKG}\}$ into the form of HiDR that $\mathcal{G}^{HiDR}_{NKG}\,\text{=}\,\{\mathcal{V},\mathcal{R},\mathcal{F}^{HiDR}_{NKG}\}$ following the definition, where $\mathcal{F}_{a}\,\text{=}\,\{(h_{i},r_{i},t_{i})\,|\,(h_{i},r_{i},t_{i})\in$ $\mathcal{F}_{NKG}\}$, $\mathcal{F}_{c}\,\text{=}\{(f_{i},has\,head\,entity,h_{i}),(f_{i},has\,tail$
+$entity,t_{i}),(f_{i},has\,relation,e_{r_{i}})|f_{i}\,\text{=}\,(h_{i},r_{i},$ $t_{i})\in\mathcal{F}_{NKG}\}$ and $\mathcal{F}_{n}\text{=}\{(f_{1},R,f_{2})|f_{i}\in\mathcal{F}_{NKG}\}$.
 For temporal knowledge graphs, we regard the TKG as a special HKG, and convert timestamps to auxiliary key-value pairs in HKGs by adding two special atomic relations: begin and end, regarding timestamps as special numerical atomic nodes.
 Thus, we firstly translate the temporal facts in TKGs $\mathcal{F}_{TKG}\text{=}\,\{(h,r,t,[\tau_{b},\tau_{e}])\}$ into H-Facts form $\mathcal{F}_{TKG}^{HKG}\text{=}\,\{(h,r,t,begin\text{:}\tau_{b},end\text{:}\tau_{e})\}$.
-Then according to the previous transformation in HKG, it can be translated into the HiDR form that $\mathcal{G}_{TKG}^{HiDR}\,\text{=}\,\{\mathcal{V},\mathcal{R},\mathcal{F}_{TKG}^{HiDR}\}$ following the definition,where $\mathcal{F}_{a}\,\text{=}\{(h,r,t)\,|\,(h,r,t,begin\text{:}\,\tau_{b},end\text{:}$ $\tau_{e})\in\mathcal{F}_{TKG}^{HKG}\}$, $\mathcal{F}_{c}\,\text{=}\,\{(f,has\,relation,e_{r}),(f,$ $has\,head\,entity,h),\,(f,has\,tail\,entity,t),\,(f,$ $begin,\tau_{b}),(f,end,\tau_{e})\,|\,f\,\text{=}\,(h,r,t,begin\text{:}\,\tau_{b},end\text{:}$ $\,\tau_{e})\in\mathcal{F}_{TKG}^{HKG}\}$ and $\mathcal{F}_{n}\,\text{=}\,\varnothing$.
+Then according to the previous transformation in HKG, it can be translated into the HiDR form that $\mathcal{G}_{TKG}^{HiDR}\,\text{=}\,\{\mathcal{V},\mathcal{R},\mathcal{F}_{TKG}^{HiDR}\}$ following the definition, where $\mathcal{F}_{a}\,\text{=}\{(h,r,t)\,|\,(h,r,t,begin\text{:}\,\tau_{b},end\text{:}$ $\tau_{e})\in\mathcal{F}_{TKG}^{HKG}\}$, $\mathcal{F}_{c}\,\text{=}\,\{(f,has\,relation,e_{r}),(f,$ $has\,head\,entity,h),\,(f,has\,tail\,entity,t),\,(f,$ $begin,\tau_{b}),(f,end,\tau_{e})\,|\,f\,\text{=}\,(h,r,t,$
+$begin\text{:}\,\tau_{b},end\text{:}\,\tau_{e})\in\mathcal{F}_{TKG}^{HKG}\}$ and $\mathcal{F}_{n}\,\text{=}\,\varnothing$.
 In summary, we could convert all above KGs into the HiDR form, and preserve the semantics in the original KGs without loss of information.
+Figure 3: HiSL module for intra-fact and inter-fact message passing.
 4.2 Hierarchical Structure Learning
-In this section, we illustrate how various KGs in the form of HiDR can be encoded by the Hierarchical Structure Learning module, abbreviated as HiSL shown in Fig. 3.
-Figure 3: HiSL module for intra-fact and inter-fact MP.
+It’s evident that HiDR form introduces many additional relation nodes and fact nodes.
+To avoid significantly increasing the model’s training parameters while fully capturing the hierarchy of HiDR form, we design a Hierarchical Structure Learning module, abbreviated as HiSL shown in Fig.
+3.
 Representation Initialization.
 We first initialize the embedding matrices $\mathbf{H}_{a}\in\mathbb{R}^{|\mathcal{V}_{a}|\times d}$ and $\mathbf{E}\in\mathbb{R}^{|\mathcal{R}^{HiDR}|\times d}$ for atomic nodes and all relation edges.
-Then we also initialize the embedding of relation node $\mathbf{H}_{r}\in\mathbb{R}^{|\mathcal{V}_{r}|\times d}$, which can be transformed from the relation edge $r$, define as:
-$\mathbf{H}_{r}=\mathbf{E_{a}\cdot W_{r}},$ (1)
-where $\mathbf{E_{a}}\subseteq\mathbf{E},\mathbf{W_{r}}\in\mathbb{R}^{d\times d}$ denote the atomic relation embeddings and a projection matrix.
+Then we also initialize the embedding of relation node $\mathbf{H}_{r}\in\mathbb{R}^{|\mathcal{V}_{r}|\times d}$, which can be transformed from the relation edge $r$ with a projection matrix $\mathbf{W_{r}}\in\mathbb{R}^{d\times d}$: $\mathbf{H}_{r}=\mathbf{E_{a}\cdot W_{r}}$, where $\mathbf{E_{a}}\subseteq\mathbf{E}$ is the atomic relation embeddings.
 Then we initialize the fact node embeddings $\mathbf{H}_{f}$ to explicitly capture key information within facts by utilizing the embedding of (main) triple:
 $\mathbf{h}_{f}=f_{m}([\mathbf{h}_{h};\mathbf{h}_{r};\mathbf{h}_{t}]),$ (1)
-where $(h,r,t)\in\mathcal{F}_{a}$, the operation [;][$\cdot$;$\cdot$][ ⋅ ; ⋅ ] is the vector concatenation, $\mathbf{h}_{h},\mathbf{h}_{t}\subseteq\mathbf{H}_{a},\mathbf{h}_{r}\subseteq\mathbf{H}_{r}$ denote (main) triple embedding and $f_{m}\text{:}\;\mathbb{R}^{3d}\rightarrow\mathbb{R}^{d}$ is a 1-layer MLP.
-For numerical atomic nodes, namely timestamps in temporal knowledge graphs, we utilize the Time2Vec Kazemi et al. (2019) to encode the timestamp $\tau$ into an embedding:
+where $(h,r,t)\in\mathcal{F}_{a}$, [;][$\cdot$;$\cdot$][ ⋅ ; ⋅ ] is the concatenation operation, $\mathbf{h}_{h},\mathbf{h}_{t}\subseteq\mathbf{H}_{a},\mathbf{h}_{r}\subseteq\mathbf{H}_{r}$ denote (main) triple embedding and $f_{m}\text{:}\;\mathbb{R}^{3d}\rightarrow\mathbb{R}^{d}$ is a 1-layer MLP.
+Therefore, the initialization of relation nodes and fact nodes is sufficiently parameter-efficient.
+For numerical atomic nodes, namely timestamps in temporal knowledge graphs, we encode the timestamp $\tau$ into an embedding with Time2Vec [18]:
 $\mathbf{h}_{\tau}=\omega_{p}\sin\left(f_{p}(\tau)\right)+f_{np}(\tau),$ (2)
@@
 Then we retain the edges directly connected to fact node $f_{k}$, thereby constructing a subgraph $\mathcal{G}_{k}\,\text{=}\,\{\mathcal{V}_{k},\mathcal{R}_{k},\mathcal{F}_{k}\}\subseteq\mathcal{G}^{HiDR}$.
-For this subgraph, we employ the graph attention network (GAT) Brody et al.
-(2021) to aggregate local information, computing the attention score $\alpha_{i,j}$ between node $i\in\mathcal{V}_{k}$ and its neighbor $j$.
+For this subgraph, we employ the graph attention network [5] to aggregate local information, computing the attention score $\alpha_{i,j}$ between node $i\in\mathcal{V}_{k}$ and its neighbor $j$.
 The formula for calculating $\alpha_{i,j}$ in the $l$-th layer is as follows:
@@
 Then, the updated node embeddings are obtained by aggregating the information of neighbors according to the attention scores:
-$\mathbf{h}_{i}^{l}=\mathbf{h}_{i}^{l}+\sum\limits_{j\in\mathcal{N}_{i}}\alpha_{i,j}^{l}\cdot\mathbf{W}_{out}^{l}\mathbf{h}_{j}^{l}.$
+$\mathbf{h}_{i}^{l}=\mathbf{h}_{i}^{l}+\sum\limits_{j\in\mathcal{N}_{i}}\alpha_{i,j}^{l}\cdot\mathbf{W}_{out}^{l}\mathbf{h}_{j}^{l}.$ (4)
 Inter-fact Message Passing.
 At this stage, message passing is conducted on the whole graph $\mathcal{G}^{HiDR}$.
-Similar to previous work Vashishth et al.
-(2019), we use a non-parametric aggregation operator $\phi\left(\cdot\right)\text{:}\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ to obtain messages from neighbouring nodes and edges.
-Specifically, we employ the circular-correlation operator inspired from HolE Nickel et al.
-(2016), defined as:
-$\phi\left(\mathbf{h}_{j},\mathbf{e}_{r}\right)=\mathbf{h}_{j}\star\mathbf{e}_{r}=\mathbf{F}^{-1}\left(\left(\mathbf{F}\mathbf{h}_{j}\right)\odot\overline{\left(\mathbf{F}\mathbf{e}_{r}\right)}\right)$
+Similar to previous work [31], we use a non-parametric aggregation operator $\phi\left(\cdot\right)\text{:}\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ to obtain messages from neighbouring nodes and edges.
+We employ the circular-correlation operator, defined as:
+$\phi\left(\mathbf{h}_{j},\mathbf{e}_{r}\right)=\mathbf{h}_{j}\star\mathbf{e}_{r}=\mathbf{F}^{-1}\left(\left(\mathbf{F}\mathbf{h}_{j}\right)\odot\overline{\left(\mathbf{F}\mathbf{e}_{r}\right)}\right),$ (5)
 where $\mathbf{F}$ and $\mathbf{F}^{-1}$ denote the discrete fourier transform (DFT) matrix and its inverse matrix, the $\odot$ is the element-wise (Hadamard) product.
-Furthermore, in order to fully capture the heterogeneity of the graph, we classify edges along two dimensions: $\lambda(r)\in\left\{forward,reverse\right\}$ and $\tau(r)\in\{connected\,\,relation,atomic\,\,relation,nested\,$ $relation\}$ and adopt two relation-type specific learnable parameters $\mathbf{W}_{\lambda\left(r\right)}\in\mathbb{R}^{d\times d}$ and $\omega_{\tau\left(r\right)}\in$ $\mathbb{R}^{1}$ for more fine-grained aggregation:
-$\mathbf{h}_{i}^{l+1}=\sum\limits_{(r,j)\in\mathcal{N}(i)}{\sigma\left(\omega_{\tau\left(r\right)}^{l}\right)\mathbf{W}_{{\lambda}\left(r\right)}^{{l}}\phi\left(\mathbf{h}_{j}^{l},\mathbf{e}_{r}^{l}\right)}+\mathbf{W}_{self}^{l}\mathbf{h}_{i}^{l}$
-$\mathbf{e}_{r}^{l+1}=\mathbf{W}_{rel}^{l}\mathbf{e}_{r}^{l}$
+Furthermore, in order to fully capture the heterogeneity of the graph, we classify edges along two dimensions: $\lambda(r)\in\left\{forward,reverse\right\}$ and $\tau(r)\in\{connected\,\,relation,atomic\,\,$
+$relation,nested\,relation\}$ and adopt two relation-type specific learnable parameters $\mathbf{W}_{\lambda\left(r\right)}\in\mathbb{R}^{d\times d}$ and $\omega_{\tau\left(r\right)}\in$ $\mathbb{R}^{1}$ for more fine-grained aggregation:
+$\mathbf{h}_{i}^{l+1}=\sum\limits_{(r,j)\in\mathcal{N}(i)}{\sigma\left(\omega_{\tau\left(r\right)}^{l}\right)\mathbf{W}_{{\lambda}\left(r\right)}^{{l}}\phi\left(\mathbf{h}_{j}^{l},\mathbf{e}_{r}^{l}\right)}+\mathbf{W}_{self}^{l}\mathbf{h}_{i}^{l},$ (6)
+$\mathbf{e}_{r}^{l+1}=\mathbf{W}_{rel}^{l}\mathbf{e}_{r}^{l},$ (7)
 where $\mathbf{W}^{l}_{self},\mathbf{W}^{l}_{rel}\in\mathbb{R}^{d\times d}$, $\sigma$ is a sigmoid activation function and $\mathcal{N}\left(i\right)$ is a set of immediate neighbors of $i$ for its outgoing edges $r$.
-We utilize $\phi\left(\cdot\right)$ to combine the information from both edge $r$ and node $j$, and then passes it to node $i$.
-Subsequently, node $i$ aggregates the information according to the types of edge $r$ separately to update its embedding, while edge $r$ is also projected into the same embedding space as the updated nodes.
+We utilize $\phi\left(\cdot\right)$ to combine the information from edge $r$ and node $j$, and then passes it to node $i$ for update.
+Through Intra-fact and Inter-fact two-stage message passing, nodes can fully capture both local semantic and global structural information.
+Moreover, its number of training parameters does not increase with the scale of the graph, thereby effectively adapting to the HiDR form.
 4.3 Link Prediction Decoder
-Since the query varies across different settings, we use the transformer Vaswani et al.
-(2017) as the decoder with the mask pattern.
+Since the query varies across different settings, we use the transformer [32] as the decoder with the mask pattern.
 Specifically, we convert the updated node and edge embeddings into a sequence of fact embeddings, mask the elements to be predicted in facts with the $\bm{\left[M\right]}$ token as the input to the transformer.
-Finally, we obtain the embedding of output $\bm{\left[M\right]}$ in the last layer to measure the plausibility of the fact, denoted as $\mathbf{h}_{pre}$, and calculate the probability distribution of candidates, followed by training it using the cross-entropy loss function:
-$P=\,\,{\rm Softmax}\left(f\left(\mathbf{h}_{pre}\right)[\mathbf{E};\mathbf{H}]^{\top}\right),$
-$\mathcal{L}=\sum_{t=0}^{\left|\mathcal{R}\right|+\left|\mathcal{V}\right|}{y_{t}\log P_{t}}$
+Finally, we obtain the embedding of output $\bm{\left[M\right]}$ in the last layer to measure the plausibility of the fact, denoted as $\mathbf{h}_{pre}$, and calculate the probability distribution of candidates, followed by training it using the cross-entropy loss:
+$P=\,{\rm Softmax}\left(f\left(\mathbf{h}_{pre}\right)[\mathbf{E};\mathbf{H}]^{\top}\right),$ (8)
+$\mathcal{L}=\sum_{t=0}^{\left|\mathcal{R}\right|+\left|\mathcal{V}\right|}{y_{t}\log P_{t}},$ (9)
 where $P\in\mathbb{R}^{\left|\mathcal{R}\right|+\left|\mathcal{V}\right|}$ represents the confidence scores of all candidates, $f\text{:}\;\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ is a 1-layer MLP, and $[\mathbf{E};\mathbf{H}]\in\mathbb{R}^{(\left|\mathcal{R}\right|+\left|\mathcal{V}\right|)\times d}$ is the embedding matrix of all candidate edges or nodes.
@@
 The final loss function $\mathcal{L}$ includes both node loss and edge loss during the predictions.
-Table 1: Results of link prediction on HKG datasets.
-All baselines’ results are taken from Chung et al.
-(2023).
-The best results are written bold, while the second are underlined.
-# Tr.
-Params denotes the number of learnable parameters during training.
-WikiPeople WD50K
-Model # Tr. Params subject/object all entities # Tr. Params subject/object all entities
-MRR Hits@1 Hits@10 MRR Hits@1 Hits@10 MRR Hits@1 Hits@10 MRR Hits@1 Hits@10
-NaLP - 0.356 0.271 0.499 0.360 0.275 0.503 - 0.230 0.170 0.347 0.251 0.187 0.375
-tNaLP - 0.358 0.288 0.486 0.361 0.290 0.490 - 0.221 0.163 0.331 0.243 0.182 0.360
-RAM 27.34M 0.459 0.384 0.584 0.461 0.386 0.585 - 0.276 0.210 0.399 0.296 0.232 0.416
-HINGE - 0.393 0.309 0.547 0.395 0.311 0.549 - 0.264 0.187 0.410 0.277 0.200 0.424
-NeuInfer - 0.357 0.247 0.533 0.357 0.248 0.532 - 0.220 0.154 0.347 0.225 0.158 0.355
-StarE 7.84M 0.458 0.364 0.611 - - - 10.39M 0.309 0.234 0.452 - - -
-HyTransformer - 0.460 0.382 0.594 - - - - 0.304 0.231 0.443 - - -
-GRAN 15.26M 0.462 0.366 0.610 0.465 0.371 0.613 18.51M 0.330 0.255 0.472 0.361 0.286 0.501
-HyNT 23.02M 0.482 0.415 0.602 0.481 0.414 0.603 29.61M 0.333 0.259 0.474 0.360 0.287 0.500
-UniHR 8.02M 0.491 0.417 0.618 0.493 0.420 0.621 10.55M 0.348 0.278 0.482 0.382 0.313 0.513
-Table 2: Results of base link prediction (left) and triple prediction (right).
-All baselines’ results are taken from Xiong et al.
-(2024).
-For BiVE Chung and Whang (2023) and NestE Xiong et al.
-(2024), we pick their best variants.
-Model FBHE/FBH DBHE FBH FBHE DBHE
-MRR Hits@10 MRR Hits@10 #Tr. Params MR MRR Hits@10 #Tr. Params MR MRR Hits@10 #Tr. Params MR MRR Hits@10
-Base link prediction Triple prediction
-QuatE 0.354 0.581 0.264 0.440 - 145603.8 0.103 0.114 - 94684.4 0.101 0.209 - 26485.0 0.157 0.179
-BiQUE 0.356 0.583 0.274 0.446 - 81687.5 0.104 0.115 - 61015.2 0.135 0.205 - 19079.4 0.163 0.185
-Neural-LP 0.315 0.486 0.233 0.357 - 115016.6 0.070 0.073 - 90000.4 0.238 0.274 - 21130.5 0.170 0.209
-DRUM 0.317 0.490 0.237 0.359 - 115016.6 0.069 0.073 - 90000.3 0.261 0.274 - 21130.5 0.166 0.209
-AnyBURL 0.310 0.526 0.220 0.364 - 108079.6 0.096 0.108 - 83136.8 0.191 0.252 - 20530.8 0.177 0.214
-BiVE 0.370 0.607 0.274 0.422 12.66M 6.20 0.855 0.941 12.67M 8.35 0.711 0.866 10.74M 3.63 0.687 0.958
-NestE 0.371 0.608 0.289 0.443 12.11M 3.34 0.922 0.982 12.46M 3.05 0.851 0.962 10.22M 2.07 0.862 0.984
-UniHR 0.401 0.619 0.300 0.455 4.12M 2.46 0.946 0.993 4.12M 5.20 0.793 0.890 3.67M 1.90 0.862 0.987
 5 Experiment
@@
 Datasets.
-For link prediction on HKGs, we select three benchmark datasets: WikiPeople Guan et al.
-(2021) , WD50K Galkin et al.
-(2020) and JF17K Wen et al.
-(2016).
-As for the NKGs, we choose FBH, FBHE and DBHE constructed by Chung and Whang (2023).
-Lastly, we use wikidata12k Dasgupta et al.
-(2018), a subset of wikidata Vrandečić and Krötzsch (2014) for link prediction on TKGs.
-The statics of datasets are given in Appendix D.
+For link prediction on HKGs, we select three benchmark datasets: WikiPeople [15] and WD50K [12].
+For link prediction on NKGs, we choose FBH, FBHE and DBHE constructed by [7].
+For link prediction on TKGs, we use wikidata12k [8], a subset of wikidata [33].
+To further evaluate the potential of the unified representation, we further introduce the hyper-relational TKG datasets WIKI-hy and YAGO-hy [10].
+The detailed statics are given in Table 1.
+Table 1: The statistics of diverse knowledge graphs dataset, where “with Q(%)" and “Arity" column respectively denote the number of facts with auxiliary key-value pairs and the range of arity of hyper-relational facts, “N-Fact” is the number of nested fact, “N-Rela” is the number of nested relation, the “AF(%)" column denotes the number of atomic facts in nested facts.
+Dataset Fact Entities Rela Train Valid Test with Q(%) Arity N-Fact N-Rela AF(%) Period
+Hyper-relational Knowledge Graph
+WikiPeople 369866 34825 178 294439 37715 37712 9482(2.6%) 2-7 - - - -
+WD50K 236507 47155 531 166435 23913 46159 32167(13.6%) 2-67 - - - -
+Nested Factual Knowledge Graph
+FBH 310116 14541 237 248094 31011 31011 - - 27062 6 33157 -
+FBHE 310116 14541 237 248094 31011 31011 - - 34941 10 33719 -
+DBHE 68296 12440 87 54636 6830 6830 - - 6717 8 8206 -
+Temporal Knowledge Graph
+wikidata12k 40621 12554 24 32497 4062 4062 - - - - - 19-2020
+Hyper-relational Temporal Knowledge Graph
+WiKi-hy 139078 16634 147 111252 13900 13926 13335(9.59%) 2-8 - - - 1513-2020
+YAGO-hy 73143 16167 54 51193 10973 10977 5107(6.98%) 2-5 - - - 0-187
+Multiple types of Knowledge Graph
+wikimix 409566 43832 184 326936 41777 3525/37328 9098(2.2%) 2-7 - - - 19-2020
 Evaluation Metric.
@@
 We use the MR (Mean Rank), MRR (Mean Reciprocal Rank) and Hits@K (K=1,3,10) as our evaluation metrics.
-And we choose to adopt the metrics used in prior works.
-Additionally, we employ filtering settings Bordes et al.
-(2013) during the evaluation process to eliminate existing facts in the dataset.
+We abbreviate ‘Hits@K’ as ‘H@K’ and employ filtering settings [4] during the evaluation to eliminate existing facts in the dataset.
+It is worth noting that for the query $((h,r,?),{(k_{i}:v_{i})}_{i=1}^{m})$ in HKGs, there are two evaluation filtering settings in existing models: one that filters out facts satisfying $((h,r,?),{(k_{i}:v_{i})}_{i=1}^{m})$ and another that filters out facts satisfying only $(h,r,?)$ in the training set.
+Similarly, the difference in filtering settings of TKG occurs in timestamp.
+In this paper, we adopt the strict filtering setting of the former.
+To ensure fair comparison, for HKG we utilize the results of HyNT [6] with the same settings as ours.
+For TKG, we thoroughly review the original code of our baselines and reproduce the results of some methods.
 Baselines.
-For link prediction on HKG, we compare our UniHR against NaLP Guan et al.
-(2021), tNaLP Guan et al.
-(2021), RAM Liu et al.
-(2021), HINGE Rosso et al.
-(2020), NeuInfer Guan et al.
-(2020), StarE Galkin et al.
-(2020), HyTransformer Yu and Yang (2021), GRAN Wang et al.
-(2021) and HyNT Chung et al.
-(2023).
-For link prediction on NKG, QuatE Zhang et al.
-(2019), BiQUE Guo and Kok (2021), Neural-LP Yang et al.
-(2017), DRUM Sadeghian et al.
-(2019), AnyBURL Meilicke et al.
-(2019), BiVE Chung and Whang (2023) and NestE Xiong et al.
-(2024) are chosen as baselines.
-BiVE and NestE are especially designed for NKG.
-We compare against following TKG link prediction methods: ComplEx-N3 Lacroix et al.
-(2018), TTransE Leblay and Chekol (2018), HyTE Dasgupta et al.
-(2018), TA-DistMult Garcia-Duran et al.
-(2018), ATiSE Xu et al.
-(2019), TeRo Xu et al.
-(2020), TASTER Wang et al.
-(2023), TGeomE+ Xu et al.
-(2023a).
+For link prediction on HKG, we compare our UniHR with NaLP [15], tNaLP [15], RAM [22], HINGE [27], NeuInfer [14], StarE [12], HyTransformer [43], GRAN [34], HyNT [6] and HAHE [23].
+For link prediction on NKG, QuatE [45], BiQUE [16], Neural-LP [42], DRUM [28], AnyBURL [24], BiVE [7] and NestE [37] are chosen as baselines.
+Among them, BiVE and NestE are especially designed for NKG.
+We compare against following TKG link prediction methods: ComplEx-N3 [20], HyTE [8], TA-DistMult [13], ATiSE [38], TeRo [39], TGeomE+ [40], HGE [26] and ECEformer [11].
 Implementation details.
 All experiments are conducted on a single Nvidia 80G A800 GPU and implemented with PyTorch.
-For base link prediction on NKGs, we also use augmented triples from Chung and Whang (2023) for training to ensure fairness.
+For base link prediction on NKGs, we also use augmented triples from [7] for training to ensure fairness.
 For triple prediction, due to the small size of training set, we conduct training based on fixed embeddings of entities obtained from the base link prediction and set $\omega_{nested\;relation}\text{=}\ 0$ to prevent overfitting.
-We employ AdamW Kingma and Ba (2015) optimizer.
-Hyperparameters are chosen by using a grid search based on the MRR performance and details can be found in Appendix F.
+We employ AdamW [19] optimizer.
+Hyperparameter details can be found in Appendix C.
 5.2 Main Results
-Table 3: Results of link prediction on wikidata12k.
+Table 2: Results of link prediction on HKG datasets. Results are reported by us, and the others are taken from [6]. Best results are bold, while the second are underlined.
+WikiPeople WD50K
+Model subject/object all entities subject/object all entities
+MRR H@1 H@10 MRR H@1 H@10 MRR H@1 H@10 MRR H@1 H@10
+NaLP [15] 0.356 0.271 0.499 0.360 0.275 0.503 0.230 0.170 0.347 0.251 0.187 0.375
+tNaLP [15] 0.358 0.288 0.486 0.361 0.290 0.490 0.221 0.163 0.331 0.243 0.182 0.360
+RAM [22] 0.459 0.384 0.584 0.461 0.386 0.585 0.276 0.210 0.399 0.296 0.232 0.416
+HINGE [27] 0.393 0.309 0.547 0.395 0.311 0.549 0.264 0.187 0.410 0.277 0.200 0.424
+NeuInfer [14] 0.357 0.247 0.533 0.357 0.248 0.532 0.220 0.154 0.347 0.225 0.158 0.355
+StarE [12] 0.458 0.364 0.611 - - - 0.309 0.234 0.452 - - -
+HyTransformer [43] 0.465 0.395 0.598 - - - 0.308 0.239 0.452 - - -
+GRAN [34] 0.477 0.408 0.596 0.480 0.411 0.599 0.329 0.259 0.465 0.363 0.294 0.493
+HyNT [6] 0.482 0.415 0.602 0.481 0.414 0.603 0.333 0.259 0.474 0.360 0.287 0.500
+HAHE [23] 0.498 0.426 0.610 0.497 0.429 0.614 0.343 0.270 0.484 0.378 0.306 0.515
+UniHR 0.491 0.417 0.618 0.493 0.420 0.621 0.348 0.278 0.482 0.382 0.313 0.513
+Link Prediction on HKG.
+We compare our method with previous methods on the WD50K and WikiPeople datasets shown in Table 2.
+We can observe that the models based on transformers (i.e., StarE, GRAN, HyNT, HAHE) demonstrate significantly better performance compared to other models.
+We attribute this to the transformer’s superior ability to capture long distance dependencies within H-Facts.
+Among these methods, it can be seen that our proposed UniHR achieves competitive results with the state-of-the-art method HAHE, which means our method effectively captures hierarchical fact information.
+Compared to the GNN-based method StarE, we achieve improvements of 3.9 points (12.6%) in MRR, 4.4 points (18.8%) in Hits@1 and 3.0 points (6.6%) in Hits@10 on WD50K.
+This indicates that the performance of StarE’s customized graph neural network is limited by its inability to flexibly capture key-value pair information.
+Moreover, since the embeddings for newly added fact nodes and relation nodes are computed from atomic facts through simple MLPs, so our training parameters do not significantly increase.
+Table 3: Results of base link prediction (left) and triple prediction (right). All baselines’ results are taken from [37]. We pick the best variants of BiVE [7] and NestE [37].
+FBH DBHE FBH FBHE DBHE
+Model MRR H@10 MRR H@10 MR MRR H@10 MR MRR H@10 MR MRR H@10
+Base link prediction Triple prediction
+QuatE [45] 0.354 0.581 0.264 0.440 145603.8 0.103 0.114 94684.4 0.101 0.209 26485.0 0.157 0.179
+BiQUE [16] 0.356 0.583 0.274 0.446 81687.5 0.104 0.115 61015.2 0.135 0.205 19079.4 0.163 0.185
+Neural-LP [42] 0.315 0.486 0.233 0.357 115016.6 0.070 0.073 90000.4 0.238 0.274 21130.5 0.170 0.209
+DRUM [12] 0.317 0.490 0.237 0.359 115016.6 0.069 0.073 90000.3 0.261 0.274 21130.5 0.166 0.209
+AnyBURL [24] 0.310 0.526 0.220 0.364 108079.6 0.096 0.108 83136.8 0.191 0.252 20530.8 0.177 0.214
+BiVE [7] 0.370 0.607 0.274 0.422 6.20 0.855 0.941 8.35 0.711 0.866 3.63 0.687 0.958
+NestE [37] 0.371 0.608 0.289 0.443 3.34 0.922 0.982 3.05 0.851 0.962 2.07 0.862 0.984
+UniHR 0.401 0.619 0.296 0.448 2.46 0.946 0.993 5.20 0.793 0.890 1.90 0.862 0.987
+Link Prediction on NKG.
+Our experiments on the NKGs consist of two tasks: base link prediction and triple prediction.
+From the results in Table 3, we can see that our proposed UniHR obtains competitive results as the first method to capture global structural information of NKGs.
+For base link prediction task on triple-based KGs, UniHR achieves considerable improvements.
+Of particular note, the MRR of FBHE increases by 8.1%.
+For triple prediction, we perform best on FBH and DBHE datasets, especially obtaining an improvement of 2.4 points in MRR on FBH, and achieve the second-best performance on FBHE, which suggests that structural information is also valuable for NKG and UniHR can effectively capture the heterogeneity of NKG to enhance node embeddings.
+In particular, as a unified method, we do not use the customized decoder for triples, while previous state-of-the-art methods do.
+We will further illustrate the effectiveness of UniHR equipped with other decoders in Appendix B.
+Table 4: Results of link prediction on wikidata12k. Results are reported by us, and others are taken from original papers.
 Model wikidata12k
-# Tr. Params MRR Hits@1 Hits@3 Hits@10
-ComplEx-N3 - 0.248 0.143 - 0.489
-TTransE - 0.172 0.096 0.184 0.329
-HyTE 1.69M 0.253 0.147 - 0.483
-TA-DistMult - 0.230 0.130 - 0.461
-TeRo 12.91M 0.299 0.198 0.329 0.507
-ATiSE 31.46M 0.252 0.148 0.288 0.462
-TASTER 5.04M 0.327 0.239 - 0.514
-TGeomE+ - 0.333 0.232 0.361 0.546
-UniHR 3.68M 0.333 0.240 0.367 0.527
-Table 4: Ablation studies on the HiSL module for triple prediction. Best results are boldfaced.
+MRR H@1 H@3 H@10
+ComplEx-N3 [20] 0.248 0.143 - 0.489
+HyTE [8] 0.253 0.147 - 0.483
+TA-DistMult [13] 0.230 0.130 - 0.461
+TeRo [39] 0.299 0.198 0.329 0.507
+ATiSE [38] 0.252 0.148 0.288 0.462
+TGeomE+ [40] 0.333 0.232 0.361 0.546
+HGE [26] 0.290 0.176 0.323 0.514
+ECEformer [11] 0.262 0.159 0.255 0.462
+UniHR 0.333 0.240 0.367 0.527
+Link Prediction on TKG.
+As shown in Table 4, we achieve competitive results on the wikidata12k, even surpassing TGeomE+ by 3.4% on Hits@1 and 1.7% on Hits@3.
+However, existing temporal knowledge graph embedding (KGE) methods, such as TGeomE+, often employ time-aware decoders, which are challenging to generalize to other types of KGs.
+In contrast, our approach efficiently encodes timestamps as atomic nodes only during initialization and learns temporal information through message passing on graph structure, demonstrating that graph structure information is also beneficial for temporal knowledge graphs, highlighting the effectiveness of our UniHR.
+Table 5: Ablation studies on triple prediction task.
 Variant FBH FBHE DBHE
-MR MRR Hits@10 MR MRR Hits@10 MR MRR Hits@10
+MR MRR H@10 MR MRR H@10 MR MRR H@10
 w/o initial $\textbf{h}_{f}$ 5.22 0.909 0.980 5.86 0.767 0.885 2.56 0.794 0.978
@@
 UniHR 2.46 0.946 0.993 5.20 0.793 0.890 1.90 0.862 0.987
+5.3 Ablation Study on HiSL
+We conduct ablation experiments on triple prediction, the most relevant task to fact nodes.
+As shown in Table 5, all variants with certain modules or parameters removed exhibit a decrease in performance.
+We can conclude that intra-fact and inter-fact message passing modules both play crucial roles, allowing UniHR to more fully represent the current fact node with enhanced structural information.
+We also change the initialization of fact node embeddings to learnable embeddings (w/o initial $\textbf{h}_{f}$), and results indicate that initializing fact node representations is essential.
+It highlights key information in the facts, mitigating the noise introduced by excessive neighbors.
+5.4 Potential of Unified Representation
+Table 6: Results on hyper-relational TKG datasets.
+Model WiKi-hy YAGO-hy
+MRR H@1 H@3 H@10 MRR H@1 H@3 H@10
+TeRo [39] 0.572 0.473 0.640 0.727 0.760 0.720 0.782 0.822
+HGE [26] 0.602 0.507 0.666 0.765 0.790 0.760 0.814 0.837
+StarE [12] 0.565 0.491 0.599 0.703 0.765 0.737 0.776 0.820
+GRAN [34] 0.661 0.610 0.679 0.750 0.808 0.789 0.817 0.842
+HyNT [6] 0.537 0.444 0.587 0.723 0.763 0.724 0.787 0.836
+HypeTKG [10] 0.687 0.633 0.710 0.789 0.832 0.817 0.838 0.857
+UniHR 0.681 0.618 0.717 0.774 0.823 0.801 0.839 0.859
+Generalize to Hyper-relational TKGs.
+Due to the unity of representation, UniHR can easily generalize to compositional types of KGs, like hyper-relational TKGs (HTKGs) [10] that integrate HKGs and TKGs.
+Every H-fact in the HTKG is coupled with a timestamp explicitly indicating its time validity.
+From results in Table 6, UniHR offers a significant performance improvement in link prediction tasks on hyper-relational TKGs compared to TKG-specific and HKG-specific models, demonstrating a good collaborative modeling ability for auxiliary key-value pairs and timestamps.
+Our UniHR even achieves competitive performance with the specialized hyper-relational TKG model HypeTKG without complex module stacking.
+Table 7: Results of separate and joint training on NKG.
+FBH DBHE FBH DBHE
+Model MRR H@10 MRR H@10 MRR H@10 MRR H@10
+Base link prediction Triple prediction
+UniHR 0.401 0.619 0.296 0.448 0.946 0.993 0.862 0.987
+UniHRJoint 0.402 0.632 0.298 0.452 0.948 0.993 0.857 0.989
+Joint Learning on Different Tasks of KGs.
+For link prediction on NKGs, the two subtasks, namely base link prediction and triple prediction, share the same graph during the message-passing phase under our unified representation form.
+Therefore, we attempt joint training on the two tasks using the NKG dataset, as shown in Table 7.
+We observe that the results of joint training are generally superior to those of separate training, indicating that nested facts and atomic facts can mutually enhance and supplement the semantics.
 Table 8: Results of separate training and joint training on the HKG and TKG dataset, where identical entities and relations share the same embeddings.
 WikiPeople- and wikidata12k- represent the filtered test sets.
-Model WikiPeople- wikidata12k-
-subject/object all entities subject/object
-MR MRR Hits@1 Hits@3 Hits@10 MR MRR Hits@1 Hits@3 Hits@10 MR MRR Hits@1 Hits@3 Hits@10
+WikiPeople- wikidata12k-
+Model subject/object all entities subject/object
+MR MRR H@1 H@3 H@10 MR MRR H@1 H@3 H@10 MR MRR H@1 H@3 H@10
 UniHR 835.8 0.486 0.412 0.528 0.617 829.0 0.488 0.414 0.531 0.620 818.7 0.314 0.220 0.345 0.509
 UniHRJoint 692.7 0.488 0.409 0.533 0.629 686.5 0.490 0.414 0.536 0.632 489.5 0.315 0.222 0.346 0.498
-Link prediction on HKG.
-We compare our method with previous methods on the WD50K and WikiPeople datasets shown in Table 1.
-We can observe that the models based on transformers (i.e., StarE, GRAN, HyNT) demonstrate significantly better performance compared to other models.
-We attribute this to the transformer’s superior ability to capture long distance dependencies within H-Facts.
-Among these methods, it can be seen that our proposed UniHR achieves state-of-the-art results, which means our method effectively captures global structural information.
-Compared to the GNN-based method StarE, we achieve improvements of 3.9 points (12.6%) in MRR, 4.4 points (18.8%) in Hits@1 and 3.0 points (6.6%) in Hits@10 on WD50K.
-This indicates that the performance of StarE’s customized graph neural network is limited by its inability to flexibly capture key-value pair information.
-Moreover, since the embeddings for newly added fact nodes and relation nodes are computed from atomic facts, so our training parameters do not significantly increase.
-Link prediction on NKG.
-Our experiments on the NKGs consist of two tasks: base link prediction and triple prediction.
-Unlike previous quaternion-based methods Chung and Whang (2023); Xiong et al.
-(2024), UniHR significantly reduces the number of required training parameters.
-From the results in Table 2, we can see that our proposed UniHR obtains competitive results as the first method to capture global structural information of NKGs.
-For base link prediction task, UniHR achieves considerable improvements on all datasets.
-Of particular note, the MRR of FBHE/FBH and DBHE increases by 8.1% and 3.8%, respectively.
-For triple prediction, we perform best on FBH and DBHE datasets, especially obtaining an improvement of 2.4 points in MRR on FBH, and achieve the second-best performance on FBHE, which suggests that structural information is also valuable for NKG and UniHR can effectively capture the heterogeneity of NKG to enhance node embeddings.
-In particular, as a unified method, we do not use the customized decoder for triples, while previous state-of-the-art methods do.
-We will further illustrate the effectiveness of UniHR with other decoders in Appendix C.
-Link prediction on TKG.
-As shown in Table 3, we achieve competitive results on the wikidata12k, even surpassing TGeomE+ by 3.4% on Hits@1 and 1.7% on Hits@3.
-However, existing temporal knowledge graph embedding methods, such as TGeomE+, often employ time-aware decoders, which are challenging to generalize to other types of KGs.
-In contrast, our approach efficiently encodes timestamps as atomic nodes only during initialization and learns temporal information through message passing on graph structure, demonstrating that graph structure information is also beneficial for temporal knowledge graphs, highlighting the effectiveness of our UniHR.
-5.3 Further Analysis
-Ablation study on HiSL.
-We conduct ablation experiments on triple prediction, the most relevant task to fact nodes, using three NKG datasets.
-As shown in Table 4, all variant models with certain modules or parameters removed exhibit a decrease in performance.
-We can conclude that intra-fact and inter-fact message passing modules both play crucial roles, allowing UniHR to more fully represent the current fact node with enhanced structural information.
-We also change the initialization of fact node embeddings to learnable embeddings (w/o initial $\textbf{h}_{f}$) in HiSL.
-There is a performance decrease, indicating that initializing fact node representations is essential.
-It highlights key information in the facts, mitigating the noise introduced by excessive neighbors.
-Potential of Joint Learning on Different Types of KGs.
+Joint Learning on Different Types of KGs.
 We suppose that unified representation makes it possible to develop pre-trained models that integrate multiple types of KGs.
 To explore its potential, we conduct joint learning on different types of KGs.
-Therefore, we construct a hybrid dataset called wikimix which includes two subsets of Wikidata Vrandečić and Krötzsch (2014), namely HKG dataset WikiPeople and TKG dataset wikidata12k, which encompass 3547 identical entities and 18 identical relations.
+Therefore, we construct a hybrid dataset called wikimix which includes two subsets of Wikidata [33], namely HKG dataset WikiPeople and TKG dataset wikidata12k, which encompass 3547 identical entities and 18 identical relations.
 Due to the different types of facts, there are no identical facts in these two subsets.
 To further prevent data leakage, we filter out 537 entries from the wikidata12k test set whose main triples appear in the H-Facts of WikiPeople train set, and 384 entries from the WikiPeople test set whose main triples appear in wikidata12k train set.
-Statics of wikimix are given in Appendix D.
+Statics of wikimix are given in Table 1.
 From the results in Table 8, it is evident that joint learning outperforms separate learning across most metrics.
@@
 Moreover, our UniHR demonstrates good scalability and effectiveness in integrating multiple types of KGs.
-6 Conclusion
+Table 9: Number of training parameters and times.
+Dataset Model Training Params Training Times
+WD50K StarE [12] 10.81M 48h+
+HyNT [6] 29.61M 22.3h
+HAHE [23] 42.35M 29.2h
+UniHR 10.55M 23.4h
+wikidata12k ATiSE [38] 31.46M 5.8h
+HGE [26] 3.44M -
+ECEformer [11] 11.99M 5.6h
+UniHR 3.68M 3.6h
+YAGO-hy HypeTKG [10] 10.83M 37.5h
+UniHR 4.37M 17.6h
+5.5 Efficiency Analysis
+Compared to existing state-of-the-art methods, UniHR does not significantly increase the overhead in terms of model training parameters and runtime as shown in Figure 9.
+The embeddings of relation nodes and fact nodes are computed using a simple MLP based on atomic nodes and relations, which does not significantly increase the number of training parameters.
+For runtime, we do not use the whole graph during all message passing phases and adopt a dropout strategy to prevent overfitting, which can effectively enhance training efficiency.
+Therefore, our UniHR achieves a superior trade-off between model effectiveness and efficiency.
+6 Conclusion and Future Work
 In this paper, we propose UniHR, a unified hierarchical knowledge graph representation learning framework consisting of a Hierarchical Data Representation (HiDR) module and a Hierarchical Structure Learning (HiSL) module.
@@
 Moreover, HiSL captures local semantic information within facts and global structural information between facts.
-Our approach achieves the best or competitive performance on link prediction tasks across three types of KGs.
-Limitations
-The limitations of our paper are summarized as follows:
-Our UniHR
-In this paper, our UniHR framework focuses on link prediction tasks under transductive settings with a single modality.
-In the future, we will investigate how to generalize our HiDR form to more complex tasks such as inductive reasoning Teru et al.
-(2020) and multi-modal scenarios Zhang et al.
-(2024), etc.
-Joint Learning on Different Types of KGs
-Constrained by computational resources, our analysis of the potential for joint training across multiple types of knowledge graphs focuse only on HKG and TKG.
-We believe the unification of knowledge graph representation learning methods is a developing trend that makes it possible to develop unified pre-trained models based on multiple types of KGs.
-In the future, we aim to explore joint training across more types of KG to demonstrate the advantages of integrating multi-type KG data.
-Ethics Statement
-In this paper, we explore the unified knowledge graph link prediction problem, aiming to complete various types of knowledge graphs using a unified model with deep learning techniques.
-Our training and evaluation are based on publicly available and widely used datasets of different types of knowledge graphs.
-Therefore, we believe this does not violate any ethics.
-Appendix A Related Works
-A.1 Link Prediction Methods for Triple-based KGs
+Our UniHR achieves the best or competitive performance across five types of KGs, and highlight the strong potential of unified representations.
+In the future, we believe that the unification of knowledge graph representation learning methods is a developing trend that makes it possible to develop unified pre-trained models based on multiple types of KGs.
+Moreover, we will investigate how to generalize our HiDR form to more complex tasks such as inductive reasoning [30] and multi-modal scenarios, etc.
+Supplemental Material Statement:
+The dataset, code, and hyperparameter details are available at https://github.com/Lza12a/UniHR.
+Appendix
+Appendix 0.A Related Works
+0.A.1 Link Prediction on Triple-based KGs
 Most existing techniques in KG representation learning are proposed for triple-based KGs.
-Among these techniques, knowledge graph embedding (KGE) models Bordes et al.
-(2013); Sun et al.
-(2018) have received extensive attention due to their effectiveness and simplicity.
+Among these techniques, knowledge graph embedding (KGE) models [4, 29] have received extensive attention due to their effectiveness and simplicity.
 The idea is to project entities and relations in the KG to low-dimensional vector spaces, utilizing KGE scoring functions to measure the plausibility of triples in the embedding space.
-Typical methods include TransE Bordes et al.
-(2013), RotatE Sun et al.
-(2018), and ConvE Dettmers et al.
-(2018).
-Depending on the KGE model alone has limitation of capturing complex graph structures, whereas augmenting global structural information with a graph neural network (GNN) Vashishth et al.
-(2019); Nathani et al.
-(2019); Xu et al.
-(2023b) proves to be an effective approach for enhancement.
+Typical methods include TransE [4], RotatE [29], and ConvE [9].
+Depending on the KGE model alone has limitation of capturing complex graph structures, whereas augmenting global structural information with a graph neural network (GNN) [31, 25, 41] proves to be an effective approach for enhancement.
 The paradigm of combining GNN as encoder with KGE scoring function as decoder helps to enhance the performance of KGE scoring function.
 These GNN methods design elaborate message passing mechanisms to capture the global structural features.
-Typically, CompGCN Vashishth et al.
-(2019) aggregates the joint embedding of entities and relations in the neighborhood via a parameter-efficient way and MA-GNN Xu et al.
-(2023b) learns global-local structural information based on multi-attention.
+Typically, CompGCN [31] aggregates the joint embedding of entities and relations in the neighborhood via a parameter-efficient way and MA-GNN [41] learns global-local structural information based on multi-attention.
 These methods achieve impressive results on triple-based KGs but are hard to generalize to beyond-triple KGs.
-Appendix B Results on JF17K
-Table 6 shows the experimental results on JF17K.
-Due to the absence of a validation set in the JF17K dataset and the different ways of dividing the dataset across various baselines, we adopt the results reported in the original paper.
-Consistent with previous experiments on hyper-relational knowledge graphs, we also achieve state-of-the-art performance on JF17K among all baselines.
-In particular, we achieved 1.7 (2.9%) points in MRR and 1.5 (2.9%) points in Hits@1 compared to the method StarE which also utilises a graph neural network encoding and a simple transformer decoding, indicating that our hierarchical GNN HiSL could better capture the structure of hyper-relational facts.
-Table 6: Link prediction on JF17K. All results of baselines are taken from the original paper. Best results are in bold.
-Model JF17K
-subject/object all entities
-MRR Hits@1 Hits@10 MRR Hits@1 Hits@10
-m-TransH 0.206 0.206 0.462 - - -
-NaLP 0.221 0.165 0.331 0.366 0.290 0.516
-HINGE 0.431 0.342 0.611 0.517 0.436 0.675
-NeuInfer 0.449 0.361 0.624 0.473 0.397 0.618
-StarE 0.574 0.496 0.725 - - -
-HyTransformer 0.582 0.501 0.742 - - -
-UniHR 0.591 0.511 0.745 0.621 0.545 0.768
-Appendix C Decoder Analysis
-To explore the effectiveness of our UniHR encoding further, we pair UniHR with different decoders and evaluated them on triple prediction task.
-In addition to the previously mentioned unified framework UniHR + Transformer, we also experiment on UniHR + ConvE with two scoring strategies.
-The ConvE Dettmers et al.
-(2018) is the decoder customized for triples and its scoring function is $vec\left(\sigma\left(\left[\mathbf{\tilde{h}}_{h};\mathbf{\tilde{e}}_{r}\right]*\psi\right)\right)$, where $\mathbf{\tilde{h}}_{h}$ and $\mathbf{\tilde{e}}_{r}$ represent reshaped 2D embeddings of head entity $h$ and relation $r$, and * is a convolution operator.
-The $vec\left(\cdot\right)$ and $\psi$ are denoted as the vectorization function and a set of convolution kernels.
-Due to our special representation, there exists two scoring methods for atomic triples, thus we present the base link prediction results separately for each scoring method.
-The $s_{f}$ represents scoring triples $\left(f,has\;head\;entity,h\right)$and $\left(f,has\;tail\;entity,t\right)$, and $s_{t}$ represents scoring $\left(h,r,t\right)$.
-The performance of base link prediction is shown in Table 7.
-Notably, FBH and FBHE share identical atomic facts, resulting in the same performance.
-It can be observed that regardless of the scoring method employed, we both achieve competitive performance, especially with scoring $\left(h,r,t\right)$ on FBH and scoring $\left(f,has\;head\;entity,h\right)\,\left(f,has\;tail\;entity,t\right)$ on DBHE.
-We attribute the differences in performance under different scoring methods to dataset characteristics.
-The DBHE dataset is relatively smaller, and scoring method $s_{f}$ effectively alleviates overfitting problem.
-Conversely, for larger datasets FBH, scoring based on $\left(h,r,t\right)$ minimizes information loss.
-Table 8 shows the results of triple prediction on three benchmark datasets.
-Among all baselines, Quate, Bique, Neural-LP, Drum, and AnyBURL struggle to model the mapping relationship between atomic facts and nested facts.
-Furthermore, prior works Chung and Whang (2023) do not guarantee that all atomic facts in the nested fact test set are present in the training set as entities, which shifts the problem from a transductive setting to an inductive setting, leading to significant performance gaps between these baselines.
-On most metrics, our method outperforms BiVE and NestE which are specifically modeled for nested facts.
-Notably, NestE fully preserves the semantics of atomic facts.
-However, on the FBHE dataset, UniHR + ConvE achieves an improvement of 0.58 (6.4%) points in MRR and 0.24 (2.4%) points in Hits@10 compared to the state-of-the-art model NestE and the second-best performance after UniHR + Transformer on the FBH and DBHE datasets, demonstrating UniHR’s powerful graph structure encoding capabilities.
-We also carry out ablation experiments on UniHR + ConvE as shown in Table 8.
-Performance declines are observed after removing any part of the HiSL module, showing the significance of HiSL for hierarchical encoding.
 Table 10: Base link prediction on FBHE, FBH and DBHE.
-All baselines’ results are taken from Xiong et al.
-(2024).
+All baselines’ results are taken from [37].
 The best results among all models are written bold, while the second are underlined.
 The $s_{f}$ and $s_{h}$ denote $\left(f,has\;head\;entity,h\right)\ \left(f,has\;tail\;entity,t\right)$ and $\left(h,r,t\right)$ two types of scoring method respectively.
-For BiVE Chung and Whang (2023) and NestE Xiong et al.
-(2024), we pick their variants with best performance.
+For BiVE [7] and NestE [37], we pick their variants with best performance.
 Model FBHE/FBH DBHE
@@
 UniHR + ConvE $s_{f}$ 0.375 0.596 0.307 0.471
-UniHR + Transformer 0.401 0.619 0.300 0.455
+UniHR + Transformer 0.401 0.619 0.296 0.448
+Appendix 0.B Decoder Analysis
+To explore the effectiveness of our UniHR encoding further, we pair UniHR with different decoders and evaluated them on triple prediction task.
+In addition to the previously mentioned unified framework UniHR + Transformer, we also experiment on UniHR + ConvE with two scoring strategies.
+The ConvE [9] is the decoder customized for triples and its scoring function is $vec\left(\sigma\left(\left[\mathbf{\tilde{h}}_{h};\mathbf{\tilde{e}}_{r}\right]*\psi\right)\right)$, where $\mathbf{\tilde{h}}_{h}$ and $\mathbf{\tilde{e}}_{r}$ represent reshaped 2D embeddings of head entity $h$ and relation $r$, and * is a convolution operator.
+The $vec\left(\cdot\right)$ and $\psi$ are denoted as the vectorization function and a set of convolution kernels.
+Due to our special representation, there exists two scoring methods for atomic triples, thus we present the base link prediction results separately for each scoring method.
+The $s_{f}$ represents scoring triples $\left(f,has\;head\;entity,h\right)$and $\left(f,has\;tail\;entity,t\right)$, and $s_{t}$ represents scoring $\left(h,r,t\right)$.
+The performance of base link prediction is shown in Table 10.
+Notably, FBH and FBHE share identical atomic facts, resulting in the same performance.
+It can be observed that regardless of the scoring method employed, we both achieve competitive performance, especially with scoring $\left(h,r,t\right)$ on FBH and scoring $\left(f,has\;head\;entity,h\right)\,\left(f,has\;tail\;entity,t\right)$ on DBHE.
+We attribute the differences in performance under different scoring methods to dataset characteristics.
+The DBHE dataset is relatively smaller, and scoring method $s_{f}$ effectively alleviates overfitting problem.
+Conversely, for larger datasets FBH, scoring based on $\left(h,r,t\right)$ minimizes information loss.
 Table 11: Triple prediction on FBHE, FBH and DBHE.
-All baselines’ results are taken from Xiong et al.
-(2024).
+All baselines’ results are taken from [37].
 The best results among all models are written bold.
-For BiVE Chung and Whang (2023) and NestE Xiong et al.
-(2024), we pick their variants with best performance.
+For BiVE [7] and NestE [37], we pick their variants with best performance.
 Model FBH FBHE DBHE
 MR MRR Hits@10 MR MRR Hits@10 MR MRR Hits@10
-QuatE Zhang et al. (2019) 145603.8 0.103 0.114 94684.4 0.101 0.209 26485.0 0.157 0.179
-BiQUE Guo and Kok (2021) 81687.5 0.104 0.115 61015.2 0.135 0.205 19079.4 0.163 0.185
-Neural-LP Yang et al. (2017) 115016.6 0.070 0.073 90000.4 0.238 0.274 21130.5 0.170 0.209
-DRUM Sadeghian et al. (2019) 115016.6 0.069 0.073 90000.3 0.261 0.274 21130.5 0.166 0.209
-AnyBURL Meilicke et al. (2019) 108079.6 0.096 0.108 83136.8 0.191 0.252 20530.8 0.177 0.214
-BiVE Chung and Whang (2023) 6.20 0.855 0.941 8.35 0.711 0.866 3.63 0.687 0.958
-NestE Xiong et al. (2024) 3.34 0.922 0.982 3.05 0.851 0.962 2.07 0.862 0.984
+QuatE [45] 145603.8 0.103 0.114 94684.4 0.101 0.209 26485.0 0.157 0.179
+BiQUE [16] 81687.5 0.104 0.115 61015.2 0.135 0.205 19079.4 0.163 0.185
+Neural-LP [42] 115016.6 0.070 0.073 90000.4 0.238 0.274 21130.5 0.170 0.209
+DRUM [28] 115016.6 0.069 0.073 90000.3 0.261 0.274 21130.5 0.166 0.209
+AnyBURL [24] 108079.6 0.096 0.108 83136.8 0.191 0.252 20530.8 0.177 0.214
+BiVE [7] 6.20 0.855 0.941 8.35 0.711 0.866 3.63 0.687 0.958
+NestE [37] 3.34 0.922 0.982 3.05 0.851 0.962 2.07 0.862 0.984
 UniHR + Transformer 2.46 0.946 0.993 5.20 0.793 0.890 1.90 0.862 0.987
@@
 UniHR + ConvE w/o inter-fact 12.56 0.864 0.961 20.56 0.864 0.966 10.75 0.764 0.951
-Table 9: The statistics of diverse knowledge graphs dataset, where “with Q(%)" and “Arity" column respectively denote the number of facts with auxiliary key-value pairs and the range of arity of hyper-relational facts, the “with AF(%)" column denotes the number of atomic facts in nested facts.
-Dataset Atomic Fact Entities Relations Train Valid Test with Q(%) Arity Nested Fact Nested Relation with AF(%) Period
-Hyper-relational Knowledge Graph
-WikiPeople 369866 34825 178 294439 37715 37712 9482(2.6%) 2-7 - - - -
-WD50K 236507 47155 531 166435 23913 46159 32167(13.6%) 2-67 - - - -
-JF17K 100947 28645 501 76379 - 24568 46320(45.9%) 2-6 - - - -
-Nested Factual Knowledge Graph
-FBH 310116 14541 237 248094 31011 31011 - - 27062 6 33157 -
-FBHE 310116 14541 237 248094 31011 31011 - - 34941 10 33719 -
-DBHE 68296 12440 87 54636 6830 6830 - - 6717 8 8206 -
-Temporal Knowledge Graph
-wikidata12k 40621 12554 24 32497 4062 4062 - - - - - 19-2020
-Multiple types of Knowledge Graph
-wikimix 409566 43832 184 326936 41777 3525(TKG)/37328(HKG) 9098(2.2%) 2-7 - - - 19-2020
+Table 11 shows the results of triple prediction on three benchmark datasets.
+Among all baselines, Quate, Bique, Neural-LP, Drum, and AnyBURL struggle to model the mapping relationship between atomic facts and nested facts.
+Furthermore, prior works [7] do not guarantee that all atomic facts in the nested fact test set are present in the training set as entities, which shifts the problem from a transductive setting to an inductive setting, leading to significant performance gaps between these baselines.
+On most metrics, our method outperforms BiVE and NestE which are specifically modeled for nested facts.
+Notably, NestE fully preserves the semantics of atomic facts.
+However, on the FBHE dataset, UniHR + ConvE achieves an improvement of 0.58 (6.4%) points in MRR and 0.24 (2.4%) points in Hits@10 compared to the state-of-the-art model NestE and the second-best performance after UniHR + Transformer on the FBH and DBHE datasets, demonstrating UniHR’s powerful graph structure encoding capabilities.
+We also carry out ablation experiments on UniHR + ConvE as shown in Table 11.
+Performance declines are observed after removing any part of the HiSL module, showing the significance of HiSL for hierarchical encoding.
 Table 12: The major hyperparameters of our approach for all link prediction tasks.
@@
 learning rate 5e-4 5e-4 5e-4 5e-4 5e-4 5e-4 5e-4 5e-4 5e-4
-Algorithm 1 Message passing process of HiSL
-Input: $\mathcal{G}^{HiDR}=\{\mathcal{V}_{a}\cup\mathcal{V}_{r}\cup\mathcal{V}_{f},\mathcal{R}_{a}\cup\mathcal{R}_{n}\cup\mathcal{R}_{c},\mathcal{F}_{a}\cup\mathcal{F}_{c}\cup\mathcal{F}_{n}\}$; the number of encoder layers $L$.
-Output: The node embedding matrix $\mathbf{H}^{L}$; the edge embedding matrix $\mathbf{E}^{L}$.
-1:  Initialize the embedding matrix of atomic nodes $\mathbf{H}_{a}^{0}\in\mathbb{R}^{\left|\mathcal{V}_{a}\right|\times d}$.
-2:  Initialize the embedding matrix of three types of relations $\mathbf{E}^{0}=\{\mathbf{E}_{a}^{0},\mathbf{E}_{c}^{0},\mathbf{E}_{n}^{0}\}\in\mathbb{R}^{\left|\mathcal{R}_{a}\cup\mathcal{R}_{c}\cup\mathcal{R}_{n}\right|\times d}$.
-3:  Initialize the embedding matrix of relation nodes $\mathbf{H}_{r}^{0}\leftarrow\mathbf{E}_{a}^{0}\cdot\mathbf{W}_{r}\in\mathbb{R}^{\left|\mathcal{V}_{r}\right|\times d}$.
-4:  for  $h,r,t\in\mathcal{F}_{a}$ do
-5:     Initialize the embedding of fact nodes $\mathbf{h}_{f}^{0}\leftarrow f_{m}([\mathbf{h}_{h};\mathbf{h}_{r};\mathbf{h}_{t}])$. $\triangleright$ Representation Initialization
-6:  end for
-7:  for  $l\leftarrow 1$ to $L$ do
-8:     for $i\in\mathcal{V}$ do
-9:        $\mathcal{M}^{l}_{i}\leftarrow\mathbf{h}_{i}^{l-1}$. $\triangleright$ Intra-fact Message Passing
-10:        for  $i,r,j\in\mathcal{F}_{c}$ do
-11:           $\alpha_{i,j}^{l-1}\leftarrow\frac{\exp\left(\mathbf{W}^{l-1}\left(\sigma\left(\mathbf{W}_{in}^{l-1}\mathbf{h}_{i}^{l-1}+\mathbf{W}_{out}^{l-1}\mathbf{h}_{j}^{l-1}\right)\right)\right)}{\sum_{j^{\prime}\in\mathcal{N}_{i}}{\exp\left(\mathbf{W}^{l-1}\left(\sigma\left(\mathbf{W}_{in}^{l-1}\mathbf{h}_{i}^{l-1}+\mathbf{W}_{out}^{l-1}\mathbf{h}_{j^{\prime}}^{l-1}\right)\right)\right)}}$.
-12:           $\mathcal{M}^{l}_{i}\leftarrow\{\alpha_{i,j}^{l-1}\cdot\mathbf{W}_{out}^{l-1}\mathbf{h}_{j}^{l-1}\}\cup\mathcal{M}^{l}_{i}$.
-13:        end for
-14:     end for
-15:     for $i\in\mathcal{V}$ do
-[…]
-17:     end $\triangleright$ Inter-fact Message Passing
-18:     for $i\in\mathcal{V}$ do
-19:        $\mathcal{M}^{l}_{i}\leftarrow\{\mathbf{W}_{self}^{{l-1}}\mathbf{h}_{i}^{{l-1}}\}$.
-20:        for  $r,j\in\mathcal{N}_{i}$ do
-21:           $\mathbf{m}^{l-1}_{(i,r,j)}\leftarrow\sigma\left(\omega_{\tau\left(r\right)}^{l-1}\right)\mathbf{W}_{{\lambda}\left(r\right)}^{{l-1}}\phi\left(\mathbf{h}_{{j}}^{{l-1}},\mathbf{e}_{{r}}^{{l-1}}\right)$.
-22:           $\mathcal{M}^{l}_{i}\leftarrow\{\mathbf{m}^{l-1}_{(i,r,j)}\}\cup\mathcal{M}^{l}_{i}$.
-23:        end for
-24:     end for
-25:     for $i\in\mathcal{V}$ do
-[…]
-27:     end for
-28:     $\mathbf{E}^{l}\leftarrow\mathbf{W}_{rel}^{l-1}\mathbf{E}^{l-1}$.
-29:  end for
-30:  return  $\mathbf{H}^{L},\mathbf{E}^{L}$
-Appendix D Datasets Statistics
-Table 9 shows the details of the three hyper-relational knowledge graph benchmark datasets: WikiPeople, WD50K, JF17K, three nested factual knowledge graph benchmark datasets: FBH, FBHE, DBHE,and the temporal knowledge graph benchmark dataset wikidata12k.
-Among them, WikiPeople is a dataset derived from Wikidata Vrandečić and Krötzsch (2014) concerning entities type “human”.
-WikiPeople filter out the elements which have at least 30 mentions as key-value pairs.
-WD50K is a high-quality dataset extracting from Wikidata statements and avoiding the potential data leakage which allows triple-based models to memorize main fact in the H-Facts of test set.
-The “with Q(%)" column in Table 9 denote the number of facts with auxiliary key-value pairs and the “Arity" column denote range of the number of entities in hyper-relational facts.
-The nested factual knowledge graph datasets FBH and FBHE Chung and Whang (2023) are constructed based on FB15k237 Toutanova and Chen (2015) from Freebase Bollacker et al.
-(2008) while DBHE is based on DB15K Liu et al.
-(2019) from DBpedia Lehmann et al.
-(2015).
-FBH contains nested facts that can be only inferred inside the atomic facts, while FBHE and DBHE contain externally-sourced nested relation crawling from Wikipedia articles, e.g., NextAlmaMater and SucceededBy.
-Temporal knowledge graph dataset wikidata12K is also a subset of Wikidata Vrandečić and Krötzsch (2014), which represents the time information $\tau\in\mathcal{T}$ as time intervals.
-Appendix E Pseudo-code of HiSL
-The pseudo code of HiSL is outlined in Algorithm 1.
-Appendix F Hyperparameter Settings
+Appendix 0.C Hyperparameter Settings
 Here, we show the hyperparameter details for each link prediction task.

### v4

@@
 Abstract
-Beyond-triple fact representations including hyper-relational facts with auxiliary key-value pairs, temporal facts with additional timestamps, and nested facts implying relationships between facts, are gaining significant attention.
-However, constrained by complex fact representation forms, existing link prediction models for beyond-triple facts have difficulty achieving hierarchical fact modeling and generalizing the modules for one specific facts to other fact types.
-To overcome this limitation, we propose a Unified Hierarchical Representation learning framework (UniHR) for unified knowledge graph link prediction.
-It consists of a unified Hierarchical Data Representation (HiDR) module and a unified Hierarchical Structure Learning (HiSL) module as graph encoder.
+Real-world knowledge graphs (KGs) contain not only standard triple-based facts, but also more complex, heterogeneous types of facts, such as hyper-relational facts with auxiliary key-value pairs, temporal facts with additional timestamps, and nested facts that imply relationships between facts.
+These richer forms of representation have attracted significant attention due to their enhanced expressiveness and capacity to model complex semantics in real-world scenarios.
+However, most existing studies suffer from two main limitations: (1) they typically focus on modeling only specific types of facts, thus making it difficult to generalize to real-world scenarios with multiple fact types; and (2) they struggle to achieve generalizable hierarchical (inter-fact and intra-fact) modeling due to the complexity of these representations.
+To overcome these limitations, we propose UniHR, a Unified Hierarchical Representation learning framework, which consists of a learning-optimized Hierarchical Data Representation (HiDR) module and a unified Hierarchical Structure Learning (HiSL) module.
 The HiDR module unifies hyper-relational KGs, temporal KGs, and nested factual KGs into triple-based representations.
-Then HiSL incorporates intra-fact and inter-fact message passing, focusing on enhancing the semantic information within individual facts and enriching the structural information between facts.
-Empirical results demonstrate the effectiveness of UniHR and highlight the strong potential of unified representations.
-Code and data are available at https://github.com/Lza12a/UniHR.
-Keywords:
-Knowledge graph representation learning Beyond-triple knowledge graph Link prediction.
-1 Introduction
-Large-scale knowledge graphs (KGs) such as Freebase [3] and Wikidata [33] have been widely applied in many areas like question answering [17] and natural language processing [2].
-However, the presence of missing facts within these KGs inevitably limit their applications.
-Therefore, the link prediction task has been introduced to predict missing elements within factual data.
-Current link prediction methods mainly focus on the fact in the form of triple e.g., $(head\,entity,relation,$
-$tail\,entity)$.
-Figure 1: Comparison between the existing link prediction methods for specific beyond-triple knowledge graphs and our unified methods.
-Despite the simplicity and unity of triple-based representation, it is difficult to adequately express complex facts, such as “Oppenheimer is educated at Harvard University for a bachelor degree in chemistry" shown in Figure 1.
-Therefore, existing researches [34, 37, 38] contribute to focusing on semantically richer facts.
-Figure 1 illustrates three specific types of facts: hyper-relational fact ((Oppenheimer, educated at, Harvard University), degree: bachelor, major: chemistry), temporal fact (Oppenheimer, honored with, Fermi Prize, 1963), nested fact ((Oppenheimer, born in, New York), imply, (Oppenheimer, nationality, The United States)).
-These forms of facts allow for expression of complex semantics and revelation of relationships between facts, extending beyond the triple-based representation.
-Thus in recent years, Hyper-relational KGs (HKG) [6], Temporal KGs (TKG) [40], and Nested factual KGs (NKG) [37] attract wide research interests.
-Recent studies have demonstrated the effectiveness of various embedding strategies for these beyond-triple representations [36].
-However, these methods are hard to achieve hierarchical fact modeling and usually designed for specific representation forms, e.g., StarE [12] customizes graph neural network to enhance inter-fact structure information between hyper-relational facts.
-For NKGs, BiVE [7] bridges intra-fact semantics between two levels of facts through a simple MLP.
-Moreover, GeomE+ [40] et al.
-TKG methods employ time-aware scoring functions to capture intra-fact collaborative information between factual semantics and temporal information in the timestamp.
-Although these methods perform well on specific types of facts, it is evident that such customized methods are difficult to generalize to other types of KGs for intra-fact structure and inter-fact semantic enhancements.
-Therefore, establishing a unified hierarchical representation learning method for multiple types of KGs is worth to investigate.
-To overcome the challenges mentioned above, we propose a Unified Hierarchical Representation learning method (UniHR), which includes a Hierarchical Data Representation (HiDR) module and a Hierarchical Structure Learning (HiSL) module as the graph encoder.
+Then HiSL incorporates intra-fact and inter-fact message passing, focusing on enhancing both semantic information within individual facts and enriching the structural information between facts.
+To go beyond the unified method itself, we further explore the potential of unified representation in complex real-world scenarios, including joint modeling of multi-task, compositional and hybrid facts.
+Extensive experiments on 9 datasets across 5 types of KGs demonstrate the effectiveness of UniHR and highlight the strong potential of unified representations.
+Introduction
+Real-world large-scale knowledge graphs (KGs) such as Wikidata (Vrandečić and Krötzsch 2014) and DBpedia (Lehmann et al.
+2015) have been widely applied in many areas including question answering (Kaiser, Saha Roy, and Weikum 2021) and natural language processing (Annervaz, Chowdhury, and Dukkipati 2018; Liu et al.
+2025a).
+To faithfully represent complex real-world knowledge, these KGs usually incorporate not only standard triple-based facts, but also more complex and heterogeneous types of facts such as hyper-relational, temporal and nested facts.
+Figure 1: Comparison between the existing link prediction methods for specific beyond-triple KGs and our unified methods for more realistic KGs with multiple fact types.
+Despite the simplicity of triple-based representation (i.e., $(head,relation,tail)$), such forms struggle to capture the complexity of real-world facts, e.g., “Oppenheimer is educated at Harvard University for a bachelor degree in chemistry”.
+Consequently, recent studies (Xiong et al.
+2023b) have focused on semantically richer beyond-triple facts, including: hyper-relational fact ((Oppenheimer, educated at, Harvard University), degree: bachelor, major: chemistry), temporal fact (Oppenheimer, honored with, Fermi Prize, 1963), nested fact ((Oppenheimer, born in, New York), imply, (Oppenheimer, nationality, The United States)).
+These fact types allow for expression of complex semantics and revelation of relationships between facts.
+Thus in recent years, Hyper-relational KGs (HKG) (Luo et al.
+2023), Temporal KGs (TKG) (Zhang et al.
+2025a), and Nested factual KGs (NKG) (Li et al.
+2025) attract wide research interests.
+As shown in Figure 1, we find that existing research mainly suffers from two major limitations: (1) They fail to reflect real-world scenarios with multiple heterogeneous fact types (Xiong et al.
+2023b), instead artificially dividing and only modeling a single KG type; (2) Earlier triple-based studies (Chen et al.
+2021) have demonstrated the effectiveness of hierarchical fact semantic modeling (inter-fact and intra-fact).
+But due to the complexity of beyond-triple representations, they struggle to achieve comprehensive hierarchical semantic modeling, even generalizing to other fact types.
+Specifically, for HKGs, StarE (Galkin et al.
+2020) customizes GNN to enhance inter-fact interactions, while GRAN (Wang et al.
+2021) designs attention variants with edge-bias to capture intra-fact heterogeneity.
+For NKGs, NestE (Xiong et al.
+2024) et al.
+NKG methods connect bi-level facts and only score intra-fact semantics.
+For TKGs, either by explicitly incorporating temporal information into the score function like GeomE+ (Xu et al.
+2023), or by unfolding entity neighborhood subgraphs along temporal chain and capturing inter-fact semantics to model temporal information like ECEformer (Fang et al.
+2024).
+Although advanced methods like HAHE (Luo et al.
+2023) begin to capture hierarchical semantics for HKGs, their heterogeneity in representation limits their scalability to other fact types.
+Therefore, establishing a unified hierarchical representation learning method for real-world KG with multiple fact types is worth investigating.
+To fill this research gap, we propose UniHR, a Unified Hierarchical Representation learning framework, which includes a Hierarchical Data Representation (HiDR) module and a Hierarchical Structure Learning (HiSL) module.
 HiDR module standardizes hyper-relational facts, nested factual facts, and temporal facts into the form of triples without loss of information.
@@
 Finally, the updated embeddings are fed into decoders for link prediction.
-Empirical results demonstrate the UniHR achieves state-of-the-art or competitive performance on most metrics and highlight the strong potential of unified representations.
-Our contributions can be summarized as follows.
-  • 1.
-    We emphasize the value of investigating unified KG representation method, including unified symbolic representation and unfied representation learning method for different KGs.
-  • 2.
+Apart from the unification of method itself, UniHR’s unified representation enables flexible extensions.
+Unlike previous KG-specific models, UniHR accommodates more complex scenarios, such as compositional knowledge graphs, multi-task learning, and joint training on hybrid facts, thereby paving the way for pre-trained models across diverse KG types.
+Empirical results demonstrate the UniHR achieves competitive performance on most metrics without excessive increase in computational cost and highlight the strong potential of unified representations: Our contributions can be summarized as follows.
+  • We emphasize the value of investigating unified KG representation learning method, including unified symbolic representation and unfied learning for different KGs.
   • To our knowledge, we propose the first unified KG representation learning framework UniHR, across different types of KGs, including a hierarchical data representation module and a hierarchical structure learning module.
-  • 3.
   • We conduct link prediction experiments on 9 datasets across 5 types of KGs.
-Compared to methods designed for one kind of KG, UniHR achieves the best or competitive results, verifying strong generalization capability.
-2 Preliminaries
-In this section, we introduce the definition of four types of existing knowledge graphs (KGs): triple-based KG, hyper-relational KG, nested factual KG and temporal KG, along with link prediction tasks on these types of KGs.
+Compared to KG-specific methods, UniHR achieves the best or competitive results, verifying strong generalization capability.
+Preliminaries
 Link Prediction on Triple-based KG.
-A triple-based KG $\mathcal{G}_{KG}\,\text{=}\,\{\mathcal{V},\mathcal{R},\mathcal{F}\}$ represents facts as triples, denoted as =$\mathcal{F}\,\text{=}\{\left(h,r,t\right)|h,t\in\mathcal{V},$ $r\in\mathcal{R}\}$, where $\mathcal{V}$ is the set of entities and $\mathcal{R}$ is the set of relations.
-The link prediction on triple-based KGs involves answering a query $\left(h,r,?\right)$ or (?$,r,t\right)$, where the missing element ‘????’ is an entity in $\mathcal{V}$.
+A triple-based KG $\mathcal{G}_{KG}\,\text{=}\,\{\mathcal{V},\mathcal{R},\mathcal{F}\}$ represents facts as triples, denoted as $\mathcal{F}\,\text{=}\{\left(h,r,t\right)|h,t\in\mathcal{V},$ $r\in\mathcal{R}\}$, where $\mathcal{V}$ is the set of entities and $\mathcal{R}$ is the set of relations.
+The link prediction on triple-based KGs involves answering a query $\left(h,r,?\right)$ or $\left(?,r,t\right)$, where the missing element ‘??’ is an entity in $\mathcal{V}$.
 Link Prediction on Hyper-relational KG.
-A hyper-relational KG (HKG) $\mathcal{G}_{HKG}\,\text{=}\,\{\mathcal{V},\mathcal{R},\mathcal{F}\}$ consists of hyper-relational facts, abbreviated as H-Facts, denoted as =$\mathcal{F}\,\text{=}\{((h,r,t),\{(k_{i}\text{:}\,v_{i})\}_{i=1}^{m})$ $|\,h,t,v_{i}\in\mathcal{V},r,k_{i}\in\mathcal{R}\}$.
-Typically, we refer to $\left(h,r,t\right)$ as the main triple in the H-Fact and $\left\{\left(k_{i}\text{:}\,v_{i}\right)\right\}_{i=1}^{m}$ as $m$ auxiliary key-value pairs.
-Similar to link prediction on triple-based KGs, the link prediction on HKGs aims to predict entities in the main triple or the key-value pairs.
+A hyper-relational KG (HKG) $\mathcal{G}_{\text{HKG}}\,\text{=}\,\{\mathcal{V},\mathcal{R},\mathcal{F}\}$ consists of hyper-relational facts (H-Facts) $\mathcal{F}$, denoted as $\mathcal{F}\,\text{=}\{((h,r,t),$ $\{(k_{i}\text{:}\,v_{i})\}_{i=1}^{m})|\,h,t,v_{i}\in\mathcal{V},r,k_{i}\in\mathcal{R}\}$.
+Typically, we refer to $\left(h,r,t\right)$ as the main triple and $\left\{\left(k_{i}\text{:}\,v_{i}\right)\right\}_{i=1}^{m}$ as $m$ auxiliary key-value pairs.
+Link prediction on HKGs aims to predict entities in the main triple or the key-value pairs.
 Symbolically, the aim is to predict the missing element, denoted as ‘?’ for queries $((h,r,t),(k_{1}\text{:}\,v_{1}),\ldots(k_{i}\text{:}\,?)),$ $((?,r,t),\{(k_{i}\text{:}v_{i})\}_{i=1}^{m})$ or $((h,r,?),\{(k_{i}\text{:}v_{i})\}_{i=1}^{m})$.
 Link Prediction on Nested Factual KG.
-A nested factual KG (NKG) can be represented as $\mathcal{G}_{NKG}\,\text{=}\,\{\mathcal{V},\mathcal{R},\mathcal{F},\hat{\mathcal{R}},\hat{\mathcal{F}}\}$, which is composed of two levels of facts, called atomic facts and nested facts.
-=$\mathcal{F}\,\text{=}\left\{\left(h,r,t\right)|h,t\in\mathcal{V},r\in\mathcal{R}\right\}$ is the set of atomic facts, where $\mathcal{V}$ is a set of atomic entities and $\mathcal{R}$ is a set of atomic relations.
-^=$\hat{\mathcal{F}}\,\text{=}\,\{\left(\mathcal{F}_{i},\hat{r},\mathcal{F}_{j}\right)|$ $\mathcal{F}_{i},\mathcal{F}_{j}\in\mathcal{F},\hat{r}\in\hat{\mathcal{R}}\}$ is the set of nested facts, where $\hat{\mathcal{R}}$ is the set of nested relations.
-The link prediction on the NKGs is performed on the atomic facts or nested facts.
-We refer to the link prediction on atomic facts as Base Link Prediction, and the link prediction on nested facts as Triple Prediction.
-For base link prediction, given a query $\left(h,r,?\right)$ or (?$,r,t\right)$, the aim is to predict the missing atomic entity ‘????’ from $\mathcal{V}$.
-For triple prediction, given a query (?$,\hat{r},\mathcal{F}_{j}\right)$ or $\left(\mathcal{F}_{i},\hat{r},?\right)$, the aim is to predict the missing atomic fact ‘????’ from $\mathcal{F}$.
+[…]
 Link Prediction on Temporal KG.
-A temporal KG (TKG) $\mathcal{G}_{TKG}\,\text{=}\,\{\mathcal{V},\mathcal{R},\mathcal{F},\mathcal{T}\}$ is composed of quadruple-based facts, which can be represented as =$\mathcal{F}\,\text{=}\,\{(h,r,t,$
-$[\tau_{b},\tau_{e}])|h,t\in\mathcal{V},r\in\mathcal{R},\tau_{b},\tau_{e}\in\mathcal{T}\}$, where $\tau_{b}$ is the begin time, $\tau_{e}$ is the end time, $\mathcal{V}$ is the set of entities, $\mathcal{R}$ is the set of relations and $\mathcal{T}$ is the set of timestamps.
-The link prediction on TKGs aims to predict missing entities ‘????’ in $\mathcal{V}$ for two types of queries (?$,r,t,[\tau_{b},\tau_{e}]\right)$ or $\left(h,r,?,[\tau_{b},\tau_{e}]\right)$.
-3 Related Works
+A temporal KG (TKG) $\mathcal{G}_{\text{TKG}}\,\text{=}\,\{\mathcal{V},\mathcal{R},\mathcal{F},\mathcal{T}\}$ is composed of quadruple-based facts, represented as $\mathcal{F}\,\text{=}\,\{(h,r,t,[\tau_{b},\tau_{e}])|h,t\in\mathcal{V},$ $r\in\mathcal{R},\tau_{b},\tau_{e}\in\mathcal{T}\}$, where $\tau_{b}$ is the begin time, $\tau_{e}$ is the end time, $\mathcal{V}$ is the set of entities, $\mathcal{R}$ is the set of relations and $\mathcal{T}$ is the set of timestamps.
+The link prediction on TKGs aims to predict missing entities ‘??’ in $\mathcal{V}$ for two types of queries $\left(?,r,t,[\tau_{b},\tau_{e}]\right)$ or $\left(h,r,?,[\tau_{b},\tau_{e}]\right)$.
+Related Works
+This section discusses the limitations of existing methods in hierarchical modeling. See Appendix C for more discussion.
 Link Prediction on Hyper-relational Knowledge Graph.
-Earlier HKG representation learning methods e.g., m-TransH [35], RAE [44] have generalized the triple-based approach to HKG and loosely represent the combinations of key-value pairs.
+Early HKG methods typically focus on modeling either local or global information.
 Galkin et al.
-first customize StarE [12] based on CompGCN [31] for H-Facts to capture global structure information between H-Facts in the message passing stage, and achieves impressive results, demonstrating that the structure information of the graph in HKGs is also important.
-GRAN [15] introduces edge-aware bias into the vanilla transformer attention [32], while HyNT [6] designs a qualifier encoder for HKG.
-They both focus on intra-fact semantic dependencies but ignore the global structure information.
-In order to fully capture inter-fact and intra-fact information, HAHE [23] first adopts dual graph attention and edge-aware bias in the transformer attention layer for hierarchical modeling and achieves significant performance improvement.
-We consider the hierarchical attention mechanism of HAHE a good beginning, but it is customized for HKG and is difficult to generalize to other facts.
+(2020) customize StarE based on CompGCN (Vashishth et al.
+2019) for H-Facts to capture global structure information among H-Facts, demonstrating the importance of structure information in HKGs.
+GRAN (Wang et al.
+2021) with edge-aware bias in attention (Vaswani et al.
+2017) layer, HyNT (Chung, Lee, and Whang 2023) with qualifier-aware encoder, and ShrinkE with relation-specific box (Xiong et al.
+2023a) all focus on modeling intra-fact semantic information.
+Recent advanced methods aim to comprehensively capture both inter-fact and intra-fact information.
+For example, HAHE (Luo et al.
+2023) employs dual-graph attention and edge-aware bias in the transformer attention layer for hierarchical modeling, achieving significant performance improvements.
+Similarly, HyperSAT (Wang, Chen, and Zhang 2025) accomplishes this through a combination of graph sampling and a key-value joint attention mechanism.
+We consider hierarchical modeling of KGs to be a promising direction, existing approaches are customized for HKG form and difficult to generalize to other types of facts.
+Figure 2: Diverse beyond-triple facts are translated into the hierarchical data representation (HiDR) form.
 Link Prediction on Nested Factual Knowledge Graph.
 Chung et al.
-[7] first introduced nested facts to imply relationships between facts.
-They also propose BiVE which bridges semantics between atomic facts and fact nodes in the encoding phase via a simple MLP and scores both atomic facts and nested facts using the quaternion-based KGE scoring functions like QuatE [45] or BiQUE [16].
-Based on BiVE, NestE [37] represents the fact nodes as a $1\times 3$ embedding matrix and the nested relations as a $3\times 3$ matrix to avoid information loss, embedding them into hyperplanes with different dimensions.
-These methods only capture the intra-fact semantic information between atomic facts and nested facts.
-Meanwhile, due to the complexity of this representation, common triple-based GNNs have difficulty in message passing between atomic fact and nested fact for capturing inter-fact global information.
+(Chung and Whang 2023) are the first to introduce nested facts to model relationships between facts, and also propose BiVE which bridges semantics between atomic facts and fact nodes via a simple MLP and scores both atomic facts and nested facts using quaternion-based KGE scoring functions like QuatE (Zhang et al.
+2019) or BiQUE (Guo and Kok 2021).
+Based on BiVE, NestE (Xiong et al.
+2024) represents fact nodes using a $1\times 3$ embedding matrix and the nested relations as a $3\times 3$ matrix to avoid information loss.
+GRADATE (Li et al.
+2025) enhances entity and fact representation learning by mining latent intra-fact semantics.
+However, due to the complexity of NKG representations, existing methods have so far primarily focused on capturing intra-fact semantic information.
 Link Prediction on Temporal Knowledge Graph.
-Recent studies in temporal knowledge graph representation learning have focused on designing special time-aware scoring functions for inter-fact semantic modeling.
-Models such as TTransE [21], HyTE [8], TeRo [39], and TGeomE+ [40] incorporate temporal-aware module into the KGE score function in various ways.
-However, existing models seldom regard timestamps as node and directly utilize inter-fact message passing mechanism to perceive time information for enhancing entity and relation embeddings.
-Therefore, we believe it is a direction worth exploring.
-4 Methodology
-In this section, we introduce our method, a Unified Hierarchical Representation learning framework (UniHR), which includes a Hierarchical Data Representation (HiDR) module and a Hierarchical Structure Learning (HiSL) module.
-Our workflow can be divided into the following three steps: 1) Given a KG $\mathcal{G}$ of any type, we represent it into $\mathcal{G}^{HiDR}$ under the HiDR form.
-2) The $\mathcal{G}^{HiDR}$ will be encoded by HiSL module to enhance the semantic information within individual facts and structural information between facts on the whole graph.
-3) In the phase of decoding, the updated embeddings of nodes and edges are fed into transformer decoders to obtain the plausibility score of facts.
-Figure 2: Diverse facts are translated into HiDR form.
-4.1 Hierarchical Data Representation
-To overcome the differences in the representation of multiple types KGs, we introduce a Hierarchical Data Representation module, abbreviated as HiDR.
-Different from labelled RDF representation [1], we constrain “triple" to be considered as the basic units of HiDR form, then HiDR could continuous benefit from the model developments of triple-based KGs, which is the most active area about link prediction over KGs.
-As shown in Fig.
-2, in order to ensure comprehensive representation of facts, we introduce three hierarchical types of nodes and three connected relations in HiDR.
-Firstly, we denote original entities within three types of KGs as atomic nodes and complement fact nodes for HKGs and TKGs lacking a designated fact node.
-To facilitate the interaction between fact nodes and relations explicitly, we incorporate relation nodes into the graph, represented as $e_{r}$ for each $r$.
-These relation nodes are derived from transforming the relation edges in the original KG.
-It is important to facilitate direct access of fact nodes to the relevant atomic nodes during message passing process.
-To achieve this, we introduce three connected relations: has relation, has head entity and has tail entity, which establish directly connections between atomic nodes and fact nodes.
-Ultimately, we denote the (main) triple $\left(h,r,t\right)$ in original fact as three connected facts: $\left(f,has\,relation,e_{r}\right),\left(f,has\,head\,entity,h\right),$ $(f,has\,tail\,entity,t)$, and an atomic fact $(h,r,t)$, where $f$ is fact node.
-Formally, the definition of HiDR form is as follows:
-Definition 1
-Hierarchical Data Representation (HiDR): A KG represented as the HiDR form is denoted as $\mathcal{G}^{HiDR}\,\text{=}\,\{\mathcal{V}^{HiDR},\mathcal{R}^{HiDR},\mathcal{F}^{HiDR}\},$ where $\mathcal{V}^{HiDR}\,\text{=}\,$
-$\mathcal{V}_{a}\cup\mathcal{V}_{r}\cup\mathcal{V}_{f}$ is a joint set of atomic node set ($\mathcal{V}_{a}$), relation node set ($\mathcal{V}_{r}$), fact node set ($\mathcal{V}_{f}$).
-$\mathcal{R}^{HiDR}\,\text{=}\,\mathcal{R}_{a}\cup\mathcal{R}_{n}\cup\mathcal{R}_{c}$ is a joint set of atomic relation set ($\mathcal{R}_{a}$), nested relation set ($\mathcal{R}_{n}$), connected relation set $\mathcal{R}_{c}\,\text{=}\{has\,relation,has\,head\,entity,has\,$
-$tail\,entity\}$.
-The fact set $\mathcal{F}^{HiDR}\,\text{=}\,\mathcal{F}_{a}\cup\mathcal{F}_{c}\cup\mathcal{F}_{n}$ is jointly composed of three types of triple-based facts: atomic facts ($\mathcal{F}_{a}$), connected facts ($\mathcal{F}_{c}$) and nested facts ($\mathcal{F}_{n}$), where $\mathcal{F}_{a}\text{=}\,\{(v_{1},r,$ $v_{2})|\,v_{1},v_{2}\in\mathcal{V}_{a},r\in\mathcal{R}_{a}\}$, $\mathcal{F}_{c}\,\text{=}\,\{(v_{1},r,v_{2})|\,v_{1}\in\mathcal{V}_{f},r\in\mathcal{R}_{c},v_{2}\in\mathcal{V}_{a}\}$, $\mathcal{F}_{n}\,\text{=}\,\{(v_{1},r,v_{2})|\,v_{1},v_{2}\in\mathcal{V}_{f},r\in\mathcal{R}_{n}\}$.
-Next, we introduce how to transform different types of KGs into HiDR form.
-For hyper-relational knowledge graphs, we regard key-value pairs as complementary information for facts. Thus, we translate the H-Facts $\mathcal{F}_{HKG}\,\text{=}$
-$\{((h,r,t),\{(k_{i}\text{:}\,v_{i})\}_{i=1}^{m}\}$ into the HiDR form that $\mathcal{G}^{HiDR}_{HKG}\,\text{=}\,\{\mathcal{V},\mathcal{R},\mathcal{F}^{HiDR}_{HKG}\}$ following the definition, where $\mathcal{F}_{c}\,\text{=}\,\{(f,has\,relation,e_{r}),(f,has$ $\,head\,entity,h),(f,has$
-$tail\,entity,t),(f,k_{1},v_{1}),\ldots,(f,k_{m},v_{m})\}$, $\mathcal{F}_{a}\,\text{=}\,\{(h,r,t)\,|\,((h,r,t),\{(k_{i}\text{:}\,v_{i})\}_{i=1}^{m})\in\mathcal{F}_{HKG})\}$ and $\mathcal{F}_{n}\text{=}\,\varnothing$.
-For nested factual knowledge graphs, HiDR can naturally represent hierarchical facts, so we translate the atomic facts $\mathcal{F}_{NKG}\,\text{=}\,\{(h_{i},r_{i},t_{i})\}$ and the nested facts ^$\hat{\mathcal{F}}_{NKG}\,\text{=}\,\{((h_{1},r_{1},t_{1}),R,(h_{2},$ $r_{2},t_{2}))|(h_{i},r_{i},t_{i})\in\mathcal{F}_{NKG}\}$ into the form of HiDR that $\mathcal{G}^{HiDR}_{NKG}\,\text{=}\,\{\mathcal{V},\mathcal{R},\mathcal{F}^{HiDR}_{NKG}\}$ following the definition, where $\mathcal{F}_{a}\,\text{=}\,\{(h_{i},r_{i},t_{i})\,|\,(h_{i},r_{i},t_{i})\in$ $\mathcal{F}_{NKG}\}$, $\mathcal{F}_{c}\,\text{=}\{(f_{i},has\,head\,entity,h_{i}),(f_{i},has\,tail$
-$entity,t_{i}),(f_{i},has\,relation,e_{r_{i}})|f_{i}\,\text{=}\,(h_{i},r_{i},$ $t_{i})\in\mathcal{F}_{NKG}\}$ and $\mathcal{F}_{n}\text{=}\{(f_{1},R,f_{2})|f_{i}\in\mathcal{F}_{NKG}\}$.
-For temporal knowledge graphs, we regard the TKG as a special HKG, and convert timestamps to auxiliary key-value pairs in HKGs by adding two special atomic relations: begin and end, regarding timestamps as special numerical atomic nodes.
-Thus, we firstly translate the temporal facts in TKGs $\mathcal{F}_{TKG}\text{=}\,\{(h,r,t,[\tau_{b},\tau_{e}])\}$ into H-Facts form $\mathcal{F}_{TKG}^{HKG}\text{=}\,\{(h,r,t,begin\text{:}\tau_{b},end\text{:}\tau_{e})\}$.
-Then according to the previous transformation in HKG, it can be translated into the HiDR form that $\mathcal{G}_{TKG}^{HiDR}\,\text{=}\,\{\mathcal{V},\mathcal{R},\mathcal{F}_{TKG}^{HiDR}\}$ following the definition, where $\mathcal{F}_{a}\,\text{=}\{(h,r,t)\,|\,(h,r,t,begin\text{:}\,\tau_{b},end\text{:}$ $\tau_{e})\in\mathcal{F}_{TKG}^{HKG}\}$, $\mathcal{F}_{c}\,\text{=}\,\{(f,has\,relation,e_{r}),(f,$ $has\,head\,entity,h),\,(f,has\,tail\,entity,t),\,(f,$ $begin,\tau_{b}),(f,end,\tau_{e})\,|\,f\,\text{=}\,(h,r,t,$
-$begin\text{:}\,\tau_{b},end\text{:}\,\tau_{e})\in\mathcal{F}_{TKG}^{HKG}\}$ and $\mathcal{F}_{n}\,\text{=}\,\varnothing$.
-In summary, we could convert all above KGs into the HiDR form, and preserve the semantics in the original KGs without loss of information.
-Figure 3: HiSL module for intra-fact and inter-fact message passing.
-4.2 Hierarchical Structure Learning
+Recent studies on TKG representation learning have mainly focused on designing elegant time-aware modules to enhance representation capability.
+Advanced models like TGeomE+ (Xu et al.
+2023) improve the modeling of local semantics in TKGs through 4th-order tensor factorization and linear temporal regularization.
+Similarly, HGE (Pan et al.
+2024) and 5EL (Zhang et al.
+2025a) enhance the expressiveness of temporal spaces by introducing geometric spaces.
+In contrast, ECEformer (Fang et al.
+2024) captures only inter-fact semantics by unfolding entity neighborhood subgraphs along the temporal chain to model temporal information.
+However, TKG representation learning methods that can simultaneously capture both intra-fact temporal semantics and global structural information remain largely unexplored.
+UniHR achieves this by regarding timestamps as nodes and directly employing hierarchical message passing.
+Methodology
+In this section, we introduce UniHR, a Unified Hierarchical Representation learning framework, which includes a Hierarchical Data Representation (HiDR) module and a Hierarchical Structure Learning (HiSL) module.
+Our workflow includes the following three steps: (1) Given a KG $\mathcal{G}$ of any type, we represent it into HiDR form $\mathcal{G}^{\text{HiDR}}$.
+(2) The $\mathcal{G}^{\text{HiDR}}$ will be encoded by HiSL module to enhance the semantic information within individual facts and structural information between facts on the whole graph.
+(3) In the phase of decoding, the updated node and edge embeddings are serialized and fed into the transformer to optimize the model.
+Hierarchical Data Representation
+To overcome the limitations of beyond-triple representations in hierarchical modeling, we introduce a Hierarchical Data Representation module (HiDR), which is optimized for representation learning.
+Unlike existing triple-based systems (Ali et al.
+2022) including RDF (triple) reification, RDF-star and labeled RDF representation techniques, we constrain “triple” to serve as the basic units of HiDR form, and split “nodes” and “relations” into three types respectively, making it more suitable for graph representation learning.
+Meanwhile, “triple” form makes HiDR could continuously benefit from the model developments of triple-based KGs, which is the most active area of research for link prediction over KGs.
+As shown in Figure 2, we denote the entities in original KGs as atomic nodes and abstract fact nodes for HKGs and TKGs lacking a designated fact node.
+To facilitate the interaction between fact nodes and relations explicitly, we incorporate relation nodes into the graph, represented as $e_{r}$ for each relation $r$.
+To facilitate direct access of fact nodes to the relevant atomic nodes during message passing, we also introduce three connected relations: has relation, has head entity and has tail entity, which establish directly connections between atomic nodes and fact nodes.
+Ultimately, we denote the (main) triple $\left(h,r,t\right)$ in original fact as three connected facts: $\left(f,has\,relation,e_{r}\right),\left(f,has\,head\,entity,h\right),(f,has\,tail\,$
+$entity,t)$, and an atomic fact $(h,r,t)$, where $f$ is fact node. Formally, the definition of HiDR form is as follows:
+Definition 1.
+Hierarchical Data Representation (HiDR): A KG represented as the HiDR form is denoted as $\mathcal{G}^{\text{HiDR}}\,\text{=}\,$
+$\{\mathcal{V}^{\text{HiDR}},\mathcal{R}^{\text{HiDR}},\mathcal{F}^{\text{HiDR}}\},$ where $\mathcal{V}^{\text{HiDR}}\,\text{=}\,\mathcal{V}_{a}\cup\mathcal{V}_{r}\cup\mathcal{V}_{f}$ is a joint set of atomic node set ($\mathcal{V}_{a}$), relation node set ($\mathcal{V}_{r}$), fact node set ($\mathcal{V}_{f}$).
+$\mathcal{R}^{\text{HiDR}}\,\text{=}\,\mathcal{R}_{a}\cup\mathcal{R}_{n}\cup\mathcal{R}_{c}$ is a joint set of atomic relation set ($\mathcal{R}_{a}$), nested relation set ($\mathcal{R}_{n}$), connected relation set $\mathcal{R}_{c}\,\text{=}\{has\,relation,has\,head\,entity,has\,tail\,entity\}$.
+The fact set $\mathcal{F}^{\text{HiDR}}\,\text{=}\,\mathcal{F}_{a}\cup\mathcal{F}_{c}\cup\mathcal{F}_{n}$ is jointly composed of three types of triple-based facts: atomic facts ($\mathcal{F}_{a}$), connected facts ($\mathcal{F}_{c}$) and nested facts ($\mathcal{F}_{n}$), where $\mathcal{F}_{a}\text{=}\,\{(v_{1},r,$
+$v_{2})|\,v_{1},v_{2}\in\mathcal{V}_{a},r\in\mathcal{R}_{a}\}$, $\mathcal{F}_{c}\,\text{=}\,\{(v_{1},r,v_{2})|\,v_{1}\in\mathcal{V}_{f},r\in\mathcal{R}_{c},v_{2}\in\mathcal{V}_{a}\}$, $\mathcal{F}_{n}\,\text{=}\,\{(v_{1},r,v_{2})|\,v_{1},v_{2}\in\mathcal{V}_{f},r\in\mathcal{R}_{n}\}$.
+Figure 3: The overview of UniHR, including HiSL module for intra-fact and inter-fact message passing.
+Next, we introduce how to transform different types of beyond-triple KGs into HiDR form.
+For HKGs, we regard the key-value pairs as complementary information for facts.
+Thus, we translate the H-Facts $\mathcal{F}_{\text{HKG}}\,\text{=}\{((h,r,t),\{(k_{i}\text{:}\,v_{i})\}_{i=1}^{m}\}$ into the HiDR form that $\mathcal{G}^{\text{HiDR}}_{\text{HKG}}\,\text{=}\,\{\mathcal{V},\mathcal{R},\mathcal{F}^{\text{HiDR}}_{\text{HKG}}\}$ following the definition, where $\mathcal{F}_{c}\,\text{=}$
+$\{(f,has\,relation,e_{r}),(f,has\,head\,entity,h),(f,has\,tail\,$
+$entity,t),(f,k_{1},v_{1}),\ldots,(f,k_{m},v_{m})\}$, $\mathcal{F}_{a}\text{=}\{(h,r,t)\,|((h,$
+$r,t),\{(k_{i}\text{:}\,v_{i})\}_{i=1}^{m})\in\mathcal{F}_{\text{HKG}})\}$ and $\mathcal{F}_{n}\text{=}\,\varnothing$.
+For NKGs, HiDR can naturally represent the hierarchical facts, so we translate the atomic facts $\mathcal{F}_{\text{NKG}}\,\text{=}\,\{(h_{i},r_{i},t_{i})\}$ and the nested facts $\hat{\mathcal{F}}_{\text{NKG}}\,\text{=}\,\{((h_{1},r_{1},t_{1}),R,(h_{2},r_{2},t_{2}))|$
+$(h_{i},r_{i},t_{i})\in\mathcal{F}_{\text{NKG}}\}$ into the form of HiDR that $\mathcal{G}^{\text{HiDR}}_{\text{NKG}}\,\text{=}\,\{\mathcal{V},$
+$\mathcal{R},\mathcal{F}^{\text{HiDR}}_{\text{NKG}}\}$ following the definition, where $\mathcal{F}_{a}\,\text{=}\,\{(h_{i},r_{i},t_{i})|$
+$(h_{i},r_{i},t_{i})\in\mathcal{F}_{\text{NKG}}\}$, $\mathcal{F}_{c}\,\text{=}\{(f_{i},has\,head\,entity,h_{i}),(f_{i},$
+$has\,tail\,entity,t_{i}),(f_{i},has\,relation,e_{r_{i}})|f_{i}\,\text{=}\,(h_{i},r_{i},t_{i})\in$ $\mathcal{F}_{\text{NKG}}\}$ and $\mathcal{F}_{n}\text{=}\{(f_{1},R,f_{2})|f_{i}\in\mathcal{F}_{\text{NKG}}\}$.
+For TKGs, we regard the TKG as a special HKG, and convert timestamps to auxiliary key-value pairs in HKGs by adding two special atomic relations: begin and end, regarding timestamps as special numerical atomic nodes.
+Thus, we firstly translate the temporal facts in TKGs $\mathcal{F}_{\text{TKG}}\text{=}\,\{(h,r,t,[\tau_{b},\tau_{e}])\}$ into H-Facts form $\mathcal{F}_{\text{TKG}}^{\text{HKG}}\text{=}\,\{(h,r,$
+$t,begin\text{:}\tau_{b},end\text{:}\tau_{e})\}$.
+Then according to the previous transformation in HKG, it can be translated into the HiDR form that $\mathcal{G}_{\text{TKG}}^{\text{HiDR}}\text{=}\{\mathcal{V},\mathcal{R},\mathcal{F}_{\text{TKG}}^{\text{HiDR}}\}$ following the definition, where $\mathcal{F}_{a}\,\text{=}\,\{(h,r,t)\,|\,(h,r,t,begin\text{:}\,\tau_{b},end\text{:}\tau_{e})\in\mathcal{F}_{\text{TKG}}^{\text{HKG}}\}$, $\mathcal{F}_{c}\,\text{=}$
+$\{(f,has\,relation,e_{r}),(f,has\,head\,entity,h),\,(f,has\,tail\,$
+$entity,t),\,(f,begin,\tau_{b}),(f,end,\tau_{e})\,|\,f\,\text{=}\,(h,r,t,begin\text{:}\,\tau_{b},$
+$end\text{:}\,\tau_{e})\in\mathcal{F}_{\text{TKG}}^{\text{HKG}}\}$ and $\mathcal{F}_{n}\,\text{=}\,\varnothing$.
+In summary, HiDR serves as a module that dynamically transforms the original data into a unified representation optimized for the model, without altering the storage format of the original data.
+Moreover, from the perspective of graph learning, it fully preserves the semantics of the original knowledge graphs without any loss of information.
+Hierarchical Structure Learning
 It’s evident that HiDR form introduces many additional relation nodes and fact nodes.
-To avoid significantly increasing the model’s training parameters while fully capturing the hierarchy of HiDR form, we design a Hierarchical Structure Learning module, abbreviated as HiSL shown in Fig.
-3.
+To avoid significantly increasing the model’s training parameters while fully capturing the hierarchy of HiDR form, we design a Hierarchical Structure Learning module, abbreviated as HiSL shown in Figure 3.
 Representation Initialization.
-We first initialize the embedding matrices $\mathbf{H}_{a}\in\mathbb{R}^{|\mathcal{V}_{a}|\times d}$ and $\mathbf{E}\in\mathbb{R}^{|\mathcal{R}^{HiDR}|\times d}$ for atomic nodes and all relation edges.
+We first initialize the embedding matrices $\mathbf{H}_{a}\in\mathbb{R}^{|\mathcal{V}_{a}|\times d}$ and $\mathbf{E}\in\mathbb{R}^{|\mathcal{R}^{\text{HiDR}}|\times d}$ for atomic nodes and all relation edges.
 Then we also initialize the embedding of relation node $\mathbf{H}_{r}\in\mathbb{R}^{|\mathcal{V}_{r}|\times d}$, which can be transformed from the relation edge $r$ with a projection matrix $\mathbf{W_{r}}\in\mathbb{R}^{d\times d}$: $\mathbf{H}_{r}=\mathbf{E_{a}\cdot W_{r}}$, where $\mathbf{E_{a}}\subseteq\mathbf{E}$ is the atomic relation embeddings.
@@
 $\mathbf{h}_{f}=f_{m}([\mathbf{h}_{h};\mathbf{h}_{r};\mathbf{h}_{t}]),$ (1)
-where $(h,r,t)\in\mathcal{F}_{a}$, [;][$\cdot$;$\cdot$][ ⋅ ; ⋅ ] is the concatenation operation, $\mathbf{h}_{h},\mathbf{h}_{t}\subseteq\mathbf{H}_{a},\mathbf{h}_{r}\subseteq\mathbf{H}_{r}$ denote (main) triple embedding and $f_{m}\text{:}\;\mathbb{R}^{3d}\rightarrow\mathbb{R}^{d}$ is a 1-layer MLP.
+where $(h,r,t)\in\mathcal{F}_{a}$, $f_{m}\text{:}\;\mathbb{R}^{3d}\rightarrow\mathbb{R}^{d}$ is a 1-layer MLP, $\cdot$$\cdot$] is the concatenation, $\mathbf{h}_{h},\mathbf{h}_{t}\subseteq\mathbf{H}_{a},\mathbf{h}_{r}\subseteq\mathbf{H}_{r}$ denote (main) triple embedding.
 Therefore, the initialization of relation nodes and fact nodes is sufficiently parameter-efficient.
-For numerical atomic nodes, namely timestamps in temporal knowledge graphs, we encode the timestamp $\tau$ into an embedding with Time2Vec [18]:
+For numerical atomic nodes, namely timestamps in temporal knowledge graphs, we encode the timestamp $\tau$ into an embedding with Time2Vec (Kazemi et al. 2019):
 $\mathbf{h}_{\tau}=\omega_{p}\sin\left(f_{p}(\tau)\right)+f_{np}(\tau),$ (2)
-where $f_{p}\text{:}\;\mathbb{R}^{1}\rightarrow\mathbb{R}^{d}$ is a 1-layer MLP as periodic function, $f_{np}\text{:}\;\mathbb{R}^{1}\rightarrow\mathbb{R}^{d}$ is a 1-layer MLP as non-periodic function, and $\omega_{p}\in\mathbb{R}^{1}$ is a learnable parameter for scaling the periodic features.
+where $f_{p}\text{:}\;\mathbb{R}^{1}\rightarrow\mathbb{R}^{d}$ and $f_{np}\text{:}\;\mathbb{R}^{1}\rightarrow\mathbb{R}^{d}$ are both 1-layer MLPs as periodic and non-periodic functions, and $\omega_{p}\in\mathbb{R}^{1}$ is a learnable parameter for scaling the periodic features.
 Intra-fact Message Passing.
@@
 Given a fact node $f_{k}\in\mathcal{V}_{f}$, we construct its constituent elements, i.e., one-hop neighbors, as the node set $\mathcal{V}_{k}\,\text{=}\,\{v\in\mathcal{N}_{f_{k}}\,|\,v\in\mathcal{V}_{a}\cup\mathcal{V}_{r}\}$, where $\mathcal{N}_{f_{k}}$ is the set of one-hop neighbors of fact node $f_{k}$.
-Then we retain the edges directly connected to fact node $f_{k}$, thereby constructing a subgraph $\mathcal{G}_{k}\,\text{=}\,\{\mathcal{V}_{k},\mathcal{R}_{k},\mathcal{F}_{k}\}\subseteq\mathcal{G}^{HiDR}$.
-For this subgraph, we employ the graph attention network [5] to aggregate local information, computing the attention score $\alpha_{i,j}$ between node $i\in\mathcal{V}_{k}$ and its neighbor $j$.
+Then we retain the edges directly connected to fact node $f_{k}$, thereby constructing a subgraph $\mathcal{G}_{k}\,\text{=}\,\{\mathcal{V}_{k},\mathcal{R}_{k},\mathcal{F}_{k}\}\subseteq\mathcal{G}^{\text{HiDR}}$.
+WikiPeople WD50K
+Model subject/object all entities subject/object all entities
+MRR H@1 H@10 MRR H@1 H@10 MRR H@1 H@10 MRR H@1 H@10
+NaLP (Guan et al. 2019) 0.356 0.271 0.499 0.360 0.275 0.503 0.230 0.170 0.347 0.251 0.187 0.375
+StarE (Galkin et al. 2020) 0.458 0.364 0.611 - - - 0.309 0.234 0.452 - - -
+GRAN (Wang et al. 2021) 0.477 0.408 0.596 0.480 0.411 0.599 0.329 0.259 0.465 0.363 0.294 0.493
+tNaLP (Guan et al. 2023) 0.358 0.288 0.486 0.361 0.290 0.490 0.221 0.163 0.331 0.243 0.182 0.360
+HyNT (Chung, Lee, and Whang 2023) 0.479 0.411 0.601 0.478 0.409 0.601 0.328 0.256 0.468 0.356 0.285 0.493
+ShrinkE (Xiong et al. 2023a) 0.485 0.431 0.601 - - - 0.345 0.275 0.482 - - -
+HAHE (Luo et al. 2023) 0.498 0.418 0.610 0.497 0.421 0.614 0.343 0.269 0.484 0.378 0.306 0.515
+NYLON (Yu, Yang, and Yang 2024) 0.385 0.299 0.527 0.384 0.300 0.520 0.326 0.262 0.446 0.291 0.226 0.414
+HyperSAT (Wang, Chen, and Zhang 2025) 0.493 0.427 0.610 0.496 0.430 0.613 0.345 0.270 0.489 0.380 0.306 0.520
+UniHR 0.496 0.419 0.619 0.496 0.420 0.621 0.348 0.278 0.482 0.382 0.313 0.517
+Table 1: Results on HKG datasets, are reproduced by us and others are taken from (Wang, Chen, and Zhang 2025).
+For this subgraph, we employ the graph attention to aggregate local information, computing the attention score $\alpha_{i,j}$ between node $i\in\mathcal{V}_{k}$ and its neighbor $j$.
 The formula for calculating $\alpha_{i,j}$ in the $l$-th layer is as follows:
@@
 Then, the updated node embeddings are obtained by aggregating the information of neighbors according to the attention scores:
-$\mathbf{h}_{i}^{l}=\mathbf{h}_{i}^{l}+\sum\limits_{j\in\mathcal{N}_{i}}\alpha_{i,j}^{l}\cdot\mathbf{W}_{out}^{l}\mathbf{h}_{j}^{l}.$ (4)
+$\mathbf{h}_{i}^{l}=\mathbf{h}_{i}^{l}+\sum\limits_{j\in\mathcal{N}_{i}}\alpha_{i,j}^{l}\cdot\mathbf{W}_{out}^{l}\mathbf{h}_{j}^{l}.$
 Inter-fact Message Passing.
-At this stage, message passing is conducted on the whole graph $\mathcal{G}^{HiDR}$.
-Similar to previous work [31], we use a non-parametric aggregation operator $\phi\left(\cdot\right)\text{:}\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ to obtain messages from neighbouring nodes and edges.
+At this stage, message passing is conducted on the whole graph $\mathcal{G}^{\text{HiDR}}$.
+Specifically, we use a non-parametric aggregation operator $\phi\left(\cdot\right)\text{:}\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ to obtain messages from neighbouring nodes and edges.
 We employ the circular-correlation operator, defined as:
-$\phi\left(\mathbf{h}_{j},\mathbf{e}_{r}\right)=\mathbf{h}_{j}\star\mathbf{e}_{r}=\mathbf{F}^{-1}\left(\left(\mathbf{F}\mathbf{h}_{j}\right)\odot\overline{\left(\mathbf{F}\mathbf{e}_{r}\right)}\right),$ (5)
-where $\mathbf{F}$ and $\mathbf{F}^{-1}$ denote the discrete fourier transform (DFT) matrix and its inverse matrix, the $\odot$ is the element-wise (Hadamard) product.
-Furthermore, in order to fully capture the heterogeneity of the graph, we classify edges along two dimensions: $\lambda(r)\in\left\{forward,reverse\right\}$ and $\tau(r)\in\{connected\,\,relation,atomic\,\,$
-$relation,nested\,relation\}$ and adopt two relation-type specific learnable parameters $\mathbf{W}_{\lambda\left(r\right)}\in\mathbb{R}^{d\times d}$ and $\omega_{\tau\left(r\right)}\in$ $\mathbb{R}^{1}$ for more fine-grained aggregation:
-$\mathbf{h}_{i}^{l+1}=\sum\limits_{(r,j)\in\mathcal{N}(i)}{\sigma\left(\omega_{\tau\left(r\right)}^{l}\right)\mathbf{W}_{{\lambda}\left(r\right)}^{{l}}\phi\left(\mathbf{h}_{j}^{l},\mathbf{e}_{r}^{l}\right)}+\mathbf{W}_{self}^{l}\mathbf{h}_{i}^{l},$ (6)
-$\mathbf{e}_{r}^{l+1}=\mathbf{W}_{rel}^{l}\mathbf{e}_{r}^{l},$ (7)
-where $\mathbf{W}^{l}_{self},\mathbf{W}^{l}_{rel}\in\mathbb{R}^{d\times d}$, $\sigma$ is a sigmoid activation function and $\mathcal{N}\left(i\right)$ is a set of immediate neighbors of $i$ for its outgoing edges $r$.
+$\phi\left(\mathbf{h}_{j},\mathbf{e}_{r}\right)=\mathbf{h}_{j}\star\mathbf{e}_{r}=\mathbf{F}^{-1}\left(\left(\mathbf{F}\mathbf{h}_{j}\right)\odot\overline{\left(\mathbf{F}\mathbf{e}_{r}\right)}\right),$
+where $\mathbf{F}$ and $\mathbf{F}^{-1}$ denote the discrete fourier transform (DFT) matrix and its inverse matrix, and the $\odot$ is the element-wise product.
+In order to fully capture the graph’s heterogeneity, we classify edges along two dimensions: direction $\lambda(r)\in\left\{forward,reverse\right\}$ and type $\tau(r)\in\{connected\,\,relation,atomic\,\,relation,nested\,\,relation\}$ and adopt two relation-specific learnable parameters $\mathbf{W}_{\lambda\left(r\right)}$
+$\in\mathbb{R}^{d\times d}$ and $\omega_{\tau\left(r\right)}\in$ $\mathbb{R}^{1}$ for fine-grained aggregation:
+$\mathbf{h}_{i}^{l+1}=\sum\limits_{(r,j)\in\mathcal{N}(i)}{\sigma\left(\omega_{\tau\left(r\right)}^{l}\right)\mathbf{W}_{{\lambda}\left(r\right)}^{{l}}\phi\left(\mathbf{h}_{j}^{l},\mathbf{e}_{r}^{l}\right)}+\mathbf{W}_{self}^{l}\mathbf{h}_{i}^{l},$
+where $\mathbf{W}^{l}_{self}\in\mathbb{R}^{d\times d}$, $\sigma$ is a sigmoid function and $\mathcal{N}\left(i\right)$ is a set of immediate neighbors of $i$ for its outgoing edges $r$.
 We utilize $\phi\left(\cdot\right)$ to combine the information from edge $r$ and node $j$, and then passes it to node $i$ for update.
+Meanwhile, we update the relation representation as: $\mathbf{e}_{r}^{l+1}=\mathbf{W}_{rel}^{l}\mathbf{e}_{r}^{l}$.
 Through Intra-fact and Inter-fact two-stage message passing, nodes can fully capture both local semantic and global structural information.
-Moreover, its number of training parameters does not increase with the scale of the graph, thereby effectively adapting to the HiDR form.
-4.3 Link Prediction Decoder
-Since the query varies across different settings, we use the transformer [32] as the decoder with the mask pattern.
-Specifically, we convert the updated node and edge embeddings into a sequence of fact embeddings, mask the elements to be predicted in facts with the $\bm{\left[M\right]}$ token as the input to the transformer.
+Moreover, the number of training parameters does not increase with the scale of the graph, thereby effectively adapting to the HiDR form.
+Link Prediction Decoder
+Since the query varies across different settings, we use the transformer (Vaswani et al.
+2017) as the decoder with mask pattern.
+Specifically, we serialize the updated node and edge embeddings into a sequence of fact embeddings, mask the elements to be predicted in facts with the $\bm{\left[M\right]}$ token as the input.
 Finally, we obtain the embedding of output $\bm{\left[M\right]}$ in the last layer to measure the plausibility of the fact, denoted as $\mathbf{h}_{pre}$, and calculate the probability distribution of candidates, followed by training it using the cross-entropy loss:
-$P=\,{\rm Softmax}\left(f\left(\mathbf{h}_{pre}\right)[\mathbf{E};\mathbf{H}]^{\top}\right),$ (8)
 $\mathcal{L}=\sum_{t=0}^{\left|\mathcal{R}\right|+\left|\mathcal{V}\right|}{y_{t}\log P_{t}},$ (7)
-where $P\in\mathbb{R}^{\left|\mathcal{R}\right|+\left|\mathcal{V}\right|}$ represents the confidence scores of all candidates, $f\text{:}\;\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ is a 1-layer MLP, and $[\mathbf{E};\mathbf{H}]\in\mathbb{R}^{(\left|\mathcal{R}\right|+\left|\mathcal{V}\right|)\times d}$ is the embedding matrix of all candidate edges or nodes.
+where $P=\,{\rm Softmax}\left(f\left(\mathbf{h}_{pre}\right)[\mathbf{E};\mathbf{H}]^{\top}\right)\in\mathbb{R}^{\left|\mathcal{R}\right|+\left|\mathcal{V}\right|}$ represents the confidence scores of all candidates, $f\text{:}\;\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ is a 1-layer MLP, and $\mathbf{E};\mathbf{H}]\in\mathbb{R}^{(\left|\mathcal{R}\right|+\left|\mathcal{V}\right|)\times d}$ is the embedding matrix of all candidate edges or nodes.
 The $P_{t}$ and $y_{t}$ are probability and ground truth of the $t$-th candidate.
-The final loss function $\mathcal{L}$ includes both node loss and edge loss during the predictions.
-5 Experiment
-5.1 Experiment Settings
+Experiment
+Experiment Settings
 Datasets.
-For link prediction on HKGs, we select three benchmark datasets: WikiPeople [15] and WD50K [12].
-For link prediction on NKGs, we choose FBH, FBHE and DBHE constructed by [7].
-For link prediction on TKGs, we use wikidata12k [8], a subset of wikidata [33].
-To further evaluate the potential of the unified representation, we further introduce the hyper-relational TKG datasets WIKI-hy and YAGO-hy [10].
-The detailed statics are given in Table 1.
-Table 1: The statistics of diverse knowledge graphs dataset, where “with Q(%)" and “Arity" column respectively denote the number of facts with auxiliary key-value pairs and the range of arity of hyper-relational facts, “N-Fact” is the number of nested fact, “N-Rela” is the number of nested relation, the “AF(%)" column denotes the number of atomic facts in nested facts.
+For HKGs, we use WikiPeople (Guan et al.
+2019) and WD50K (Wang et al.
+2021).
+For NKGs, we select FBH, FBHE and DBHE (Chung and Whang 2023).
+For TKGs, we use wikidata12k (Dasgupta, Ray, and Talukdar 2018).
+To further evaluate the potential of the unified representation, we further introduce hyper-relational TKG datasets WIKI-hy and YAGO-hy (Ding et al.
+2024).
+Detailed statics of 9 benchmarks above are given in Appendix E.
+Evaluation Metric.
+We use the MR (Mean Rank), MRR (Mean Reciprocal Rank) and Hits@K (K=1,3,10) as our evaluation metrics.
+We abbreviate ‘Hits@K’ as ‘H@K’ and employ filtering settings (Bordes et al.
+2013) during the evaluation to eliminate existing facts in the dataset.
+It is worth noting that for the query $((h,r,?),{(k_{i}:v_{i})}_{i=1}^{m})$ in HKGs, there are two evaluation filtering settings in existing models: one that filters out facts satisfying $((h,r,?),{(k_{i}:v_{i})}_{i=1}^{m})$ and another that filters out facts satisfying only $(h,r,?)$ in the training set.
+Similarly, the difference in filtering settings of TKG occurs in timestamp.
+In this paper, we adopt the strict filtering setting of the former.
+To ensure fair comparison, for HKG we utilize the results of HyperSAT (Wang, Chen, and Zhang 2025) with the same settings as ours.
+For TKG, we thoroughly review the original code of our baselines and reproduce the results of some methods.
+Baselines.
+For HKG, we compare with NaLP (Guan et al.
+2019), StarE (Galkin et al.
+2020), GRAN (Wang et al.
+2021), tNaLP (Guan et al.
+2023), HyNT (Chung, Lee, and Whang 2023), ShrinkE (Xiong et al.
+2023a), HAHE (Luo et al.
+2023), NYLON (Yu, Yang, and Yang 2024) and HyperSAT (Wang, Chen, and Zhang 2025).
+For NKG, QuatE (Zhang et al.
+2019), BiQUE (Guo and Kok 2021), BiVE (Chung and Whang 2023), NestE (Xiong et al.
+2024), HOKE (Pirrò 2025) and GRADATE (Li et al.
+2025) are chosen as baselines.
+For TKG, we compare against following methods: ATiSE (Xu et al.
+2019), TGeomE+ (Xu et al.
+2023), HGE (Pan et al.
+2024), HyperSAT (Wang, Chen, and Zhang 2025), DuaTHP (Chen et al.
+2025), ECEformer (Fang et al.
+2024) and 5EL (Zhang et al.
+2025a).
+FBH DBHE FBH FBHE DBHE
+Model MRR H@10 MRR H@10 MR MRR H@10 MR MRR H@10 MR MRR H@10
+Base link prediction Triple prediction
+QuatE (Zhang et al. 2019) 0.354 0.581 0.264 0.440 145603.8 0.103 0.114 94684.4 0.101 0.209 26485.0 0.157 0.179
+BiQUE (Guo and Kok 2021) 0.356 0.583 0.274 0.446 81687.5 0.104 0.115 61015.2 0.135 0.205 19079.4 0.163 0.185
+BiVE (Chung and Whang 2023) 0.370 0.607 0.274 0.422 6.20 0.855 0.941 8.35 0.711 0.866 3.63 0.687 0.958
+NestE (Xiong et al. 2024) 0.371 0.608 0.289 0.443 3.34 0.922 0.982 3.05 0.851 0.962 2.07 0.862 0.984
+HOKE (Pirrò 2025) - - - - 3.06 0.719 0.777 2.82 0.674 0.764 2.10 0.674 0.777
+GRADATE (Li et al. 2025) - - - - 18.15 0.780 0.871 26.81 0.603 0.757 4.72 0.654 0.916
+UniHR 0.401 0.619 0.296 0.448 2.46 0.946 0.993 5.20 0.793 0.890 1.90 0.862 0.987
+Table 2: Results of base link prediction and triple prediction. Results of NKG-specific methods are taken from original papers.
+Implementation details.
+All experiments are conducted on a single Nvidia 80G A800 GPU and implemented with PyTorch.
+For base link prediction on NKGs, we also use augmented triples from (Chung and Whang 2023) for training to ensure fairness.
+For triple prediction, due to the small size of training set, we conduct training based on fixed embeddings of entities obtained from the base link prediction and set $\omega_{nested\;relation}\text{=}\ 0$ to prevent overfitting.
+Hyperparameter details can be found in Appendix F.
+Main Results
+Link Prediction on HKG.
+We compare our method with previous methods on the WD50K and WikiPeople datasets shown in Table 1.
+Among these methods, it can be seen that our proposed UniHR achieves competitive results with the state-of-the-art method HAHE and HyperSAT, which means our method effectively captures hierarchical fact information.
+Compared to GNN-based method StarE, we achieve improvements of 3.9 points (12.6%) in MRR, 4.4 points (18.8%) in Hits@1 and 3.0 points (6.6%) in Hits@10 on WD50K.
+This indicates that the performance of StarE’s customized GNN is limited by its inability to flexibly capture key-value pair information and hierarchical semantics.
+Model wikidata12k
+MRR H@1 H@3 H@10
+ATiSE (Xu et al. 2019) 0.252 0.148 0.288 0.462
+TGeomE+ (Xu et al. 2023) 0.333 0.232 0.361 0.546
+HGE (Pan et al. 2024) 0.290 0.176 0.323 0.514
+ECEformer (Fang et al. 2024) 0.262 0.159 0.255 0.462
+DuaTHP (Chen et al. 2025) 0.304 0.209 0.331 0.509
+HyperSAT (Wang, Chen, and Zhang 2025) 0.313 0.240 0.348 0.456
+5EL (Zhang et al. 2025a) 0.311 0.237 0.355 0.546
+UniHR 0.334 0.242 0.368 0.527
+Table 3: Results of link prediction on wikidata12k. Results are reported by us, and others are taken from original papers.
+Link Prediction on NKG.
+From the results in Table 2, we can see that our proposed UniHR obtains competitive results as the first method to capture global structural information of NKGs.
+For base link prediction task on triple-based KGs, UniHR achieves considerable improvements.
+Of particular note, the MRR of FBHE increases by 8.1%.
+For triple prediction, we perform best on FBH and DBHE datasets, especially obtaining an improvement of 2.4 points in MRR on FBH, and achieve the second-best performance on FBHE, which suggests that structural information is also valuable for NKG and UniHR can effectively capture the heterogeneity of NKG to enhance node embeddings.
+Unlike previous methods that use customized decoders for triples, our unified approach does not.
+The effectiveness of UniHR with other decoders is shown in Appendix A.
+Variant FBHE (NKG) DB15K (HKG) wikidata12k (TKG)
+MRR H@10 MRR H@10 MRR H@10
+w/o initial $\textbf{h}_{f}$ 0.767 0.885 0.346 0.481 0.333 0.525
+w/o $\textbf{W}_{r}$ 0.792 0.885 0.346 0.480 0.331 0.521
+w/o intra-fact MP 0.754 0.883 0.341 0.471 0.321 0.515
+w/o $\omega_{\tau(r)}$ 0.782 0.888 0.342 0.476 0.328 0.522
+w/o $\textbf{W}_{\lambda(r)}$ 0.778 0.889 0.341 0.474 0.327 0.521
+w/o inter-fact MP 0.776 0.887 0.338 0.468 0.319 0.511
+UniHR 0.793 0.890 0.348 0.482 0.334 0.527
+Table 4: Results of ablation studies on three KG types.
+Link Prediction on TKG.
+As shown in Table 3, we achieve competitive results on wikidata12k, even surpassing TGeomE+ by 4.3% on Hits@1 and 1.9% on Hits@3.
+However, existing TKG methods (e.g, TGeomE+ and HGE with temporal-augmented triple encoding, or ECEformer with temporal-guided subgraph encoding) only focus on partial factual semantics.
+In contrast, our approach efficiently encodes timestamps as atomic nodes only during initialization and learns temporal information through message passing on graph structure, demonstrating that graph structure information is also beneficial for temporal knowledge graphs, highlighting the effectiveness of our UniHR.
+Figure 4: (a) improvements of joint training on hybrid tasks.
+(b) improvements of joint training on wikimix dataset with hybrid fact forms.
+Yellow region indicates improvements achieved by joint training.
+Detailed results are shown in Appendix D.
+Ablation Study on HiSL
+To analyze the contribution of different modules across various KG types, we present ablation results in Table 4, with more studies provided in Appendix B.
+It can be observed that both intra-fact and inter-fact message passing contribute to performance improvement.
+In particular, intra-fact message passing proves to be more beneficial for NKGs.
+We attribute this to fact nodes in NKGs being inherently composed of other atomic nodes, making triple prediction rely heavily on comprehensive bi-level fact semantics.
+In contrast, HKGs and TKGs focus solely on atomic nodes, whose representations are not dependent on other nodes.
+Therefore, inter-fact message passing, by capturing the global context among facts, works more effective for HKGs and TKGs, leading to further performance improvements.
+Potential of Unified Representation
+Generalize to Compositional KGs.
+Owing to its unified representation, UniHR can flexibly generalize to compositional knowledge graphs, such as hyper-relational temporal KGs (HTKGs) (Ding et al.
+2024), which integrate the characteristics of both HKGs and TKGs.
+In HTKGs, each hyper-relational fact is associated with a timestamp that explicitly indicates its temporal validity.
+As shown in Table 5, UniHR offers a performance improvement in link prediction tasks on HTKGs, outperforming both TKG-specific and HKG-specific models.
+This result illustrates the strong ability of UniHR to jointly model auxiliary key-value pairs and temporal information.
+Furthermore, UniHR achieves competitive performance with the specialized model HypeTKG, despite not relying on complex module stacking.
+Model WiKi-hy YAGO-hy
+MRR H@1 H@3 H@10 MRR H@1 H@3 H@10
+HGE 0.602 0.507 0.666 0.765 0.790 0.760 0.814 0.837
+StarE 0.565 0.491 0.599 0.703 0.765 0.737 0.776 0.820
+GRAN 0.661 0.610 0.679 0.750 0.808 0.789 0.817 0.842
+HyNT 0.537 0.444 0.587 0.723 0.763 0.724 0.787 0.836
+HypeTKG 0.687 0.633 0.710 0.789 0.832 0.817 0.838 0.857
+UniHR 0.692 0.626 0.716 0.792 0.841 0.810 0.841 0.862
+Table 5: Results on hyper-relational TKG datasets.
+Joint Learning on Different Tasks of KGs.
+For link prediction on NKGs, the two subtasks, namely base link prediction and triple prediction, share the same graph during the message-passing phase under our unified representation form.
+Therefore, we attempt joint training on two tasks using the NKG dataset, as shown in Figure 4(b).
+Consistent with previous studies (Li et al.
+2025), we also observe that results of joint training are generally superior to those of separate training, further confirming that nested and atomic facts can mutually enhance and complement each other’s semantics.
+Joint Learning on Different Types of KGs.
+We believe unified representation is key to develop pre-trained models that integrate multiple KG types.
+To explore this potential, we jointly train on different KG types.
+Notably, real-world KGs like Wikidata (Vrandečić and Krötzsch 2014) naturally contain diverse fact types.
+Thus, we construct a hybrid dataset wikimix, by filtering two Wikidata subsets: HKG WikiPeople and TKG wikidata12k, which share 3,546 entities and 18 relations but have no overlapping facts.
+To prevent data leakage, we remove test entries whose main triples appear in the other subset’s train set (537 from wikidata12k and 384 from WikiPeople).
+Details are in Appendix E.
+Quantitative Analysis.
+As shown in Figure 4(a), we find joint learning outperforms separate learning across most metrics.
+Notably, MR improves by 17.1% on HKG and 39.7% on TKG, indicating that leveraging richer structural interactions across different fact types facilitates more effective representation learning.
+Visualization Analysis.
+As shown in Figure 5, entity embeddings from different categories are more coherently clustered and better separated under joint training compared to separate training, demonstrating that joint learning enables the model to acquire more structured and discriminative representations across diverse fact types.
+Figure 5: t-SNE visualization of shared entity’s embeddings.
+Figure 6: Results of efficiency analysis.
+Efficiency Analysis
+For memory usage, HiDR as a data preprocessing module, incurs minimal additional storage overhead.
+Although some extra nodes and relations are introduced, only embeddings for three “connected relations” need to be stored, while embeddings for other nodes can be derived from existing atomic elements.
+For runtime efficiency, as shown in Figure 6, UniHR does not significantly increase the number of model parameters or runtime compared to state-of-the-art methods.
+The embeddings of newly introduced nodes are computed from atomic elements, thus avoiding parameter inflation.
+During message passing, we employ subgraph sampling instead of using the entire graph, and apply dropout to prevent overfitting, which effectively improves training efficiency.
+Overall, UniHR achieves a better trade-off between effectiveness and efficiency.
+Conclusion
+In this paper, we propose UniHR, a unified hierarchical KG representation learning framework consisting of a learning-optimized Hierarchical Data Representation (HiDR) module and a unified Hierarchical Structure Learning (HiSL) module.
+The HiDR module unifies hyper-relational, nested and temporal facts into the triple form.
+Moreover, HiSL captures local semantic information within facts and global structural information between facts.
+Extensive experiments show UniHR achieves the best or competitive performance across 5 types of KGs over 9 datasets and further highlight the strong potential of unified representations across 3 complex scenarios.
+We hope UniHR will encourage research on unified KG representation learning and facilitate the development of pre-trained models for real-world KGs with heterogeneous fact types.
+Appendix A Appendix
+Appendix B A Decoder Analysis
+To explore the effectiveness of our UniHR encoding further, we pair UniHR with different decoders and evaluated them on triple prediction task.
+In addition to the previously mentioned unified framework UniHR + Transformer, we also experiment on UniHR + ConvE with two scoring strategies.
+The ConvE (Dettmers et al.
+2018) is the decoder customized for triples and its scoring function is $vec\left(\sigma\left(\left[\mathbf{\tilde{h}}_{h};\mathbf{\tilde{e}}_{r}\right]*\psi\right)\right)$, where $\mathbf{\tilde{h}}_{h}$ and $\mathbf{\tilde{e}}_{r}$ represent reshaped 2D embeddings of head entity $h$ and relation $r$, and * is a convolution operator.
+The $vec\left(\cdot\right)$ and $\psi$ are denoted as the vectorization function and a set of convolution kernels.
+Model FBHE/FBH DBHE
+MRR Hits@10 MRR Hits@10
+QuatE 0.354 0.581 0.264 0.440
+BiQUE 0.356 0.583 0.274 0.446
+BiVE 0.370 0.607 0.274 0.422
+NestE 0.371 0.608 0.289 0.443
+UniHR + ConvE $s_{h}$ 0.397 0.622 0.289 0.443
+UniHR + ConvE $s_{f}$ 0.375 0.596 0.307 0.471
+UniHR + Transformer 0.401 0.619 0.296 0.448
+Table 6: Base link prediction on FBHE, FBH and DBHE.
+All baselines’ results are taken from (Xiong et al.
+2024).
+The best results among all models are written bold, while the second are underlined.
+The $s_{f}$ and $s_{h}$ denote $\left(f,has\;head\;entity,h\right)\ \left(f,has\;tail\;entity,t\right)$ and $\left(h,r,t\right)$ two types of scoring method respectively.
+Due to our special representation, there exists two scoring methods for atomic triples, thus we present the base link prediction results separately for each scoring method.
+The $s_{f}$ represents scoring triples $\left(f,has\;head\;entity,h\right)\$and $\left(f,has\;tail\;entity,t\right)$, and $s_{t}$ represents scoring $\left(h,r,t\right)$.
+The performance of base link prediction is shown in Table 6.
+It can be observed that regardless of the scoring method employed, we both achieve competitive performance, especially with scoring $\left(h,r,t\right)$ on FBH and scoring $\left(f,has\;head\;entity,h\right)\,\left(f,has\;tail\;entity,t\right)$ on DBHE.
+We attribute the differences in performance under different scoring methods to dataset characteristics.
+The DBHE dataset is relatively smaller, and scoring method $s_{f}$ effectively alleviates overfitting problem.
+Conversely, for larger datasets FBH, scoring based on $\left(h,r,t\right)$ minimizes information loss.
+Table 8 shows the results of triple prediction on three benchmark datasets.
+Among all baselines, Quate and Bique struggle to model the mapping relationship between atomic facts and nested facts.
+Furthermore, prior works (Chung and Whang 2023) do not guarantee that all atomic facts in the nested fact test set are present in the training set as entities, which shifts the problem from a transductive setting to an inductive setting, leading to significant performance gaps between these baselines.
+On most metrics, our method outperforms BiVE and NestE which are specifically modeled for nested facts.
+Notably, NestE fully preserves the semantics of atomic facts.
+However, on the FBHE dataset, UniHR + ConvE achieves an improvement of 0.58 (6.4%) points in MRR and 0.24 (2.4%) points in Hits@10 compared to the state-of-the-art model NestE and the second-best performance after UniHR + Transformer on the FBH and DBHE datasets, demonstrating UniHR’s powerful graph structure encoding capabilities.
+We also carry out ablation experiments on UniHR + ConvE as shown in Table 8.
+Performance declines are observed after removing any part of the HiSL module, showing the significance of HiSL for hierarchical encoding.
+Variant FBH DBHE
+MR MRR H@10 MR MRR H@10
+w/o initial $\textbf{h}_{f}$ 5.22 0.909 0.980 2.56 0.794 0.978
+w/o $\textbf{W}_{r}$ 3.06 0.944 0.992 2.30 0.850 0.976
+w/o intra-fact MP 3.97 0.897 0.972 2.02 0.842 0.983
+w/o $\omega_{\tau(r)}$ 2.70 0.934 0.992 2.69 0.810 0.973
+w/o $\textbf{W}_{\lambda(r)}$ 2.50 0.941 0.992 2.37 0.810 0.978
+w/o inter-fact MP 2.61 0.913 0.991 2.11 0.827 0.986
+UniHR 2.46 0.946 0.993 1.90 0.862 0.987
+Table 7: Ablation studies on triple prediction task.
+Model FBH FBHE DBHE
+MR MRR Hits@10 MR MRR Hits@10 MR MRR Hits@10
+QuatE 145603.8 0.103 0.114 94684.4 0.101 0.209 26485.0 0.157 0.179
+BiQUE 81687.5 0.104 0.115 61015.2 0.135 0.205 19079.4 0.163 0.185
+BiVE 6.20 0.855 0.941 8.35 0.711 0.866 3.63 0.687 0.958
+NestE 3.34 0.922 0.982 3.05 0.851 0.962 2.07 0.862 0.984
+HOKE 3.06 0.719 0.777 2.82 0.674 0.764 2.10 0.674 0.777
+GRADATE 18.15 0.780 0.871 26.81 0.603 0.757 4.72 0.654 0.916
+UniHR + Transformer 2.46 0.946 0.993 5.20 0.793 0.890 1.90 0.862 0.987
+UniHR + ConvE 3.00 0.900 0.983 6.27 0.909 0.986 2.06 0.876 0.978
+UniHR + ConvE w/o $\textbf{h}_{f}$ 4.39 0.887 0.979 10.10 0.865 0.970 2.76 0.798 0.961
+UniHR + ConvE w/o intra-fact 6.54 0.859 0.959 18.10 0.871 0.968 5.82 0.665 0.900
+UniHR + ConvE w/o inter-fact 12.56 0.864 0.961 20.56 0.864 0.966 10.75 0.764 0.951
+Table 8: Triple prediction on FBHE, FBH and DBHE.
+All baselines’ results are taken from (Xiong et al.
+2024).
+The best results among all models are written bold.
+For BiVE (Chung and Whang 2023) and NestE (Xiong et al.
+2024), we pick their variants with best performance.
+Appendix C B More Ablation Study
+We present more ablation experiments on triple prediction.
+As shown in Table 7, all variants with certain modules or parameters removed exhibit a decrease in performance.
+We can conclude that intra-fact and inter-fact message passing modules both play crucial roles, allowing UniHR to more fully represent the current fact node with enhanced structural information.
+We also change the initialization of fact node embeddings to learnable embeddings (w/o initial $\textbf{h}_{f}$), and results indicate that initializing fact node representations is essential.
+It highlights key information in the facts, mitigating the noise introduced by excessive neighbors.
 Dataset Fact Entities Rela Train Valid Test with Q(%) Arity N-Fact N-Rela AF(%) Period
@@
 wikimix 409566 43832 184 326936 41777 3525/37328 9098(2.2%) 2-7 - - - 19-2020
-Evaluation Metric.
-We conduct link prediction across multiple settings, evaluating performance based on the rank of predicted facts.
-We use the MR (Mean Rank), MRR (Mean Reciprocal Rank) and Hits@K (K=1,3,10) as our evaluation metrics.
-We abbreviate ‘Hits@K’ as ‘H@K’ and employ filtering settings [4] during the evaluation to eliminate existing facts in the dataset.
-It is worth noting that for the query $((h,r,?),{(k_{i}:v_{i})}_{i=1}^{m})$ in HKGs, there are two evaluation filtering settings in existing models: one that filters out facts satisfying $((h,r,?),{(k_{i}:v_{i})}_{i=1}^{m})$ and another that filters out facts satisfying only $(h,r,?)$ in the training set.
-Similarly, the difference in filtering settings of TKG occurs in timestamp.
-In this paper, we adopt the strict filtering setting of the former.
-To ensure fair comparison, for HKG we utilize the results of HyNT [6] with the same settings as ours.
-For TKG, we thoroughly review the original code of our baselines and reproduce the results of some methods.
-Baselines.
-For link prediction on HKG, we compare our UniHR with NaLP [15], tNaLP [15], RAM [22], HINGE [27], NeuInfer [14], StarE [12], HyTransformer [43], GRAN [34], HyNT [6] and HAHE [23].
-For link prediction on NKG, QuatE [45], BiQUE [16], Neural-LP [42], DRUM [28], AnyBURL [24], BiVE [7] and NestE [37] are chosen as baselines.
-Among them, BiVE and NestE are especially designed for NKG.
-We compare against following TKG link prediction methods: ComplEx-N3 [20], HyTE [8], TA-DistMult [13], ATiSE [38], TeRo [39], TGeomE+ [40], HGE [26] and ECEformer [11].
-Implementation details.
-All experiments are conducted on a single Nvidia 80G A800 GPU and implemented with PyTorch.
-For base link prediction on NKGs, we also use augmented triples from [7] for training to ensure fairness.
-For triple prediction, due to the small size of training set, we conduct training based on fixed embeddings of entities obtained from the base link prediction and set $\omega_{nested\;relation}\text{=}\ 0$ to prevent overfitting.
-We employ AdamW [19] optimizer.
-Hyperparameter details can be found in Appendix C.
-5.2 Main Results
-Table 2: Results of link prediction on HKG datasets. Results are reported by us, and the others are taken from [6]. Best results are bold, while the second are underlined.
-WikiPeople WD50K
-Model subject/object all entities subject/object all entities
-MRR H@1 H@10 MRR H@1 H@10 MRR H@1 H@10 MRR H@1 H@10
-NaLP [15] 0.356 0.271 0.499 0.360 0.275 0.503 0.230 0.170 0.347 0.251 0.187 0.375
-tNaLP [15] 0.358 0.288 0.486 0.361 0.290 0.490 0.221 0.163 0.331 0.243 0.182 0.360
-RAM [22] 0.459 0.384 0.584 0.461 0.386 0.585 0.276 0.210 0.399 0.296 0.232 0.416
-HINGE [27] 0.393 0.309 0.547 0.395 0.311 0.549 0.264 0.187 0.410 0.277 0.200 0.424
-NeuInfer [14] 0.357 0.247 0.533 0.357 0.248 0.532 0.220 0.154 0.347 0.225 0.158 0.355
-StarE [12] 0.458 0.364 0.611 - - - 0.309 0.234 0.452 - - -
-HyTransformer [43] 0.465 0.395 0.598 - - - 0.308 0.239 0.452 - - -
-GRAN [34] 0.477 0.408 0.596 0.480 0.411 0.599 0.329 0.259 0.465 0.363 0.294 0.493
-HyNT [6] 0.482 0.415 0.602 0.481 0.414 0.603 0.333 0.259 0.474 0.360 0.287 0.500
-HAHE [23] 0.498 0.426 0.610 0.497 0.429 0.614 0.343 0.270 0.484 0.378 0.306 0.515
-UniHR 0.491 0.417 0.618 0.493 0.420 0.621 0.348 0.278 0.482 0.382 0.313 0.513
-Link Prediction on HKG.
-We compare our method with previous methods on the WD50K and WikiPeople datasets shown in Table 2.
-We can observe that the models based on transformers (i.e., StarE, GRAN, HyNT, HAHE) demonstrate significantly better performance compared to other models.
-We attribute this to the transformer’s superior ability to capture long distance dependencies within H-Facts.
-Among these methods, it can be seen that our proposed UniHR achieves competitive results with the state-of-the-art method HAHE, which means our method effectively captures hierarchical fact information.
-Compared to the GNN-based method StarE, we achieve improvements of 3.9 points (12.6%) in MRR, 4.4 points (18.8%) in Hits@1 and 3.0 points (6.6%) in Hits@10 on WD50K.
-This indicates that the performance of StarE’s customized graph neural network is limited by its inability to flexibly capture key-value pair information.
-Moreover, since the embeddings for newly added fact nodes and relation nodes are computed from atomic facts through simple MLPs, so our training parameters do not significantly increase.
-Table 3: Results of base link prediction (left) and triple prediction (right). All baselines’ results are taken from [37]. We pick the best variants of BiVE [7] and NestE [37].
-FBH DBHE FBH FBHE DBHE
-Model MRR H@10 MRR H@10 MR MRR H@10 MR MRR H@10 MR MRR H@10
-Base link prediction Triple prediction
-QuatE [45] 0.354 0.581 0.264 0.440 145603.8 0.103 0.114 94684.4 0.101 0.209 26485.0 0.157 0.179
-BiQUE [16] 0.356 0.583 0.274 0.446 81687.5 0.104 0.115 61015.2 0.135 0.205 19079.4 0.163 0.185
-Neural-LP [42] 0.315 0.486 0.233 0.357 115016.6 0.070 0.073 90000.4 0.238 0.274 21130.5 0.170 0.209
-DRUM [12] 0.317 0.490 0.237 0.359 115016.6 0.069 0.073 90000.3 0.261 0.274 21130.5 0.166 0.209
-AnyBURL [24] 0.310 0.526 0.220 0.364 108079.6 0.096 0.108 83136.8 0.191 0.252 20530.8 0.177 0.214
-BiVE [7] 0.370 0.607 0.274 0.422 6.20 0.855 0.941 8.35 0.711 0.866 3.63 0.687 0.958
-NestE [37] 0.371 0.608 0.289 0.443 3.34 0.922 0.982 3.05 0.851 0.962 2.07 0.862 0.984
-UniHR 0.401 0.619 0.296 0.448 2.46 0.946 0.993 5.20 0.793 0.890 1.90 0.862 0.987
-Link Prediction on NKG.
-Our experiments on the NKGs consist of two tasks: base link prediction and triple prediction.
-From the results in Table 3, we can see that our proposed UniHR obtains competitive results as the first method to capture global structural information of NKGs.
-For base link prediction task on triple-based KGs, UniHR achieves considerable improvements.
-Of particular note, the MRR of FBHE increases by 8.1%.
-For triple prediction, we perform best on FBH and DBHE datasets, especially obtaining an improvement of 2.4 points in MRR on FBH, and achieve the second-best performance on FBHE, which suggests that structural information is also valuable for NKG and UniHR can effectively capture the heterogeneity of NKG to enhance node embeddings.
-In particular, as a unified method, we do not use the customized decoder for triples, while previous state-of-the-art methods do.
-We will further illustrate the effectiveness of UniHR equipped with other decoders in Appendix B.
-Table 4: Results of link prediction on wikidata12k. Results are reported by us, and others are taken from original papers.
-Model wikidata12k
-MRR H@1 H@3 H@10
-ComplEx-N3 [20] 0.248 0.143 - 0.489
-HyTE [8] 0.253 0.147 - 0.483
-TA-DistMult [13] 0.230 0.130 - 0.461
-TeRo [39] 0.299 0.198 0.329 0.507
-ATiSE [38] 0.252 0.148 0.288 0.462
-TGeomE+ [40] 0.333 0.232 0.361 0.546
-HGE [26] 0.290 0.176 0.323 0.514
-ECEformer [11] 0.262 0.159 0.255 0.462
-UniHR 0.333 0.240 0.367 0.527
-Link Prediction on TKG.
-As shown in Table 4, we achieve competitive results on the wikidata12k, even surpassing TGeomE+ by 3.4% on Hits@1 and 1.7% on Hits@3.
-However, existing temporal knowledge graph embedding (KGE) methods, such as TGeomE+, often employ time-aware decoders, which are challenging to generalize to other types of KGs.
-In contrast, our approach efficiently encodes timestamps as atomic nodes only during initialization and learns temporal information through message passing on graph structure, demonstrating that graph structure information is also beneficial for temporal knowledge graphs, highlighting the effectiveness of our UniHR.
-Table 5: Ablation studies on triple prediction task.
-Variant FBH FBHE DBHE
-MR MRR H@10 MR MRR H@10 MR MRR H@10
-w/o initial $\textbf{h}_{f}$ 5.22 0.909 0.980 5.86 0.767 0.885 2.56 0.794 0.978
-w/o $\textbf{W}_{r}$ 3.06 0.944 0.992 5.98 0.792 0.885 2.30 0.850 0.976
-w/o intra-fact MP 3.97 0.897 0.972 5.26 0.754 0.883 2.02 0.842 0.983
-w/o $\omega_{\tau(r)}$ 2.70 0.934 0.992 5.46 0.782 0.888 2.69 0.810 0.973
-w/o $\textbf{W}_{\lambda(r)}$ 2.50 0.941 0.992 5.69 0.778 0.889 2.37 0.810 0.978
-w/o inter-fact MP 2.61 0.913 0.991 4.56 0.776 0.887 2.11 0.827 0.986
-UniHR 2.46 0.946 0.993 5.20 0.793 0.890 1.90 0.862 0.987
-5.3 Ablation Study on HiSL
-We conduct ablation experiments on triple prediction, the most relevant task to fact nodes.
-As shown in Table 5, all variants with certain modules or parameters removed exhibit a decrease in performance.
-We can conclude that intra-fact and inter-fact message passing modules both play crucial roles, allowing UniHR to more fully represent the current fact node with enhanced structural information.
-We also change the initialization of fact node embeddings to learnable embeddings (w/o initial $\textbf{h}_{f}$), and results indicate that initializing fact node representations is essential.
-It highlights key information in the facts, mitigating the noise introduced by excessive neighbors.
-5.4 Potential of Unified Representation
-Table 6: Results on hyper-relational TKG datasets.
-Model WiKi-hy YAGO-hy
-MRR H@1 H@3 H@10 MRR H@1 H@3 H@10
-TeRo [39] 0.572 0.473 0.640 0.727 0.760 0.720 0.782 0.822
-HGE [26] 0.602 0.507 0.666 0.765 0.790 0.760 0.814 0.837
-StarE [12] 0.565 0.491 0.599 0.703 0.765 0.737 0.776 0.820
-GRAN [34] 0.661 0.610 0.679 0.750 0.808 0.789 0.817 0.842
-HyNT [6] 0.537 0.444 0.587 0.723 0.763 0.724 0.787 0.836
-HypeTKG [10] 0.687 0.633 0.710 0.789 0.832 0.817 0.838 0.857
-UniHR 0.681 0.618 0.717 0.774 0.823 0.801 0.839 0.859
-Generalize to Hyper-relational TKGs.
-Due to the unity of representation, UniHR can easily generalize to compositional types of KGs, like hyper-relational TKGs (HTKGs) [10] that integrate HKGs and TKGs.
-Every H-fact in the HTKG is coupled with a timestamp explicitly indicating its time validity.
-From results in Table 6, UniHR offers a significant performance improvement in link prediction tasks on hyper-relational TKGs compared to TKG-specific and HKG-specific models, demonstrating a good collaborative modeling ability for auxiliary key-value pairs and timestamps.
-Our UniHR even achieves competitive performance with the specialized hyper-relational TKG model HypeTKG without complex module stacking.
-Table 7: Results of separate and joint training on NKG.
-FBH DBHE FBH DBHE
-Model MRR H@10 MRR H@10 MRR H@10 MRR H@10
-Base link prediction Triple prediction
-UniHR 0.401 0.619 0.296 0.448 0.946 0.993 0.862 0.987
-UniHRJoint 0.402 0.632 0.298 0.452 0.948 0.993 0.857 0.989
-Joint Learning on Different Tasks of KGs.
-For link prediction on NKGs, the two subtasks, namely base link prediction and triple prediction, share the same graph during the message-passing phase under our unified representation form.
-Therefore, we attempt joint training on the two tasks using the NKG dataset, as shown in Table 7.
-We observe that the results of joint training are generally superior to those of separate training, indicating that nested facts and atomic facts can mutually enhance and supplement the semantics.
-Table 8: Results of separate training and joint training on the HKG and TKG dataset, where identical entities and relations share the same embeddings.
-WikiPeople- and wikidata12k- represent the filtered test sets.
-WikiPeople- wikidata12k-
-Model subject/object all entities subject/object
-MR MRR H@1 H@3 H@10 MR MRR H@1 H@3 H@10 MR MRR H@1 H@3 H@10
-UniHR 835.8 0.486 0.412 0.528 0.617 829.0 0.488 0.414 0.531 0.620 818.7 0.314 0.220 0.345 0.509
-UniHRJoint 692.7 0.488 0.409 0.533 0.629 686.5 0.490 0.414 0.536 0.632 489.5 0.315 0.222 0.346 0.498
-Joint Learning on Different Types of KGs.
-We suppose that unified representation makes it possible to develop pre-trained models that integrate multiple types of KGs.
-To explore its potential, we conduct joint learning on different types of KGs.
-Therefore, we construct a hybrid dataset called wikimix which includes two subsets of Wikidata [33], namely HKG dataset WikiPeople and TKG dataset wikidata12k, which encompass 3547 identical entities and 18 identical relations.
-Due to the different types of facts, there are no identical facts in these two subsets.
-To further prevent data leakage, we filter out 537 entries from the wikidata12k test set whose main triples appear in the H-Facts of WikiPeople train set, and 384 entries from the WikiPeople test set whose main triples appear in wikidata12k train set.
-Statics of wikimix are given in Table 1.
-From the results in Table 8, it is evident that joint learning outperforms separate learning across most metrics.
-Notably, there are improvements of 17.1% and 40.2% in MR metric on wikipeople- and wikidata12k- datasets, respectively.
-This indicates that more complex structural interactions and diverse types of training data are beneficial.
-Moreover, our UniHR demonstrates good scalability and effectiveness in integrating multiple types of KGs.
-Table 9: Number of training parameters and times.
-Dataset Model Training Params Training Times
-WD50K StarE [12] 10.81M 48h+
-HyNT [6] 29.61M 22.3h
-HAHE [23] 42.35M 29.2h
-UniHR 10.55M 23.4h
-wikidata12k ATiSE [38] 31.46M 5.8h
-HGE [26] 3.44M -
-ECEformer [11] 11.99M 5.6h
-UniHR 3.68M 3.6h
-YAGO-hy HypeTKG [10] 10.83M 37.5h
-UniHR 4.37M 17.6h
-5.5 Efficiency Analysis
-Compared to existing state-of-the-art methods, UniHR does not significantly increase the overhead in terms of model training parameters and runtime as shown in Figure 9.
-The embeddings of relation nodes and fact nodes are computed using a simple MLP based on atomic nodes and relations, which does not significantly increase the number of training parameters.
-For runtime, we do not use the whole graph during all message passing phases and adopt a dropout strategy to prevent overfitting, which can effectively enhance training efficiency.
-Therefore, our UniHR achieves a superior trade-off between model effectiveness and efficiency.
-6 Conclusion and Future Work
-In this paper, we propose UniHR, a unified hierarchical knowledge graph representation learning framework consisting of a Hierarchical Data Representation (HiDR) module and a Hierarchical Structure Learning (HiSL) module.
-The HiDR form unifies the hyper-relational facts, nested facts and temporal facts into the form of triples, overcoming the limitations of customized encoders for different forms of facts.
-Moreover, HiSL captures local semantic information within facts and global structural information between facts.
-Our UniHR achieves the best or competitive performance across five types of KGs, and highlight the strong potential of unified representations.
-In the future, we believe that the unification of knowledge graph representation learning methods is a developing trend that makes it possible to develop unified pre-trained models based on multiple types of KGs.
-Moreover, we will investigate how to generalize our HiDR form to more complex tasks such as inductive reasoning [30] and multi-modal scenarios, etc.
-Supplemental Material Statement:
-The dataset, code, and hyperparameter details are available at https://github.com/Lza12a/UniHR.
-Appendix
-Appendix 0.A Related Works
-0.A.1 Link Prediction on Triple-based KGs
-Most existing techniques in KG representation learning are proposed for triple-based KGs.
-Among these techniques, knowledge graph embedding (KGE) models [4, 29] have received extensive attention due to their effectiveness and simplicity.
-The idea is to project entities and relations in the KG to low-dimensional vector spaces, utilizing KGE scoring functions to measure the plausibility of triples in the embedding space.
-Typical methods include TransE [4], RotatE [29], and ConvE [9].
-Depending on the KGE model alone has limitation of capturing complex graph structures, whereas augmenting global structural information with a graph neural network (GNN) [31, 25, 41] proves to be an effective approach for enhancement.
-The paradigm of combining GNN as encoder with KGE scoring function as decoder helps to enhance the performance of KGE scoring function.
-These GNN methods design elaborate message passing mechanisms to capture the global structural features.
-Typically, CompGCN [31] aggregates the joint embedding of entities and relations in the neighborhood via a parameter-efficient way and MA-GNN [41] learns global-local structural information based on multi-attention.
-These methods achieve impressive results on triple-based KGs but are hard to generalize to beyond-triple KGs.
-Table 10: Base link prediction on FBHE, FBH and DBHE.
-All baselines’ results are taken from [37].
-The best results among all models are written bold, while the second are underlined.
-The $s_{f}$ and $s_{h}$ denote $\left(f,has\;head\;entity,h\right)\ \left(f,has\;tail\;entity,t\right)$ and $\left(h,r,t\right)$ two types of scoring method respectively.
-For BiVE [7] and NestE [37], we pick their variants with best performance.
-Model FBHE/FBH DBHE
-MRR Hits@10 MRR Hits@10
-QuatE 0.354 0.581 0.264 0.440
-BiQUE 0.356 0.583 0.274 0.446
-Neural-LP 0.315 0.486 0.233 0.357
-DRUM 0.317 0.490 0.237 0.359
-AnyBURL 0.310 0.526 0.220 0.364
-BiVE 0.370 0.607 0.274 0.422
-NestE 0.371 0.608 0.289 0.443
-UniHR + ConvE $s_{h}$ 0.397 0.622 0.289 0.443
-UniHR + ConvE $s_{f}$ 0.375 0.596 0.307 0.471
-UniHR + Transformer 0.401 0.619 0.296 0.448
-Appendix 0.B Decoder Analysis
-To explore the effectiveness of our UniHR encoding further, we pair UniHR with different decoders and evaluated them on triple prediction task.
-In addition to the previously mentioned unified framework UniHR + Transformer, we also experiment on UniHR + ConvE with two scoring strategies.
-The ConvE [9] is the decoder customized for triples and its scoring function is $vec\left(\sigma\left(\left[\mathbf{\tilde{h}}_{h};\mathbf{\tilde{e}}_{r}\right]*\psi\right)\right)$, where $\mathbf{\tilde{h}}_{h}$ and $\mathbf{\tilde{e}}_{r}$ represent reshaped 2D embeddings of head entity $h$ and relation $r$, and * is a convolution operator.
-The $vec\left(\cdot\right)$ and $\psi$ are denoted as the vectorization function and a set of convolution kernels.
-Due to our special representation, there exists two scoring methods for atomic triples, thus we present the base link prediction results separately for each scoring method.
-The $s_{f}$ represents scoring triples $\left(f,has\;head\;entity,h\right)$and $\left(f,has\;tail\;entity,t\right)$, and $s_{t}$ represents scoring $\left(h,r,t\right)$.
-The performance of base link prediction is shown in Table 10.
-Notably, FBH and FBHE share identical atomic facts, resulting in the same performance.
-It can be observed that regardless of the scoring method employed, we both achieve competitive performance, especially with scoring $\left(h,r,t\right)$ on FBH and scoring $\left(f,has\;head\;entity,h\right)\,\left(f,has\;tail\;entity,t\right)$ on DBHE.
-We attribute the differences in performance under different scoring methods to dataset characteristics.
-The DBHE dataset is relatively smaller, and scoring method $s_{f}$ effectively alleviates overfitting problem.
-Conversely, for larger datasets FBH, scoring based on $\left(h,r,t\right)$ minimizes information loss.
-Table 11: Triple prediction on FBHE, FBH and DBHE.
-All baselines’ results are taken from [37].
-The best results among all models are written bold.
-For BiVE [7] and NestE [37], we pick their variants with best performance.
-Model FBH FBHE DBHE
-MR MRR Hits@10 MR MRR Hits@10 MR MRR Hits@10
-QuatE [45] 145603.8 0.103 0.114 94684.4 0.101 0.209 26485.0 0.157 0.179
-BiQUE [16] 81687.5 0.104 0.115 61015.2 0.135 0.205 19079.4 0.163 0.185
-Neural-LP [42] 115016.6 0.070 0.073 90000.4 0.238 0.274 21130.5 0.170 0.209
-DRUM [28] 115016.6 0.069 0.073 90000.3 0.261 0.274 21130.5 0.166 0.209
-AnyBURL [24] 108079.6 0.096 0.108 83136.8 0.191 0.252 20530.8 0.177 0.214
-BiVE [7] 6.20 0.855 0.941 8.35 0.711 0.866 3.63 0.687 0.958
-NestE [37] 3.34 0.922 0.982 3.05 0.851 0.962 2.07 0.862 0.984
-UniHR + Transformer 2.46 0.946 0.993 5.20 0.793 0.890 1.90 0.862 0.987
-UniHR + ConvE 3.00 0.900 0.983 6.27 0.909 0.986 2.06 0.876 0.978
-UniHR + ConvE w/o $\textbf{h}_{f}$ 4.39 0.887 0.979 10.10 0.865 0.970 2.76 0.798 0.961
-UniHR + ConvE w/o intra-fact 6.54 0.859 0.959 18.10 0.871 0.968 5.82 0.665 0.900
-UniHR + ConvE w/o inter-fact 12.56 0.864 0.961 20.56 0.864 0.966 10.75 0.764 0.951
-Table 11 shows the results of triple prediction on three benchmark datasets.
-Among all baselines, Quate, Bique, Neural-LP, Drum, and AnyBURL struggle to model the mapping relationship between atomic facts and nested facts.
-Furthermore, prior works [7] do not guarantee that all atomic facts in the nested fact test set are present in the training set as entities, which shifts the problem from a transductive setting to an inductive setting, leading to significant performance gaps between these baselines.
-On most metrics, our method outperforms BiVE and NestE which are specifically modeled for nested facts.
-Notably, NestE fully preserves the semantics of atomic facts.
-However, on the FBHE dataset, UniHR + ConvE achieves an improvement of 0.58 (6.4%) points in MRR and 0.24 (2.4%) points in Hits@10 compared to the state-of-the-art model NestE and the second-best performance after UniHR + Transformer on the FBH and DBHE datasets, demonstrating UniHR’s powerful graph structure encoding capabilities.
-We also carry out ablation experiments on UniHR + ConvE as shown in Table 11.
-Performance declines are observed after removing any part of the HiSL module, showing the significance of HiSL for hierarchical encoding.
-Table 12: The major hyperparameters of our approach for all link prediction tasks.
+Table 9: The statistics of diverse knowledge graphs dataset, where “with Q(%)” and “Arity” column respectively denote the number of facts with auxiliary key-value pairs and the range of arity of hyper-relational facts, “N-Fact” is the number of nested fact, “N-Rela” is the number of nested relation, the “AF(%)” column denotes the number of atomic facts in nested facts.
 Hyperparameter WikiPeople WD50K wikidata12k FBHEbase FBHbase DBHEbase FBHEtriple FBHtriple DBHEtriple
@@
 learning rate 5e-4 5e-4 5e-4 5e-4 5e-4 5e-4 5e-4 5e-4 5e-4
-Appendix 0.C Hyperparameter Settings
+Table 10: The major hyperparameters of our approach for all link prediction tasks.
+WikiPeople wikidata12k
+Model subject/object all entities subject/object
+MR MRR H@1 H@3 H@10 MR MRR H@1 H@3 H@10 MR MRR H@1 H@3 H@10
+UniHR 835.8 0.486 0.412 0.528 0.617 829.0 0.488 0.414 0.531 0.620 818.7 0.314 0.220 0.345 0.509
+UniHRJoint 692.7 0.489 0.409 0.530 0.629 686.5 0.493 0.418 0.536 0.632 493.6 0.317 0.224 0.348 0.503
+Table 11: Results of separate training and joint training on hybrid KG dataset wikimix, where identical entities and relations share the same embeddings.
+WikiPeople and wikidata12k represent the filtered test sets.
+FBH DBHE FBH DBHE
+Model MRR H@10 MRR H@10 MRR H@10 MRR H@10
+Base link prediction Triple prediction
+UniHR 0.401 0.619 0.296 0.448 0.946 0.993 0.862 0.987
+UniHRJoint 0.404 0.632 0.298 0.454 0.949 0.994 0.860 0.989
+Table 12: Results of separate and joint training on NKG.
+Appendix D C More Related Works
+Link Prediction on Triple-based KGs.
+Most existing techniques in KG representation learning are proposed for triple-based KGs.
+Among these techniques, knowledge graph embedding (KGE) models (Bordes et al.
+2013; Sun et al.
+2018) have received extensive attention due to their effectiveness and simplicity.
+The idea is to project entities and relations in the KG to low-dimensional vector spaces, utilizing KGE scoring functions to measure the plausibility of triples in the embedding space.
+Typical methods include TransE (Bordes et al.
+2013), RotatE (Sun et al.
+2018), and ConvE (Dettmers et al.
+2018).
+Depending on the KGE model alone has limitation of capturing complex graph structures, whereas augmenting global structural information with a graph neural network (GNN) (Vashishth et al.
+2019; Nathani et al.
+2019; Xu, Bao, and Liu 2023) proves to be an effective approach for enhancement.
+The paradigm of combining GNN as encoder with KGE scoring function as decoder helps to enhance the performance of KGE scoring function.
+These GNN methods design elaborate message passing mechanisms to capture the global structural features.
+Typically, CompGCN (Vashishth et al.
+2019) aggregates the joint embedding of entities and relations in the neighborhood via a parameter-efficient way and MA-GNN (Xu, Bao, and Liu 2023) learns global-local structural information based on multi-attention.
+These methods achieve impressive results on triple-based KGs but are challenging to generalize to beyond-triple KGs.
+Additionally, these GNNs usually focus more on modeling global information while neglecting local information.
+Why We Need Unified Representation?
+Firstly, real-world data sources naturally contain heterogeneous data forms (Liu et al.
+2025b).
+For example, Wikidata (Vrandečić and Krötzsch 2014) includes triple-based facts, hyper-relational facts, temporal facts, and even nested facts, financial reports and research papers usually contain various modalities such as tables, images, and text.
+Unified representation aims to integrate multiple heterogeneous data within a single semantic structure, avoiding the need for cross-type data conversion or independent processing during use.
+In many real-world scenarios (e.g., question answering, recommendation, or transfer learning), the research community is constantly exploring the benefits of unified representations.
+For question answering (QA), TrustUQA (Zhang et al.
+2025b) demonstrates a unified representation allows the knowledge base to contain a broader range of knowledge types, thereby simplifying the burden of the knowledge retrieval module and enabling the QA system to cover a wider variety of questions and answers.
+For recommendation systems, DURation (Lu et al.
+2022) with unified representation effectively eliminate information redundancy caused by separate module designs and help capture cross-dimensional relational knowledge, thus improving the comprehensiveness and accuracy of recommendations.
+Moreover, UniHR not only meets the requirements of real-world KG reasoning tasks (involving multiple types of facts), but we also believe that unified representation itself is beneficial for transfer learning.
+It not only facilitates transfer learning across different types of KGs, but also enables pre-training across multiple KG types (as demonstrated in Section 5.4).
+This further promotes more effective pre-training and transfer learning in the field of knowledge graphs.
+Appendix E D Detailed Results of Our “Potential”
+Table 11 shows the results of “Joint Learning on Different Tasks of KGs” section. Table 12 shows the results of “Joint Learning on Different Types of KGs” section.
+Appendix F E Dataset Details
+Table 9 shows the details of the two hyper-relational knowledge graph (HKG) benchmark datasets: WikiPeople, WD50K three nested factual knowledge graph (NKG) benchmark datasets: FBH, FBHE, DBHE, and the temporal knowledge graph (TKG) benchmark dataset wikidata12k.
+Among them, WikiPeople is a dataset derived from Wikidata (Vrandečić and Krötzsch 2014) concerning entities type “human”.
+WikiPeople filter out the elements which have at least 30 mentions as key-value pairs.
+WD50K is a high-quality dataset extracting from Wikidata statements and avoiding the potential data leakage which allows triple-based models to memorize main fact in the H-Facts of test set.
+The “with Q(%)” column in Table 9 denote the number of facts with auxiliary key-value pairs and the “Arity” column denote range of the number of entities in hyper-relational facts.
+Nested factual knowledge graph datasets FBH and FBHE (Chung and Whang 2023) are constructed based on FB15k237 from Freebase (Bollacker et al.
+2008) while DBHE is based on DB15K from DBpedia (Lehmann et al.
+2015).
+FBH contains nested facts that can be only inferred inside the atomic facts, while FBHE and DBHE contain externally-sourced nested relation crawling from Wikipedia articles, e.g., NextAlmaMater and SucceededBy.
+Temporal knowledge graph dataset wikidata12K is also a subset of Wikidata (Vrandečić and Krötzsch 2014), which represents the time information $\tau\in\mathcal{T}$ as time intervals.
+Appendix G F Hyperparameter Settings
 Here, we show the hyperparameter details for each link prediction task.